### v1

# Reflection of Nichols Algebras over Coquasi-Hopf Algebras

Abstract
This paper extends the foundational reflection theory of Nichols algebras to the setting of some certain coquasi-Hopf algebras. Our primary motivation arises from the classification of pointed finite-dimensional coquasi-Hopf algebras. We develop a reflection theory for tuples of simple Yetter-Drinfeld modules in the category $^{G}_{G}\mathcal{YD}^{\Phi}$}, where $G$ is a finite group and $\Phi$ is a 3-cocycle on $G$. We prove that such a tuple gives rise to a semi-Cartan graph if admitting all reflections. Consequently, its Weyl groupoid is well-defined. We further establish several criteria for the finite-dimensionality of Nichols algebras in terms of the associated semi-Cartan graph. As an application, we provide a new proof for the infinite-dimensionality of a specific class of Nichols algebras previously studied in huang2024classification, bypassing extensive computational arguments.
1 Introduction
The classification of finite-dimensional pointed Hopf algebras has been one of the central problems in the theory of Hopf algebras over the past three decades. A crucial breakthrough in this direction was achieved through the systematic study of Nichols algebras. Heckenberger introduced the revolutionary concepts of semi-Cartan graphs and Weyl groupoids [Hec06, Hec09, Hec10], which provide powerful combinatorial tools for understanding the structure and classification of Nichols algebras.
The reflection theory of Nichols algebras represents one of the most important developments in the field. The fundamental theorem states that a tuple of simple Yetter-Drinfeld modules over Hopf algebra with bijective antipode will give rise to a semi-Cartan graph provided that it admits all reflections [AHS10,HS10]. This result was proved again using a categorical approach in the setting of Hopf algebras in HSdualpair.
We recall that a semi-Cartan graph consists of the following data,
Definition 1.1.
Let $\mathbb{I}$ be a non-empty finite set, $\mathcal{X}$ a non-empty set, $r:\mathbb{I}\times\mathcal{X}\rightarrow\mathcal{X}$, $A:\mathbb{I}\times\mathbb{I}\times\mathcal{X}\rightarrow\mathbb{Z}$ maps. For all $i,j\in\mathbb{I}$ and ${X}\in\mathcal{X}$ we write $r_{i}(X)=r(i,X)$, $a_{ij}^{X}=A(i,j,X)$ and $A^{X}=(a_{ij}^{X})_{i,j\in\mathbb{I}}\in\mathbb{Z}^{\mathbb{I}\times\mathbb{I}}$. The quadruple $\mathcal{G}=\mathcal{G}(\mathbb{I},\mathcal{X},(r_{i}),(A^{X})_{X\in\mathcal{X}})$ is called a semi-Cartan graph if for all $X\in\mathcal{X}$, the matrix $A^{X}$ is a generalized Cartan matrix, and the following axioms hold.
$(CG1)$ For all $i\in I$, $r_{i}^{2}=\mathrm{id}_{\mathcal{X}}$.
$(CG2)$ For all $i\in I$, $X\in\mathcal{X}$, $A^{X}$ and $A^{r_{i}(X)}$ have the same $i$-th row.
The structure underlying a semi-Cartan graph encodes the combinatorial data necessary to define the Weyl groupoid $\mathcal{W}(\mathcal{G})$, which parametrizes the action of reflections on the set of real roots. Given a semi-Cartan graph $\mathcal{G}=(\mathbb{I},\mathcal{X},r,(A^{X})_{X\in\mathcal{X}})$, the Weyl groupoid of $\mathcal{G}$ is defined as follows.
Definition 1.2.
We denote by $\mathcal{D}(\mathcal{X},\operatorname{End}(\mathbb{Z}^{\mathbb{I}}))$ to be the category with objects $\operatorname{Ob}\mathcal{D}(\mathcal{X},\operatorname{End}(\mathbb{Z}^{\mathbb{I}}))=\mathcal{X}$, and morphisms
$\operatorname{Hom}(X,Y)=\{(Y,f,X)\mid f\in\operatorname{End}(\mathbb{Z}^{\mathbb{I}})\},$
where the composition of morphisms is defined by
$(Z,g,Y)\circ(Y,f,X)=(Z,gf,X),\ \text{for all}\ X,Y,Z\in\mathcal{X},\ f,g\in\operatorname{End}(\mathbb{Z}^{\mathbb{I}}).$
Let $\alpha_{i}$, $1\leq i\leq\theta$ be the standard basis of $\mathbb{Z}^{\mathbb{I}}$, and
$s_{i}^{X}\in\operatorname{Aut}(\mathbb{Z}^{\mathbb{I}}),\ s_{i}^{X}(\alpha_{j})=\alpha_{j}-a_{ij}^{X}\alpha_{i},\text{for all}\ j.$
We call the smallest subcategory of $\mathcal{D}(\mathcal{X},\operatorname{End}(\mathbb{Z}^{\mathbb{I}}))$ which contains all morphisms $(r_{i}(X),s_{i}^{X},X)$ with $i\in\mathbb{I}$, $X\in\mathcal{X}$ the Weyl groupoid of $\mathcal{G}$, denoted by $\mathcal{W}(\mathcal{G})$.
In the context of Nichols algebras, the Weyl groupoid captures the symmetries and reflection properties of the underlying braided vector space, establishing a bridge between the algebraic structure of Nichols algebras and the combinatorics of root systems [AHS10].
The interplay between Weyl groupoid theory and reflection theory has led to spectacular progress in the classification of finite-dimensional Nichols algebras. The classification program, initiated by Andruskiewitsch and Schneider in a series of papers [AS98,AS00,AS02,AS10], aims to determine all finite-dimensional pointed Hopf algebras over algebraically closed fields of characteristic zero whose group of group-like elements is abelian. A cornerstone of this classification is Heckenberger’s complete classification of finite-dimensional Nichols algebras of diagonal type [Hec09]. Building on this classification, Angiono made a breakthrough contribution by providing explicit presentations by generators and relations for all Nichols algebras of diagonal type [A13, A15]. The lifting problem of Nichols algebra of diagonal type has been addressed through cocycle deformation [AA+14, AG19, AA17]. The classification of finite-dimensional Nichols algebras over non-abelian groups has achieved excellent progress as well [AGn99,AGn03,AHS10,MS00,FK99,Gn00,Gn+11,HLV12,HS10,HS10b,HS10c,HS14,Hs15,HV17].
In this paper, we extend the reflection theory to the setting of some special coquasi-Hopf algebras. Our motivation comes from the classification of coradically graded pointed finite-dimensional coquasi-Hopf algebras, which has been studied in [rank1,coquasitriangular,tame-rep,FQQR2,QQG,huang2024classification,LiLiu1]. In the classification of pointed Hopf algebras over abelian groups, Nichols algebras of non-diagonal type do not appear. However, in the coquasi-Hopf algebra setting, Nichols algebras of non-diagonal type do occur and present greater challenges. The authors in huang2024classification used extensive computations to establish that a certain class of such Nichols algebras is infinite-dimensional. We later provided an alternative proof of their main result via categorical Morita equivalence LiLiu1. Nonetheless, these arguments are not entirely natural, and we hope that the theory of semi-Cartan graphs can be extended to coquasi-Hopf algebras. If such a generalization holds, it could be applied to the classification of finite-dimensional coquasi-Hopf algebras over non-abelian groups.
Our approach combines the methods of Simon and HSdualpair together. The next theorem generalizes [HSdualpair, Theorem 7.1] to the case of coquasi-Hopf algebras with bijective antipodes under some minor conditions, which is crucial to description of reflection of Nichols algebras.
Theorem A.
(Theorem 3.8) Let $H$ be a coquasi-Hopf algebra with bijective antipode, and $\mathcal{C}={{}^{H}_{H}\mathcal{YD}}$, (see Definition 2.5 for details). Suppose $N\in{{}^{H}_{H}\mathcal{YD}}$, and $\mathcal{B}(N)$ is finite-dimensional. We have the following equivalence of braided monoidal categories
$\Omega:{{}^{\mathcal{B}(N)}_{\mathcal{B}(N)}\mathcal{YD}(\mathcal{C})}$ $\cong{{}^{\mathcal{B}(N^{*})}_{\mathcal{B}(N^{*})}\mathcal{YD}(\mathcal{C})}.$
It is worth mentioning that in our cases, the above two categories are not strict. This makes the definition of the functor $\Omega$ more complicated and adds some challenges to studying its properties.
Now we restrict to the case of pointed cosemisimple coquasi-Hopf algebras. Let $G$ be a finite group, and $\Phi$ be a $3$-cocycle on $G$, which gives rise to the Yetter-Drinfeld module category $^{G}_{G}\mathcal{YD}^{\Phi}$} QQG. Let $\theta\geq 1$ be a positive integer, $\mathbb{I}=\{1,2,..,\theta\}$ and $M=(M_{1},\ldots,M_{\theta})$, where $M_{1},\ldots,M_{\theta}\in{{}^{G}_{G}\mathcal{YD}^{\Phi}}$ are irreducible Yetter-Drinfeld modules. We always denote $\mathcal{B}(M)$ to be the Nichols algebra of $M_{1}\oplus\cdots\oplus M_{\theta}$ in the following context. Suppose $M$ admits $i$-th reflection for some $1\leq i\leq\theta$ (see Definition 5.1). Then the $i$-th reflection of $M$ is denoted by $R_{i}(M)=(V_{1},\ldots,V_{\theta})$, which is a tuple consisting of finite-dimensional irreducible Yetter-Drinfeld modules. There is a natural explanation of reflection of Nichols algebra in terms of $\Omega$:
Theorem B.
(Theorem 5.7) With above assumptions on $M$. Suppose $M$ admits the $i$-th reflection for $1\leq i\leq\theta$, and $\mathcal{B}(M_{i})$ is finite-dimensional. Then there is an isomorphism of Hopf algebras in $^{G}_{G}\mathcal{YD}^{\Phi}$}:
$\Theta_{i}:\mathcal{B}(R_{i}(M))\cong\Omega_{i}\left(\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}\right)\#\mathcal{B}(M_{i}^{*}).$ (1.1)
Readers who are familiar with reflection theory in the Hopf algebra context will recognize this formula, we emphasize that our contribution lies in establishing this result in the framework of coquasi-Hopf algebras. The extension from Hopf algebras to coquasi-Hopf algebras introduces technical challenges, particularly in sections 3 and 4 where the non-strict monoidal structure necessitates a reworking of the reflection theory. Now with the help of above theorems, our main result follows naturally.
Theorem C.
(Theorem 5.9) Suppose $M$ admits all reflections and let $\mathcal{X}=\{[P]\mid P\in\mathcal{F}_{\theta}(M)\}$, see Definition 5.2 for related notations. We denote the generalized Cartan matrix $A^{[X]}=(a_{ij}^{X})_{i,j\in\mathbb{I}}$ for each $P]\in\mathcal{X}$, where $(a_{ij}^{X})_{i,j\in\mathbb{I}}$ is defined in Lemma 5.4. Let $r$ be the following map,
$r:\mathbb{I}\times\mathcal{X}\rightarrow\mathcal{X},\ i\times X\mapsto[R_{i}(X)].$
Furthermore, we assume $\mathcal{B}(P_{i})$ is finite-dimensional for each $P\in\mathcal{X}$ and $1\leq i\leq\theta$, then
$\mathcal{G}(M)=(\mathbb{I},\mathcal{X},r,(A^{X})_{X\in\mathcal{X}}),$
is a semi-Cartan graph.
After establishing that $M$ gives rise to a semi-Cartan graph, several additional criteria for determining the finite-dimensionality of the Nichols algebra can also be derived. We show that if $\mathcal{B}(M)$ is finite-dimensional, then the semi-Cartan graph $\mathcal{G}(M)$ is a finite semi-Cartan graph. Furthermore, if $\mathcal{G}(M)$ is moreover a standard finite semi-Cartan graph, that is $A^{X}=A^{Y}$, for all $X,Y\in\mathcal{X}$, then $A^{M}$ must be a finite Cartan matrix. As an application of these results, we provide a third proof of [huang2024classification, Proposition 4.1].
The paper is organized as follows. Section 2 is dedicated to preliminaries. In Section 3, we establish a braided monoidal equivalence between categories of Yetter-Drinfeld modules. Section 4 deals with projections of Nichols algebras; the results obtained here are essential to the study of reflections of Nichols algebras. In Section 5, we prove that, under minor assumptions, a tuple of simple Yetter-Drinfeld modules gives rise to a semi-Cartan graph. Finally, in Section 6, we apply the theory of semi-Cartan graph to the classification of finite-dimensional Nichols algebras over certain coquasi-Hopf algebra.
2 Preliminaries
Throughout this paper, $\mathbbm{k}$ is an algebraically closed field of characteristic zero and all linear spaces are over $\mathbbm{k}$.
2.1 Coquasi-Hopf algebras
By definition, coquasi-Hopf algebras are exactly the dual of Drinfeld’s quasi-Hopf algebras Drinfeld. Their formal definition can be given as follows.
Definition 2.1.
A coquasi-Hopf algebra is a coalgebra $(H,\Delta,\varepsilon)$ equipped with a compatible quasi-algebra structure and an antipode $(\mathcal{S},\alpha,\beta)$. Namely, there exist:
  • Two coalgebra homomorphisms:
    $m:H\otimes H\rightarrow H,\qquad a\otimes b\mapsto ab,$
    $\mu:\mathbbm{k}\rightarrow H,\qquad\lambda\mapsto\lambda 1_{H},$
  • A convolution-invertible map $\Phi:H^{\otimes 3}\rightarrow\mathbbm{k}$ called associator,
  • A coalgebra antimorphism $\mathcal{S}:H\rightarrow H$,
  • Two linear functions $\alpha,\beta:H\rightarrow\mathbbm{k}$
such that for all $a,b,c,d\in H$ the following equalities hold:
$a_{1}(b_{1}c_{1})\Phi(a_{2},b_{2},c_{2})$ $=\Phi(a_{1},b_{1},c_{1})(a_{2}b_{2})c_{2},$ (2.1)
$1_{H}a$ $=a=a1_{H},$ (2.2)
$\Phi(a_{1},b_{1},c_{1}d_{1})\Phi(a_{2}b_{2},c_{2},d_{2})$ $=\Phi(b_{1},c_{1},d_{1})\Phi(a_{1},b_{2}c_{2},d_{2})\Phi(a_{2},b_{3},c_{3}),$ (2.3)
$\Phi(a,1_{H},b)$ $=\varepsilon(a)\varepsilon(b),$ (2.4)
$\mathcal{S}(a_{1})\alpha(a_{2})a_{3}$ $=\alpha(a)1_{H},\quad a_{1}\beta(a_{2})\mathcal{S}(a_{3})=\beta(a)1_{H},$ (2.5)
$\Phi(a_{1},\mathcal{S}(a_{3}),a_{5})\beta(a_{2})\alpha(a_{4})$ $=\Phi^{-1}(\mathcal{S}(a_{1}),a_{3},\mathcal{S}(a_{5}))\alpha(a_{2})\beta(a_{4})=\varepsilon(a).$ (2.6)
Throughout this paper, we use the Sweedler sigma notation $\Delta(a)=a_{1}\otimes a_{2}$ for the coproduct and $a_{1}\otimes a_{2}\otimes\cdots\otimes a_{n+1}$ for the result of the $n$-iterated application of $\Delta$ on $a$. We say $H$ has a bijective antipode if $\mathcal{S}$ is bijective.
A coquasi-Hopf algebra $H$ is pointed if the underlying coalgebra is so. Given a coquasi-Hopf algebra, let $\{H_{n}\}_{n\geq 0}$ be its coradical filtration, and
$\operatorname{gr}(H):=H_{0}\oplus H_{1}/H_{0}\oplus H_{2}/H_{1}\oplus\cdots$
the corresponding coradically graded coalgebra. Then naturally $(\operatorname{gr}(H),\operatorname{gr}(\Phi))$ inherits from $H$ a graded coquasi-Hopf algebra structure. Furthermore, for homogeneous elements $a,b,c\in\operatorname{gr}(H)$, $\operatorname{gr}(\Phi)(a,b,c)=0$ unless they all lie in $H_{0}$.
Following the definition in [preantipode], we now recall the definition of a preantipode, which plays an important role in further calculations.
Definition 2.2.
[preantipode, Definition 3.6] Let $(H,\Phi)$ be a coquasi-bialgebra, a preantipode for $H$ is a $\mathbbm{k}$-linear map $\mathbb{S}:H\rightarrow H$ such that, for all $h\in H$,
$\mathbb{S}(h_{1})_{1}h_{2}\otimes\mathbb{S}(h_{1})_{2}$ $=1_{H}\otimes\mathbb{S}(h),$
$\mathbb{S}(h_{2})_{1}\otimes h_{1}\mathbb{S}(h_{2})_{2}$ $=\mathbb{S}(h)\otimes 1_{H},$
$\Phi(h_{1},\mathbb{S}(h_{2}),h_{3})$ $=\varepsilon(h).$
Furthermore, [preantipode, Remark 3.7] provides us a useful property:
$h_{1}\mathbb{S}(h_{2})=\varepsilon(\mathbb{S}(h))1_{H}=\mathbb{S}(h_{1})h_{2},\ \text{for all}\ h\in H.$ (2.7)
Meanwhile, when $H$ is equipped with the structure of a coquasi-Hopf algebra with antipode $(\mathcal{S},\alpha,\beta)$, then preantipode for $H$ always exists.
Lemma 2.3.
[preantipode, Theorem 3.10] Let $H$ be a coquasi-Hopf algebra with antipode $(\mathcal{S},\alpha,\beta)$, then
$\mathbb{S}:=\beta*\mathcal{S}*\alpha$ (2.8)
is a preantipode for $H$, where * denotes convolution product.
Example 2.4.
A baby example of coquasi-Hopf algebra is $(\mathbbm{k}G,\Phi)$, where $G$ is a group and $\Phi$ a normalized 3-cocycle on $G$. It is well known that the group algebra $\mathbbm{k}G$ is a Hopf algebra with:
$\Delta(g)=g\otimes g,\qquad\mathcal{S}(g)=g^{-1},\qquad\varepsilon(g)=1\qquad\text{for any }g\in G.$
By extending $\Phi$ linearly, $\Phi:(\mathbbm{k}G)^{\otimes 3}\rightarrow\mathbbm{k}$ becomes a convolution-invertible map. Define two linear functions $\alpha,\beta:\mathbbm{k}G\rightarrow\mathbbm{k}$ by:
$\alpha(g):=\varepsilon(g),\qquad\beta(g):=\frac{1}{\Phi(g,g^{-1},g)}$
for any $g\in G$. Then $\mathbbm{k}G$ together with these $\Phi,\alpha$ and $\beta$ becomes a coquasi-Hopf algebra. We denote this resulting coquasi-Hopf algebra by $(\mathbbm{k}G,\Phi)$. For all $g\in G$, we define
$\mathbb{S}(g)=\frac{1}{\Phi(g,g^{-1},g)}g^{-1}.$
Then Lemma 2.3 shows $\mathbb{S}$ is a preantipode for $(\mathbbm{k}G,\Phi)$.
2.2 Yetter-Drinfeld module categories over coquasi-Hopf algebras
We now turn our attention to the Yetter-Drinfeld module category structure. The definition of Yetter-Drinfeld module category $^{H}_{H}\mathcal{YD}$} over an arbitrary coquasi Hopf-algebra $H$ was already given in [YD-module]. It is defined as the center $\mathcal{Z}({{{}^{H}}\mathscr{M}})$ of the comodule category $^{H}\mathscr{M}$}. Hence it is braided monoidal equivalent to $\textbf{Rep}(D(H))$. Since it plays a crucial role in our paper, we recall this definition below.
Definition 2.5.
[YD-module, Definition 3.1] Let $H$ be a coquasi-Hopf algebra with associator $\Phi$. A left-left Yetter–Drinfeld module over $H$ is a triple $(V,\delta_{V},\rhd)$ such that:
  • $(V,\delta_{V})$ is a left comodule of $H$ and we denote $\delta_{V}(v)$ by $v_{-1}\otimes v_{0}$ as usual;
  • $\rhd:H\otimes V\to V$ is a $\mathbbm{k}$-linear map satisfying for all $h,l\in H$ and $v\in V$:
    $(hl)\rhd v=\frac{\Phi(h_{2},(l_{2}\rhd v_{0})_{-1},l_{3})}{\Phi(h_{1},l_{1},v_{-1})\Phi((h_{3}\rhd(l_{2}\rhd v_{0})_{0})_{-1},h_{4},l_{4})}(h_{3}\rhd(l_{2}\rhd v_{0})_{0})_{0},$ (2.9)
    $1_{H}\rhd v=v,$ (2.10)
    $(h_{1}\rhd v)_{-1}h_{2}\otimes(h_{1}\rhd v)_{0}=h_{1}v_{-1}\otimes(h_{2}\rhd v_{0}).$ (2.11)
A morphism $f:(V,\delta_{V},\rhd)\rightarrow(V^{\prime},\delta_{V^{\prime}},\rhd^{\prime})$ is a colinear map $f:(V,\delta_{V})\rightarrow(V^{\prime},\delta_{V^{\prime}})$ such that $f(h\rhd v)=h\rhd^{\prime}f(v)$ for all $h\in H$.
The category $^{H}_{H}\mathcal{YD}$} is a $\mathbbm{k}$-linear braided monoidal abelian category over the field $\mathbbm{k}$. The unit object of $^{H}_{H}\mathcal{YD}$} is $\mathbbm{k}$, which is regarded as an object in $^{H}_{H}\mathcal{YD}$} via trivial structures. For $V,W\in{{}^{H}_{H}\mathcal{YD}}$, the tensor product of Yetter-Drinfeld modules is defined by:
$(V,\delta_{V},\rhd)\otimes(W,\delta_{W},\rhd)=(V\otimes W,\delta_{V\otimes W},\rhd),$
where $\delta_{V\otimes W}$ is given by
$\delta_{V\otimes W}(v\otimes w)=v_{-1}w_{-1}\otimes v_{0}\otimes w_{0},$
and
$h\rhd(v\otimes w)=\frac{\Phi(h_{1},v_{-1},w_{-2})\Phi((h_{2}\rhd v_{0})_{-1},(h_{4}\rhd w_{0})_{-1},h_{5})}{\Phi((h_{2}\rhd v_{0})_{-2},h_{3},w_{-1})}(h_{2}\rhd v_{0})_{0}\otimes(h_{4}\rhd w_{0})_{0}.$ (2.12)
The braiding $c_{V,W}:V\otimes W\rightarrow W\otimes V$ is given by:
$c_{V,W}(v\otimes w)=(v_{-1}\rhd w)\otimes v_{0}.$ (2.13)
[…]
$p(h,g)$ $=\Phi^{-1}(h,g_{1}\beta(g_{2}),\mathcal{S}(g_{3})),$
$q(h,g)$ $=\Phi(h,g_{3},\alpha(\mathcal{S}^{-1}(g_{2}))\mathcal{S}^{-1}(g_{1})),$
$s(h,g)$ $=\Phi(\mathcal{S}^{-1}(h_{3})\beta(\mathcal{S}^{-1}(h_{2})),h_{1},g),$
$t(h,g)$ $=\Phi^{-1}(\mathcal{S}(h_{1})\alpha(h_{2}),h_{3},g).$
for any $h,g\in H$. These maps satisfy the following properties:
$(h_{1}g_{1})\mathcal{S}(g_{3})p(h_{2},g_{2})$ $=h_{2}p(h_{1},g),$ (2.14)
$(h_{2}g_{3})\mathcal{S}^{-1}(g_{1})q(h_{1},g_{2})$ $=h_{1}q(h_{2},g),$ (2.15)
$q(h_{1}g_{1},\mathcal{S}(g_{3}))p(h_{2},g_{2})$ $=\varepsilon(h)\varepsilon(g),$ (2.16)
$\mathcal{S}(h_{1})(h_{3}g_{2})t(h_{2},g_{1})$ $=g_{1}t(h,g_{2}),$ (2.17)
$s(\mathcal{S}(h_{1}),h_{3}g_{2})t(h_{2},g_{1})$ $=\varepsilon(h)\varepsilon(g).$ (2.18)
With the above notations established, we can now prove the following lemma.
Lemma 2.6.
Suppose $H$ has a bijective antipode.
(1) The equation (2.11) is equivalent to
$(h\rhd v)_{-1}\otimes(h\rhd v)_{0}=p((h_{3}\rhd v_{0})_{-1},h_{4})q(h_{1}v_{-2},\mathcal{S}(h_{6}))(h_{2}v_{-1})\mathcal{S}(h_{5})\otimes(h_{3}\rhd v_{0})_{0}.$ (2.19)
(2) The equation (2.11) implies:
$v_{-1}\otimes(h\rhd v_{0})=t(h_{3},v_{-1})s(\mathcal{S}(h_{1}),(h_{4}\rhd v_{0})_{-1}h_{6})\mathcal{S}(h_{2})((h_{4}\rhd v_{0})_{-2}h_{5})\otimes(h_{4}\rhd v_{0})_{0}.$ (2.20)
Proof.
(1): Suppose $V\in{{}^{H}_{H}\mathcal{YD}}$ with a linear map $\rhd:H\otimes V\rightarrow V$ satisfying (2.11). For any $v\in V,h\in H$, then we have
$(h\rhd v)_{-1}\otimes(h\rhd v)_{0}=\varepsilon(h_{2})\varepsilon((h_{1}\rhd v)_{-2})(h_{1}\rhd v)_{-1}\otimes(h_{1}\rhd v)_{0}$
[…]
Conversely, if (2.19) holds, we have
$\varepsilon((h\rhd v)_{-1})(h\rhd v)_{0}=p((h_{2}\rhd v_{0})_{-1},h_{3})q(h_{1}v_{-1},\mathcal{S}(h_{4}))(h_{2}\rhd v_{0})_{0}.$ (2.21)
The origin equation comes from the following calculation,
$(h_{1}\rhd v)_{-1}h_{2}\otimes(h_{1}\rhd v)_{0}$ $=p((h_{3}\rhd v_{0})_{-1},h_{4})q(h_{1}v_{-2},\mathcal{S}(h_{6}))((h_{2}v_{-1})\mathcal{S}(h_{5}))h_{7}\otimes(h_{3}\rhd v_{0})_{0}$
[…]
(2) We have
$v_{-1}\otimes(h\rhd v_{0})$ $=\varepsilon(v_{-1})\varepsilon(h_{1})v_{-2}\otimes(h_{2}\rhd v_{0})$
$\overset{(\ref{2.18})}{=}s(\mathcal{S}(h_{1}),h_{3}v_{-1})t(h_{2},v_{-2})v_{-3}\otimes(h_{4}\rhd v_{0})$
$\overset{(\ref{2.17})}{=}s(\mathcal{S}(h_{1}),h_{5}v_{-1})t(h_{3},v_{-3})\mathcal{S}(h_{2})(h_{4}v_{-2})\otimes h_{6}\rhd v_{0}$
[…]
Example 2.7.
For our purpose, we aim to describe Yetter-Drinfeld modules over coquasi-Hopf algebra of the form $(\mathbbm{k}G,\Phi)$, where $G$ is a finite group and $\Phi$ a $3$-cocycle on $G$. Assume that $V$ is a left $kG$-comodule with comodule structure map $\delta_{L}:V\rightarrow kG\otimes V$. Define
$^{g}V:=\{v\in V\mid\delta_{L}(v)=g\otimes v\}.$
Thus
$V=\bigoplus_{g\in G}{}^{g}V.$
Here we call $g$ the degree of the elements in $^{g}V$ and denote $\deg v=g$ for $v\in{}^{g}V$.
The left $\mathbbm{k}G$-comodule $(V,\delta_{L})$ is a left-left Yetter–Drinfeld module over the coquasi-Hopf algebra $H=(\mathbbm{k}G,\Phi)$ if there is a linear map $\rhd:G\otimes V\rightarrow V$ such that for all $e,f\in G$ and $v\in{}^{g}V$:
$e\rhd(f\rhd v)=\frac{\Phi(e,f,g)\Phi(efgf^{-1}e^{-1},e,f)}{\Phi(e,fgf^{-1},f)}(ef)\rhd v,$ (2.22)
$1_{H}\rhd v=v,$ (2.23)
$e\rhd v\in{}^{ege^{-1}}V.$ (2.24)
The category of all left-left Yetter–Drinfeld modules over $(\mathbbm{k}G,\Phi)$ is denoted by $^{G}_{G}\mathcal{YD}^{\Phi}$}. Similarly, one can define left-right, right-left and right-right Yetter–Drinfeld modules over $(\mathbbm{k}G,\Phi)$.
It is well-known that $^{G}_{G}\mathcal{YD}^{\Phi}$} is a braided tensor category. More precisely, for any $M,N\in{{}^{G}_{G}\mathcal{YD}^{\Phi}}$, the structure maps of $M\otimes N$ as a left-left Yetter–Drinfeld module are given by:
$\delta_{L}(m_{g}\otimes n_{h})$ $:=gh\otimes m_{g}\otimes n_{h},$ (2.25)
$x\rhd(m_{g}\otimes n_{h})$ $:=\frac{\Phi(x,g,h)\Phi(xgx^{-1},xhx^{-1},x)}{\Phi(xgx^{-1},x,h)}x\rhd m_{g}\otimes x\rhd n_{h},$ (2.26)
for all $x,g,h\in G$ and $m_{g}\in{}^{g}M$, $n_{h}\in{}^{h}N$.
The associativity constraint $a$ and the braiding $c$ of $^{G}_{G}\mathcal{YD}^{\Phi}$} are given respectively by:
$a((u_{e}\otimes v_{f})\otimes w_{g})$ $=\Phi(e,f,g)^{-1}u_{e}\otimes(v_{f}\otimes w_{g}),$ (2.27)
$c(u_{e}\otimes v_{f})$ $=e\rhd v_{f}\otimes u_{e},$ (2.28)
for all $e,f,g\in G$, $u_{e}\in{}^{e}U$, $v_{f}\in{}^{f}V$, $w_{g}\in{}^{g}W$ and $U,V,W\in{{}^{G}_{G}\mathcal{YD}^{\Phi}}$.
In general, let $\mathcal{C}$ be a braided monoidal category and $A$ a Hopf algebra in $\mathcal{C}$. The category of Yetter-Drinfeld modules $^{A}_{A}\mathcal{YD}(\mathcal{C})$} is well-defined. The reader may refer to [rootsys, Section 3] for details.
2.3 Bosonization for coquasi-Hopf algebras
Bosonization for Hopf algebras is well-known. In [quasibon], the authors introduced bosonization for quasi-Hopf algebras. Dually, bosonization for coquasi-Hopf algebras is given in YD-module. Let $H$ be a coquasi-Hopf algebra with associator $\Phi$. Suppose $R$ is a Hopf algebra in $^{H}_{H}\mathcal{YD}$}, we denote
$r^{1}\otimes r^{2}:=\Delta_{R}(r).$
Lemma 2.8.
[YD-module, Theorem 5.2] Let us consider on $N:=R\otimes H$ the following structures:
$(r\otimes h)(s\otimes k)$ $=\frac{\Phi(h_{2},s_{-1},k_{2})\Phi(r_{-1},(h_{3}\rhd s_{0})_{-1},h_{5}k_{4})}{\Phi(r_{-2},h_{1},s_{-2}k_{1})\Phi((h_{3}\rhd s_{0})_{-2},h_{4},k_{3})}r_{0}(h_{3}\rhd s_{0})_{0}\otimes h_{6}k_{5},$
$u_{B}(k)$ $=k1_{R}\otimes 1_{M},$
$\Delta_{B}(r\otimes h)$ $=\Phi^{-1}(r_{-1}^{1},r_{-2}^{2},h_{1})r_{0}^{1}\otimes r_{-1}^{2}h_{2}\otimes r_{0}^{2}\otimes h_{3},$
$\varepsilon_{B}(r\otimes h)$ $=\varepsilon_{R}(r)\varepsilon_{M}(h),$
$\omega_{B}((r\otimes h),(s\otimes k),(t\otimes l))$ $=\varepsilon_{R}(r)\varepsilon_{R}(s)\varepsilon_{R}(t)\Phi(h,k,l),$
where $r,s,t\in R$, $h,k,l\in H$. With above operations, $N$ is a coquasi-Hopf algebra.
Now suppose $N$ are $H$ are both coquasi-Hopf algebras. Furthermore, assume there exist morphisms of coquasi-Hopf algebras
$\pi:N\rightarrow H\ \text{ and }\sigma:H\rightarrow N$
such that $\pi\sigma=\operatorname{id}_{N}$.
Lemma 2.9.
[YD-module, Theorem 5.8] Under above assumptions, $L:=N^{\operatorname{co}H}$ is a Hopf algebra in $^{H}_{H}\mathcal{YD}$}. For all $a\in N$, we denote
$\tau(a)=\Phi_{N}(a_{1},\sigma\mathbb{S}\pi(a_{3})_{1},a_{4})a_{2}\sigma\mathbb{S}\pi(a_{3})_{2}.$ (2.29)
Then $\tau:N\rightarrow L$ is a well-defined map, where, for all $r,s\in L$, $h\in H$, $k\in\mathbbm{k}$, the Yetter-Drinfeld module structure is given by
$\operatorname{ad}(h)(r)$ $:=\tau(\sigma(h)r)=\Phi_{H}(h_{1}r_{-1},\mathbb{S}(h_{3})_{1},h_{4})(h_{2}r_{0})\mathbb{S}(h_{3})_{2},$
$r_{-1}\otimes r_{0}$ $:=\rho_{R}(r):=\pi(r_{1})\otimes r_{2}.$
Moreover, $L$ is a Hopf algebra in $^{H}_{H}\mathcal{YD}$} via
$m_{L}(r\otimes s):=rs,\ \ \ \ \ \ \ u_{L}(k)=k1_{N}$
$\Delta_{L}(r):=\tau(r_{1})\otimes\tau(r_{2}),\ \ \ \ \varepsilon_{L}(r):=\varepsilon_{N}(r).$
Furthermore, there is an isomorphism of coquasi-Hopf algebra $\phi:L\#H\rightarrow N$ given by
$\phi(r\otimes h)=r\sigma(h),\ \ \phi^{-1}(a)=\tau(a_{1})\otimes\pi(a_{2}).$
For reader’s convenience, we still give an example about $^{G}_{G}\mathcal{YD}^{\Phi}$}, where $G$ is a finite group and $\Phi$ is a $3$-cocycle on $G$. Suppose $R$ is a Hopf algebra in $^{G}_{G}\mathcal{YD}^{\Phi}$}. Since $R$ is a left $G$-comodule, there is a $G$-grading on $R$:
$R=\bigoplus_{x\in G}{{}^{x}R},$
where $^{x}R:=\{X\in R\mid\rho(x)=x\otimes X\}$. For simplicity, in this paper, homogeneous elements in $^{G}_{G}\mathcal{YD}^{\Phi}$} are denoted by capital letters, say $X,Y,Z$..., and the associated degrees are denoted by their lower cases, say $x,y,z$
Example 2.10.
Let us maintain the assumptions on $R$ from above. For all $X,Y,Z\in R$, $h,g,k\in G$, we consider the following structure on $M=R\otimes\mathbbm{k}G$.
$(X\otimes h)(Y\otimes g)=\frac{\Phi(h,y,k)\Phi(x,hyh^{-1},hg)}{\Phi(x,h,yg)\Phi(hyh^{-1},h,g)}X(h\rhd Y)\otimes hg,\ \ \ u_{H}(k)=k1_{R}\otimes 1_{\mathbbm{k}G},$
$\Delta(X\otimes h)=\Phi^{-1}(x^{1},x^{2},h)X^{1}\otimes x^{2}h\otimes X^{2}\otimes h,\ \ \ \ \ \ \ \varepsilon(X\otimes h)=\varepsilon_{R}(X)\varepsilon_{\mathbbm{k}G}(h),$
$\Phi_{M}(X\otimes h,Y\otimes g,Z\otimes k)=\varepsilon_{R}(X)\varepsilon_{R}(Y)\varepsilon_{R}(Z)\Phi(h,g,k),$
$\mathcal{S}(X\otimes g)=\Phi(g^{-1}x^{-1},x,g)(1\otimes g^{-1}x^{-1})(\mathcal{S}_{R}(X)\otimes 1),$
$\alpha(1\otimes g)=1,\ \ \alpha(X\otimes g)=0,\ \ \ \ \beta(1\otimes g)=\frac{1}{\Phi(g,g^{-1},g)},\ \ \beta(X\otimes g)=0.$
With the above operations, $M$ is indeed a coquasi-Hopf algebra.
Conversely, let us now consider the reverse construction. Suppose $M$ is a coradically graded coquasi-Hopf algebra with coradical $(\mathbbm{k}G,\Phi)$. Then $R:=M^{\operatorname{co}\mathbbm{k}G}$ is an object in $^{G}_{G}\mathcal{YD}^{\Phi}$} equipped with the following structures:
$\operatorname{ad}(g)(X)=\frac{\Phi(gh,g^{-1},g)}{\Phi(g,g^{-1},g)}(gX)g^{-1},\ \$
$\rho_{R}(X)=\pi(X_{1})\otimes X_{2},\ X\in{{}^{h}R},g\in G.$
Furthermore, $R$ carries a Hopf algebra structure in $^{G}_{G}\mathcal{YD}^{\Phi}$},
$m_{R}(X\otimes Y)=m_{M}(X\otimes Y),\ \ \ \ u_{R}(k)=\mathbbm{k}1_{M},$
$\Delta_{R}(X)=\frac{1}{\Phi(x_{1},x_{2},x_{2}^{-1})}X_{1}\cdot x_{2}^{-1}\otimes X_{2},\ \ \ \varepsilon_{R}(X)=\varepsilon_{M}(X),$
$\mathcal{S}_{R}(X)=\frac{1}{\Phi(x,x^{-1},x)}x\cdot\mathcal{S}(X),$
where $X,Y\in R$.
However, this construction alone is not sufficient for our purposes. The following condition occurs frequently in this paper. Given $R,R^{\prime}$ be two Hopf algebras in $^{G}_{G}\mathcal{YD}^{\Phi}$} together with a surjective Hopf algebra map $\pi^{\prime}:R\rightarrow R^{\prime}$ and an injective map $i^{\prime}:R^{\prime}\rightarrow R$ such that $\pi\circ i=\operatorname{id}_{R}^{\prime}$. Let $M=R\#\mathbbm{k}G$, and $H=R^{\prime}\#\mathbbm{k}G$. Obviously, these maps naturally extend to yield a surjective coquasi-Hopf algebra map $\pi:M\rightarrow H$ and an injective coquasi-Hopf algebra map $i:H\rightarrow M$, such that $\pi\circ i=\operatorname{id}$.
Example 2.11.
Let $L:=M^{\operatorname{co}H}$, then $L$ is an object in $^{H}_{H}\mathcal{YD}$}, with the following structures:
$\operatorname{ad}:H\otimes L\rightarrow L,h\otimes l\mapsto\Phi_{H}(h_{1}l_{-1},\mathbb{S}_{H}(h_{3})_{1},h_{4})(h_{2}l_{0})\mathbb{S}_{H}(h_{3})_{2},$
$\operatorname{\rho}:L\rightarrow H\otimes L,l\mapsto\pi(l_{1})\otimes l_{2}.$
Our primary focus now shifts to analyzing the linear map $\operatorname{ad}$. Given $g\in G$ and $X\in{{}^{x}R}$,
$\operatorname{ad}(g)(X)=\frac{\Phi(gx,g^{-1},g)}{\Phi(g,g^{-1},g)}(gX)g^{-1}=g\rhd X.$
Now suppose $Y$ is a $(y,1)$-primitive element in $H$, we first need to compute what $\mathbb{S}(Y)$ is,
$\mathbb{S}(Y)=\beta*\mathcal{S}*\alpha(Y)=\frac{1}{\Phi(y,y^{-1},y)}\mathcal{S}(Y)$ $=\frac{1}{\Phi(y,y^{-1},y)}(1\#y^{-1})(-Y\#1)$
$=-\frac{1}{\Phi(y,y^{-1},y)}y^{-1}Y.$
Then we can proceed to calculate $\operatorname{ad}(Y)(X)$,
$\operatorname{ad}(Y)(X)=YX-\frac{1}{\Phi(y,y^{-1},y)}(yX)(y^{-1}Y)$ $=YX-\frac{\Phi(yx,y^{-1},y)}{\Phi(y,y^{-1},y)}((yX)y^{-1})Y$ (2.30)
$=YX-\operatorname{ad}(y)(X)Y.$
The definition of the bosonization naturally extends to the case that $R$ is a Hopf algebra in the braided monoidal category $^{A}_{A}\mathcal{YD}(\mathcal{C})$}, where $\mathcal{C}$ is now an arbitrary braided monoidal category. We collect several key results from B95 and AF00 that will be essential for the construction presented in Section 3.
Lemma 2.12.
[AF00, Theorem 3.2] Let $H$ and $A$ be Hopf algebras in a braided monoidal category $\mathcal{C}$. Let $\pi:H\to A$ and $\iota:A\to H$ be Hopf algebra morphisms such that $\pi\circ\iota=\mathrm{id}_{A}$. If $\mathcal{C}$ has equalizers and $A\otimes(-)$ preserves equalizers, there is a Hopf algebra $R$ in the braided monoidal category $^{A}_{A}\mathcal{YD}(\mathcal{C})$}, such that
$H\cong R\#A.$
Lemma 2.13.
[B95, Proposition 4.2.3] Let $A$ be a Hopf algebra in $\mathcal{C}$ and $K$ a Hopf algebra in $^{A}_{A}\mathcal{YD}(\mathcal{C})$}. There is an obvious isomorphism of braided monoidal categories
${{}^{K\#A}_{K\#A}\mathcal{YD}}(\mathcal{C})\cong{{}^{K}_{K}\mathcal{YD}}({{}^{A}_{A}\mathcal{YD}(\mathcal{C})}).$ (2.31)
2.4 Nichols algebras in braided monoidal category
Having established the necessary framework for bosonization, we now turn to the theory of Nichols algebras. Roughly speaking, Nichols algebras serve as the analogues of ordinary symmetric algebras in more general braided monoidal categories. This concept can be defined through various equivalent approaches. Now let $\mathcal{C}$ be an abelian braided monoidal category with braiding isomorphism $c$. We present the definition of Nichols algebras in $\mathcal{C}$.
Let $V$ be a nonzero object in $\mathcal{C}$, by $T(V)$ we denote the tensor algebra in $\mathcal{C}$ generated freely by $V$. Here we denote that
$V^{\otimes n}:=\left(\cdots\left(\left(V\otimes V\right)\otimes V\right)\cdots\otimes V\right).$
Then $T(V)$ is isomorphic to $\bigoplus_{n\geq 0}V^{\otimes n}$ as an object. It is standard that $T(V)$ is a graded Hopf algebra in $\mathcal{C}$.
Definition 2.14.
Let $V\in\mathcal{C}$ and $I(V)$ be the largest coideal of T(V){T}(V) contained in $\bigoplus_{n\geq 2}T^{n}(V)$. The Nichols algebra of $V$ is defined by
$\mathcal{B}(V):=T(V)/I(V).$
For further analysis, we shall assume that $\mathcal{C}$ is a $\mathbbm{k}$-linear braided monoidal abelian category over the field $\mathbbm{k}$. This additional structure allows us to adopt an equivalent definition which is more convenient for our purposes.
Definition 2.15.
[defofnichols, Definition 2.4] For a given object $V\in\mathcal{C}$ the Nichols algebra $\mathcal{B}(V)$ is the unique Hopf algebra in $\mathcal{C}$ that satisfies the following conditions:
$(1)$ The Hopf algebra $\mathcal{B}(V)$ is graded by the non-negative integers.
$(2)$ The zeroth component of the grading satisfies $\mathcal{B}(V)_{0}=\mathbbm{k}$.
$(3)$ The first component of the grading satisfies $\mathcal{B}(V)_{1}=V$ , and $\mathcal{B}(V)$ is generated by $V$ as an algebra in $\mathcal{C}$.
$(4)$ The subobject of primitive elements of $\mathcal{B}(V)$ is $V$.
Suppose $\mathcal{C}$ and $\mathcal{D}$ are braided monoidal abelian categories, and $F:\mathcal{C}\rightarrow\mathcal{D}$ is a braided monoidal equivalence of abelian category. The following lemma plays a crucial role in our theory.
Lemma 2.16.
Let $V\in\mathcal{C}$ be an object and $\mathcal{B}(V)$ be its corresponding Nichols algebra in $\mathcal{C}$. Then
$F(\mathcal{B}(V))\cong\mathcal{B}(F(V))$
as Hopf algebras in $\mathcal{D}$.
Proof.
Let $V\in\mathcal{C}$ and $J_{V,V}:F(V\otimes V)\cong F(V)\otimes F(V)$ the monoidal structure of $F$. Using $J$ repeatedly, we have
$F(V^{\otimes n})\cong F(V)^{\otimes n}$
for each $n\in\mathbb{N}$. Therefore we have $F(\bigoplus\limits_{n\in\mathbb{N}}V^{\otimes n})\cong\bigoplus\limits_{n\in\mathbb{N}}F(V^{\otimes n})\cong\bigoplus\limits_{n\in\mathbb{N}}F(V)^{\otimes n}$. That is $F(T(V))\cong T(F(V))$.
Recall that the Nichols algebra in $\mathcal{C}$ are of the form $T(X)/I$ where $X\in\mathcal{C}$ and $I$ is the unique maximal graded Hopf ideal in $T(V)$ generated by homogeneous elements of degree greater than or equal to $2$. According to $F(T(V))\cong T(F(V))$, we have
$F(T(V)/I)\cong T(F(V))/F(I),$
which is finite-dimensional as well. Here $F(I)$ is a homogeneous Hopf ideal of degree greater than or equal to $2$ of $T(F(V))$. Note that Nichols algebra generated by $F(V)$ must be of the form $T(F(V))/J$, where $J$ is the unique maximal homogeneous graded Hopf ideal of $T(F(V))\in\mathcal{D}$ with degree greater than or equal to $2$. Hence $F(I)\subset J$ and $T(F(V))/J\subset T(F(V))/F(I)$. On the other hand, $F^{-1}:\mathcal{D}\longrightarrow\mathcal{C}$ is inverse of $F$, which is an exact monoidal functor. Hence
$F^{-1}(T(F(V))/J)\cong T(F^{-1}(F(V))/F^{-1}(J)\cong T(V)/F^{-1}(J)\supseteq T(V)/I.$
So $F^{-1}(J)\subseteq I$, combining $F(I)\subseteq J$ implies $F(I)=J$, which leads to $F(T(V)/I)\cong T(F(V))/J$ and the proof is done. ∎
3 Monoidal equivalence between Yetter-Drinfeld module categories
We build upon the theory of Hopf pairings and partial dualization in braided monoidal categories Simon. We apply this theory to the setting of coquasi-Hopf algebras. Our main result, Theorem 3.8, establishes a key braided monoidal equivalence $\Omega$ between categories of Yetter-Drinfeld modules, which is important to our reflection theory.
3.1 Hopf pairings in braided monoidal categories
In this subsection, $\mathcal{C}$ always represents a braided monoidal category with associative isomorphism $a$ and braiding isomorphism $c$ except for special instructions. We begin by introducing the notion of Hopf pairings.
Definition 3.1.
Let $A$ and $B$ be Hopf algebras in $\mathcal{C}$. A morphism $\omega:A\otimes B\to 1$ in $\mathcal{C}$ is called a Hopf pairing, if the following identities hold:
$\omega\circ(\mu_{A}\otimes\operatorname{id}_{B})$ $=\omega\circ(\operatorname{id}_{A}\otimes\omega\otimes\operatorname{id}_{B})\circ(\operatorname{id}_{A}\otimes a^{-1}_{A,B,B})\circ a_{A,A,B\otimes B}\circ(\operatorname{id}_{A\otimes A}\otimes\Delta_{B}),$
$\omega\circ(\eta_{A}\otimes\operatorname{id}_{B})$ $=\varepsilon_{B},$
$\omega\circ(\operatorname{id}_{A}\otimes\mu_{B})$ $=\omega\circ(\operatorname{id}_{A}\otimes\omega\otimes\operatorname{id}_{B})\circ(\operatorname{id}_{A}\otimes a^{-1}_{A,B,B})\circ a_{A,A,B\otimes B}\circ(\Delta_{A}\otimes\operatorname{id}_{B\otimes B}),$
$\omega\circ(\operatorname{id}_{A}\otimes\eta_{B})$ $=\varepsilon_{A}.$
A Hopf pairing $\omega:A\otimes B\to 1$ is called non-degenerate, if there is a morphism $\omega^{\prime}:1\rightarrow B\otimes A$, such that
$(\omega\otimes\operatorname{id}_{A})\circ a^{-1}_{A,B,A}\circ(\operatorname{id}_{A}\otimes\omega^{\prime})=\operatorname{id}_{A},\ \ \text{and}\ \ (\operatorname{id}_{B}\otimes\omega)\circ a^{-1}_{B,A,B}\circ(\omega^{\prime}\otimes\operatorname{id}_{B})=\operatorname{id}_{B}.$
There are some useful properties of Hopf pairings, we list here for further use.
Remark 3.2.
[Simon, Remark 2.13, 2.16]
  • (1)
    Suppose $\omega:A\otimes B\rightarrow\textbf{1}$ is a Hopf pairing in $\mathcal{C}$, then
    $\omega\circ(\operatorname{id}\otimes\mathcal{S}_{B})=\omega\circ(\mathcal{S}_{A}\otimes\operatorname{id}).$ (3.1)
  • (2)
    If $\omega:A\otimes B\rightarrow\textbf{1}$ is a Hopf pairing, then also
    $\omega^{-}:=\omega\circ c^{-1}_{A,B}\circ(\mathcal{S}_{B}^{-1}\otimes\mathcal{S}_{A}^{-1}):B\otimes A\rightarrow\textbf{1}.$ (3.2)
    Furthermore, if $\omega$ is non-degenerate, then $\omega^{-}$ is non-degenerate as well with inverse copairing $(\mathcal{S}_{A}\otimes\mathcal{S}_{B})\circ c_{B,A}\circ\omega^{\prime}$.
  • (3)
    If $A$ has a left dual object $A^{*}$, then $A^{*}$ naturally carries the structure of a Hopf algebra in $\mathcal{C}$ such that the evaluation morphism $\operatorname{ev}_{A}:A^{*}\otimes A\rightarrow\textbf{1}$ is a Hopf pairing. Moreover, this Hopf pairing is non-degenerate with inverse copairing $\operatorname{coev}_{A}$.
Having established the properties of Hopf pairings, we now recall the important braided monoidal equivalence $\Omega^{\omega}$ induced by a certain Hopf pairing $\omega$.
Theorem 3.3.
[Simon, Theorem 3.20] Let $\omega:A\otimes B\to\mathbf{1}$ be a non-degenerate Hopf pairing. The categories $^{A}_{A}\mathcal{YD}(\mathcal{C})$} and $^{B}_{B}\mathcal{YD}(\mathcal{C})$} are braided monoidal equivalent via the monoidal functors $\Omega^{\omega}$ as follows.
$\Omega^{\omega}:{{}^{A}_{A}\mathcal{YD}(\mathcal{C})}$ $\longrightarrow{{}^{B}_{B}\mathcal{YD}(\mathcal{C})},$
$(X,\rho_{A},\delta_{A})$ $\mapsto(X,\rho_{B},\delta_{B}).$
Where $\rho_{B},\delta_{B}$ are given by
$\rho_{B}$ $=(B\otimes X\xrightarrow{c_{X,B}^{-1}}X\otimes B\xrightarrow{\delta_{A}\otimes\operatorname{id}_{B}}(A\otimes X)\otimes B\xrightarrow{a_{X,A,B}\circ(c^{-1}_{A,X}\otimes\operatorname{id}_{B})}X\otimes(A\otimes B)\xrightarrow{\operatorname{id}_{X}\otimes(\omega\circ(\operatorname{id}_{A}\otimes(\mathcal{S}_{B}^{-1})^{2}))}X)$
$\delta_{B}$ $=(X\xrightarrow{\omega^{\prime}\otimes\operatorname{id}_{X}}(B\otimes A)\otimes X\xrightarrow{a_{B,A,X}}B\otimes(A\otimes X)\xrightarrow{\operatorname{id}_{B}\otimes\rho_{A}}B\otimes X\xrightarrow{c_{X,B}\circ c_{B,X}\circ(\mathcal{S}_{B}^{2}\otimes\operatorname{id}_{X})}B\otimes X).$
Remark 3.4.
  • (1)
    Starting from a non-degenerate Hopf pairing $\omega:A\otimes B\to\mathbf{1}$. There is a non-degenerate Hopf pairing $\omega^{-}:B\otimes A\rightarrow\textbf{1}$ by Remark 3.2. Thus we obtain an equivalence:
    $\Omega^{\omega^{-}}:{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}\rightarrow{{}^{A}_{A}\mathcal{YD}(\mathcal{C})}.$ (3.3)
  • (2)
    The braided monoidal functor
    $\Omega^{\omega^{-}}\circ\Omega^{\omega}:{{}^{A}_{A}\mathcal{YD}(\mathcal{C})}\rightarrow{{}^{A}_{A}\mathcal{YD}(\mathcal{C})}$
    is isomorphic to the identity functor by [Simon, Proposition 3.22].
  • (3)
    For completeness, we recall the construction of the functor $\Omega^{\omega}$, which is investigated in [Simon, Theorem 3.16, Corollary 3.19]. Let $\omega:A\otimes B\to\mathbf{1}$ be a non-degenerate Hopf pairing. The functor $\Omega^{\omega}$ is given by
    $\Omega^{\omega}:=T\circ D.$
    The braided monoidal equivalence $T$ is given by
    $T$ $:\mathcal{YD}_{A}^{A}(\mathcal{C})\to{{}^{A}_{A}\mathcal{YD}(\mathcal{C})}$
    $T$ $(X,\rho^{r},\delta^{r})=(X,\rho^{-}\circ(\mathcal{S}_{A}^{-1}\otimes\mathrm{id}_{X}),(\mathcal{S}_{A}\otimes\mathrm{id}_{X})\circ\delta^{+}),$
    with $\rho^{-}:=\rho^{r}\circ c_{X,A}^{-1}$ and $\delta^{+}:=c_{X,A}\circ\delta^{r}$.
    Another braided monoidal equivalence $D$ is given by
    $D:{{}^{A}_{A}\mathcal{YD}(\mathcal{C})}\to{\mathcal{YD}^{B}_{B}}(\mathcal{C}),\ \ (X,\rho_{X},\delta_{X})\mapsto(X,\rho_{D(X)},\delta_{D(X)})$
    with
    $\rho_{D(X)}$ $=(\mathrm{id}\otimes\omega)\circ a_{X,A,B}\circ(c_{A,X}^{-1}\otimes S^{-1})\circ a^{-1}_{A,X,B}\circ(\delta_{X}\otimes\operatorname{id}),$
    $\delta_{D(X)}$ $=c_{B,X}\circ(\mathcal{S}_{B}\otimes\rho_{X})\circ a_{B,A,X}\circ(\omega^{\prime}\otimes\mathrm{id}).$
3.2 Partial dualization of a Hopf algebra in braided monoidal categories
We now turn our attention to the construction of partial dualization. This construction, originally developed in [Simon, Theorem 4.4], provides a powerful tool for studying Hopf algebras in braided monoidal categories.
Let us begin by recalling the setup. Consider a braided monoidal category $\mathcal{C}$ and a Hopf algebra $R$ in this category. A partial dualization datum is defined as follows.
Definition 3.5.
A partial dualization datum is denoted by $\mathcal{A}=(R\overset{\pi}{\rightarrow}A,B,\omega)$ in $\mathcal{C}$ if
$\cdot$ There is a Hopf algebra projection $\pi:R\rightarrow A$ to a Hopf subalgebra $A\subset R$ in $\mathcal{C}$.
$\cdot$ There is a Hopf algebra $B$ with a non-degenerate Hopf pairing $\omega:A\otimes B\rightarrow\textbf{1}$.
Now let $H$ be a coquasi-Hopf algebra with bijective antipode, and $\mathcal{C}={{}^{H}_{H}\mathcal{YD}}$. Suppose $R$ is a Hopf algebra in $\mathcal{C}$ and $\mathcal{A}=(R\overset{\pi}{\rightarrow}A,B,\omega)$ is a partial dualization datum. By Lemma 2.12, there is a well-defined $K\in{{}^{A}_{A}\mathcal{YD}(\mathcal{C})}$ such that $R\cong K\#A$ in $\mathcal{C}$. Combined with Theorem 3.3 discussed in the previous subsection, the following lemma is immediate.
Lemma 3.6.
With above notations, the equivalence of braided categories
$\Omega^{\omega}:{{}^{A}_{A}\mathcal{YD}(\mathcal{C})}\rightarrow{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}$
induces an equivalence of braided monoidal categories:
$^{R}_{R}\mathcal{YD}(\mathcal{C})\cong{{}^{K}_{K}\mathcal{YD}}({{}^{A}_{A}\mathcal{YD}(\mathcal{C})})\overset{\overline{\Omega}}{\longrightarrow}{{}^{L}_{L}\mathcal{YD}}({{}^{B}_{B}\mathcal{YD}(\mathcal{C})})\cong{{}^{L\#B}_{L\#B}\mathcal{YD}}(\mathcal{C}).$ (3.4)
Here $L:=\Omega(K)$ is a Hopf algebra in $^{B}_{B}\mathcal{YD}(\mathcal{C})$}.
Proof.
Since $\Omega^{\omega}$ is a braided monoidal equivalence, then $L:=\Omega(K)$ is a Hopf algebra in $^{B}_{B}\mathcal{YD}(\mathcal{C})$}. Furthermore, L#BL\#B is a Hopf algebra in $\mathcal{C}$. The braided monoidal equivalences
$^{R}_{R}\mathcal{YD}(\mathcal{C})\cong{{}^{K}_{K}\mathcal{YD}}({{}^{A}_{A}\mathcal{YD}(\mathcal{C})}),$
${{}^{L}_{L}\mathcal{YD}}({{}^{B}_{B}\mathcal{YD}(\mathcal{C})})\cong{{}^{L\#B}_{L\#B}\mathcal{YD}}(\mathcal{C})$
follow by Lemma 2.13. Each object in ${{}^{K}_{K}\mathcal{YD}}({{}^{A}_{A}\mathcal{YD}(\mathcal{C})})$ naturally belongs to $^{A}_{A}\mathcal{YD}(\mathcal{C})$}, therefore the equivalence $\Omega^{\omega}$ naturally extends to an equivalence of braided monoidal categories.
$\overline{\Omega}:{{}^{K}_{K}\mathcal{YD}}({{}^{A}_{A}\mathcal{YD}(\mathcal{C})})\longrightarrow{{}^{L}_{L}\mathcal{YD}}({{}^{B}_{B}\mathcal{YD}(\mathcal{C})}).$
We now specialize to the case which we concern.
Lemma 3.7.
Let $H$ be a coquasi-Hopf algebra with bijective antipode. Assume that $N\in{{}^{H}_{H}\mathcal{YD}}$ and $\mathcal{B}(N)$ is finite-dimensional, then we have an isomorphism of Hopf algebras in $^{H}_{H}\mathcal{YD}$}
$\mathcal{B}(N)^{*\operatorname{op,cop}}\cong\mathcal{B}(N^{*}).$
Consequently, the evaluation map $\operatorname{ev}:\mathcal{B}(N^{*})\otimes\mathcal{B}(N)\rightarrow\textbf{1}$ is a non-degenerate Hopf pairing.
Proof.
We establish this result in several steps. First, the object $\mathcal{B}(N)^{*\operatorname{op,cop}}$ with the following operations is a Hopf algebra in $^{H}_{H}\mathcal{YD}$} by [AGn99, Lemma 2.2.3]
$(\mathcal{B}(N)^{*},\mu_{{\mathcal{B}(N)^{*}}},c^{-1}_{{\mathcal{B}(N)^{*}},{\mathcal{B}(N)^{*}}},\eta_{{\mathcal{B}(N)^{*}}},c_{{\mathcal{B}(N)^{*}},{\mathcal{B}(N)^{*}}}\Delta_{{\mathcal{B}(N)^{*}}},\varepsilon_{{\mathcal{B}(N)^{*}}},\mathcal{S}_{{\mathcal{B}(N)^{*}}}).$
It is immediate that $\mathcal{B}(N)^{*}(1)=N^{*}$, thus $\mathcal{B}(N)^{*\operatorname{op,cop}}$ is generated by $N^{*}$. On the other hand, since $N^{*}\in{{}^{H}_{H}\mathcal{YD}}$, $\mathcal{B}(N^{*})$ is a Nichols algebra generated by $N^{*}$ in $^{H}_{H}\mathcal{YD}$}. By universal property of Nichols algebra, there is a surjective Hopf algebra map
$\pi:\mathcal{B}(N)^{*\operatorname{op,cop}}\longrightarrow\mathcal{B}(N^{*})$
in $^{H}_{H}\mathcal{YD}$}. The object $\mathcal{B}(N^{*})$ is the graded dual of $\mathcal{B}(N)$ by [defofnichols, Lemma 2.6]. Therefore $\operatorname{dim}(\mathcal{B}(N^{*}))=\operatorname{dim}(\mathcal{B}(N)^{*})$ since $\operatorname{dim}(\mathcal{B}(N))$ is finite-dimensional. As a result
$\mathcal{B}(N)^{*\operatorname{op,cop}}\cong\mathcal{B}(N^{*}).$
as Hopf algebras in $^{H}_{H}\mathcal{YD}$}. By Remark 3.2(3), the evaluation map $\operatorname{ev}:\mathcal{B}(N)^{*\operatorname{op},\operatorname{cop}}\otimes\mathcal{B}(N)\rightarrow\textbf{1}$ is a non-degenerate Hopf pairing. Then by an appropriate abuse of notation, we can write
$\operatorname{ev}:\mathcal{B}(N^{*})\otimes\mathcal{B}(N)\rightarrow\textbf{1},$
which confirms that this is indeed a non-degenerate Hopf pairing. ∎
This lemma will play an important role in what follows, as noted in the subsequent application of Lemma 3.6.
Theorem 3.8.
Let $H$ be a coquasi-Hopf algebra with bijective antipode, and $\mathcal{C}={{}^{H}_{H}\mathcal{YD}}$. Suppose $N=\bigoplus_{i\in I}N_{i}\in{{}^{H}_{H}\mathcal{YD}}$, and $\mathcal{B}(N_{i})$ are all finite-dimensional. For each $i\in I$, we have a partial dualization datum $\mathcal{A}_{i}=(\mathcal{B}(N)\overset{\pi_{i}}{\rightarrow}\mathcal{B}(N_{i}),\mathcal{B}(N_{i}^{*}),\operatorname{ev}_{\mathcal{B}(N_{i})}^{-})$. We set $K_{i}:=\mathcal{B}(N)^{\operatorname{co}\mathcal{B}(N_{i})}\in{{}^{\mathcal{B}(N_{i})}_{\mathcal{B}(N_{i})}\mathcal{YD}(\mathcal{C})}$. Then we have the following equivalences of braided monoidal categories
$\Omega_{i}:{{}^{\mathcal{B}(N_{i})}_{\mathcal{B}(N_{i})}\mathcal{YD}(\mathcal{C})}$ $\cong{{}^{\mathcal{B}(N_{i}^{*})}_{\mathcal{B}(N_{i}^{*})}\mathcal{YD}(\mathcal{C})},$
$\widetilde{\Omega_{i}}:{{}^{K_{i}}_{K_{i}}\mathcal{YD}}({{}^{\mathcal{B}(N_{i})}_{\mathcal{B}(N_{i})}\mathcal{YD}(\mathcal{C})})$ $\cong{{}^{\Omega_{i}(K_{i})}_{\Omega_{i}(K_{i})}\mathcal{YD}}({{}^{\mathcal{B}(N_{i}^{*})}_{\mathcal{B}(N_{i}^{*})}\mathcal{YD}(\mathcal{C})}).$
Proof.
It follows from Remark 3.4(1), Lemma 3.6 and 3.7 immediately. ∎
3.3 The functor $\Omega^{\omega}$ under $\mathbb{Z}$-gradings
We now investigate the properties of the functor $\Omega^{\omega}$.
[…]
Furthermore, suppose $H$ is a $\mathbb{N}_{0}$-graded coquasi-Hopf algebra. A $\mathbb{Z}$-graded Yetter-Drinfeld module $V$ over $H$ is by definition an object $V$ in ${{}^{H}_{H}\mathcal{YD}}(\mathbb{Z}\text{-Gr}\mathcal{M}_{k})$. In other words, $V$ is a $\mathbb{Z}$-graded vector space such that the operation $H\otimes V\rightarrow V$ and comodule structure maps $V\rightarrow H\otimes V$ are $\mathbb{Z}$-graded.
Lemma 3.9.
Suppose $H$ is a $\mathbb{Z}$-graded coquasi-Hopf algebra.
$(1)$ Let $K$ be a Nichols algebra in ${{}^{H}_{H}\mathcal{YD}}(\mathbb{Z}\operatorname{-Gr}\mathcal{M}_{\mathbbm{k}})$, and $K(1)=\oplus_{\gamma\in\mathbb{Z}}K(1)_{\gamma}$ a $\mathbb{Z}$-graded object in ${{}^{H}_{H}\mathcal{YD}}(\mathbb{Z}\operatorname{-Gr}\mathcal{M}_{\mathbbm{k}})$. Then there is a unique $\mathbb{Z}$-grading on $K$ extending the grading on $K(1)$. Moreover, $K(n)$ is $\mathbb{Z}$-graded in ${{}^{H}_{H}\mathcal{YD}}(\mathbb{Z}\operatorname{-Gr}\mathcal{M}_{\mathbbm{k}})$ for all $n\geq 0$.
$(2)$ Let $K$ be a $\mathbb{Z}$-graded braided Hopf algebra in ${{}^{H}_{H}\mathcal{YD}}(\mathbb{Z}\text{-Gr}\mathcal{M}_{k})$. Then the bosonization K#HK\#H is a $\mathbb{Z}$-graded Hopf algebra with $\deg K(\gamma)\#H(\lambda)=\gamma+\lambda$ for all $\gamma,\lambda\in\mathbb{Z}$.
$(3)$ Let $H_{0}\subseteq H$ be a coquasi-Hopf subalgebra of degree 0, and $\pi:H\rightarrow H_{0}$ a coquasi-Hopf algebra map with $\pi|_{H_{0}}=\mathrm{id}$. Define $R=H^{\operatorname{co}H_{0}}$. Then $R$ is a $\mathbb{Z}$-graded braided Hopf algebra in ${{}^{H_{0}}_{H_{0}}\mathcal{YD}}(\mathbb{Z}\text{-Gr}\mathcal{M}_{k})$ with $R(\gamma)=R\cap H(\gamma)$ for all $\gamma\in\mathbb{Z}$.
Proof.
(1) The module and comodule maps of $K(1)$ are $\mathbb{Z}$-graded and hence the infinitesimal braiding $c\in\mathrm{Aut}(K(1)\otimes K(1))$, being determined by the module and comodule maps, is $\mathbb{Z}$-graded. Moreover, the structure map of $K(n)$ for $n\in\mathbb{N}$ are determined by $c$ and $K(1)$, since $K(1)$ generates $K$. Therefore $K$ is a $\mathbb{Z}$-graded objects in ${{}^{H}_{H}\mathcal{YD}}(\mathbb{Z}\operatorname{-Gr}\mathcal{M}_{\mathbbm{k}})$.
(2) and (3) are the same as the $\mathbb{N}_{0}$-graded case. ∎
Let $A,B\in{{}^{H}_{H}\mathcal{YD}}$ be $\mathbb{N}_{0}$-graded Hopf algebras with a non-degenerate Hopf pairing $\omega$. Moreover,
$\omega(A(m),B(n))=0,\ \ \text{for all}\ m\neq n.$
Obviously, if $\mathcal{B}(N)$ is finite-dimensional and let $A=\mathcal{B}(N^{*})$, $B=\mathcal{B}(N)$, $\omega=\operatorname{ev}$. This setting satisfies the above requirement. To extend this construction to the $\mathbb{Z}$-graded setting, we extend the $\mathbb{N}_{0}$-grading to $\mathbb{Z}$-grading by
A(n)=0,B(n)=0for alln<0.A(n)=0,\ B(n)=0\ \ $\text${for all}\ n<0.
We then define a new $\mathbb{Z}$-grading on $A$ by $\operatorname{deg}(A(n))=-n$ for all $n\in\mathbb{Z}$, which ensures that $\omega:A\otimes B\rightarrow\textbf{1}$ is $\mathbb{Z}$-graded. Here, the grading of $\mathbbm{k}$ is given by $\mathbbm{k}(n)=0$, if $n\neq 0$, and $\mathbbm{k}(0)=\mathbbm{k}$.
Now we return to the purpose of this subsection. We first introduce the following notations for further use.
Definition 3.10.
Let $H$ be a coquasi-Hopf algebra with bijective antipode, and $\mathcal{C}={{}^{H}_{H}\mathcal{YD}}$. Suppose $R$ is an $\mathbb{N}_{0}$-graded Hopf algebra in $^{H}_{H}\mathcal{YD}$}. Let $X$ be a $R$-module in $\mathcal{C}$ and $Y$ be a $R$-comodule in $\mathcal{C}$. For $n\geq 0$, we define:
$\mathcal{F}_{n}X$ $=\{x\in X\mid R(i)\rhd x=0\text{ for all }i>n\},\$
$\mathcal{F}^{n}Y$ $=\{y\in Y\mid\delta_{Y}(y)\in\bigoplus_{i=0}^{n}R(i)\otimes Y\}.$
With above notations, the functor $\Omega$ relates the operator $\mathcal{F}_{n}$ to $\mathcal{F}^{n}$.
Lemma 3.11.
Let $(\mathcal{B}(N^{*}),\mathcal{B}(N),\operatorname{ev})$ be a Hopf pairing in $^{H}_{H}\mathcal{YD}$} with the equivalence of tensor category:
$\Omega:{{}^{\mathcal{B}(N)}_{\mathcal{B}(N)}\mathcal{YD}(\mathcal{C})}\cong{{}^{\mathcal{B}(N^{*})}_{\mathcal{B}(N^{*})}\mathcal{YD}(\mathcal{C})}.$
Let $V\in{{}^{\mathcal{B}(N)}_{\mathcal{B}(N)}\mathcal{YD}(\mathcal{C})}$,
(1) $\mathcal{F}_{n}\Omega(V)=\mathcal{F}^{n}V$ for all $n\geq 0$.
(2) $\mathcal{F}^{n}\Omega(V)=\mathcal{F}_{n}(V)$ for all $n\geq 0$.
(3) Suppose $V$ is a $\mathbb{Z}$-graded object in $^{\mathcal{B}(N)}_{\mathcal{B}(N)}\mathcal{YD}(\mathcal{C})$} , then $\Omega(V)$ is a $\mathbb{Z}$-graded object in $^{\mathcal{B}(N^{*})}_{\mathcal{B}(N^{*})}\mathcal{YD}(\mathcal{C})$}, where the grading of $\Omega(V)$ is given by $\Omega(V)(n)=V(-n)$ for all $n\in\mathbb{Z}$.
Proof.
Recall that the braided monoidal equivalence of abelian category
$\Omega:{{}^{\mathcal{B}(N)}_{\mathcal{B}(N)}\mathcal{YD}(\mathcal{C})}\cong{{}^{\mathcal{B}(N^{*})}_{\mathcal{B}(N^{*})}\mathcal{YD}(\mathcal{C})},$
sends
$(V,\rho_{\mathcal{B}(N)},\delta_{\mathcal{B}(N)})\mapsto(V,\rho_{\mathcal{B}(N^{*})},\delta_{\mathcal{B}(N^{*})}).$
It is induced by the non-degenerate Hopf pairing
$\operatorname{ev}^{-}:=\operatorname{ev}\circ c^{-1}_{\mathcal{B}(N^{*}),\mathcal{B}(N)}\circ(\mathcal{S}^{-1}_{\mathcal{B}(N)}\otimes\mathcal{S}^{-1}_{\mathcal{B}(N^{*})}):\mathcal{B}(N)\otimes\mathcal{B}(N^{*})\rightarrow\mathbbm{k}.$
(1) By naturality of braiding and associative isomorphism,
$\rho_{\mathcal{B}(N^{*})}$
$=(\operatorname{id}_{V}\otimes(\operatorname{ev}^{-}\circ(\operatorname{id}_{\mathcal{B}(N)}\otimes(\mathcal{S}_{\mathcal{B}(N^{*})}^{-1})^{2})))\circ a_{V,\mathcal{B}(N),\mathcal{B}(N^{*})}\circ(c^{-1}_{\mathcal{B}(N),V}\otimes\operatorname{id}_{\mathcal{B}(N^{*})})\circ(\delta_{\mathcal{B}(N)}\otimes\operatorname{id}_{\mathcal{B}(N)})\circ c^{-1}_{V,\mathcal{B}(N^{*})}$
[…]
$=((\operatorname{ev}^{-})^{-}\otimes\operatorname{id}_{V})\circ a^{-1}_{\mathcal{B}(N^{*}),\mathcal{B}(N),V}\circ(\operatorname{id}_{\mathcal{B}(N^{*})}\otimes\delta_{\mathcal{B}(N)}).$
Here $(\operatorname{ev}^{-})^{-}=\operatorname{ev}\circ c^{-1}_{\mathcal{B}(N^{*}),\mathcal{B}(N)}\circ c^{-1}_{\mathcal{B}(N),\mathcal{B}(N^{*})}\circ((\mathcal{S}^{-1}_{\mathcal{B}(N^{*})})^{2}\otimes(\mathcal{S}^{-1}_{\mathcal{B}(N)})^{2})$.
Define an auxiliary map $\rho^{\prime}:=(\operatorname{ev}\otimes\operatorname{id}_{V})\circ a^{-1}_{\mathcal{B}(N^{*}),\mathcal{B}(N),V}\circ(\operatorname{id}_{\mathcal{B}(N^{*})}\otimes\delta_{\mathcal{B}(N)})$, which is derived from $\rho_{\mathcal{B}(N^{*})}$. By [rootsys, Corollary 3.3.4, Lemma 12.2.11], $(V,\rho^{\prime})$ is still a $\mathcal{B}(N^{*})$-module in $\mathcal{C}$ and
$\mathcal{F}_{n}\Omega(V)=\mathcal{F}_{n}(V,\rho^{\prime}).$
Now for fixed $n\geq 0$, the kernel of the induced map:
$\mathcal{B}(N^{*})\otimes V\longrightarrow\operatorname{Hom}(\mathcal{B}(N)(n),V),\ \ a\otimes v\mapsto(b\mapsto\Phi(a_{-1},b_{-1},v_{-1})^{-1}\langle a_{0},b_{0}\rangle v_{0})$ (3.5)
is $\bigoplus_{m\neq n}\mathcal{B}(N^{*})(m)\otimes V$. This means, if $v\in\mathcal{F}^{n}V$, then for each $i>n$, $\rho^{\prime}(\mathcal{B}(N^{*})(i)\otimes v)=0$ by (3.5). On the other hand, if $v\in\mathcal{F}_{n}(V,\rho^{\prime})$, this will imply $\delta(v)\in\bigoplus_{i=0}^{n}\mathcal{B}(N)(i)\otimes V$. which completes the proof of (1): $\mathcal{F}_{n}\Omega(V)=\mathcal{F}_{n}(V,\rho^{\prime})=\mathcal{F}^{n}V$.
(2) For the second statement, recall $\delta_{\mathcal{B}(N^{*})}$ is given by
$\delta_{\mathcal{B}(N^{*})}=(\mathcal{S}_{\mathcal{B}(N^{*})}^{2}\otimes\operatorname{id}_{V})\circ c_{{\mathcal{B}(N^{*})},V}\circ c_{V,{\mathcal{B}(N^{*})}}\circ(\operatorname{id}_{\mathcal{B}(N^{*})}\otimes\rho_{\mathcal{B}(N)})\circ a_{{\mathcal{B}(N^{*})},{\mathcal{B}(N)},V}\circ((\operatorname{ev}^{-})^{\prime}\otimes\operatorname{id}_{V}).$
We denote $\delta^{\prime}=(\operatorname{id}_{\mathcal{B}(N^{*})}\otimes\rho_{\mathcal{B}(N)})\circ a_{{\mathcal{B}(N^{*})},{\mathcal{B}(N)},V}\circ((\operatorname{ev}^{-})^{\prime}\otimes\operatorname{id}_{V})$, which origins from $\delta_{\mathcal{B}(N)}$. Again by [rootsys, Corollary 3.3.6, Lemma 12.2.11]
$\mathcal{F}^{n}(V,\delta^{\prime})=\mathcal{F}^{n}\Omega(V).$
A similar argument to that in (1) shows that $\mathcal{F}_{n}(V)=\mathcal{F}^{n}(V,\delta^{\prime})$. This combines together shows $\mathcal{F}_{n}(V)=\mathcal{F}^{n}(V,\delta^{\prime})=\mathcal{F}^{n}\Omega(V)$.
(3) We regard $\mathcal{B}(N)$ and $\mathcal{B}(N^{*})$ as $\mathbb{Z}$-graded Hopf algebras in $^{G}_{G}\mathcal{YD}^{\Phi}$}. Let $V=\bigoplus_{n\in\mathbb{Z}}V(n)$ be its gradation, and set $\Omega(V)(n)=V(-n)$ for all $n\in\mathbb{Z}$. Then for $m\geq 0$, by definition of $\rho_{\mathcal{B}(N^{*})}$, we have
$\rho_{\mathcal{B}(N^{*})}(\mathcal{B}(N^{*})(m)\otimes\Omega(V)(n))=\rho_{\mathcal{B}(N^{*})}(\mathcal{B}(N^{*})(m)\otimes V(-n))\subseteq V(-m-n)=\Omega(V(m+n)).$
On the other hand, by definition of $\delta_{\mathcal{B}(N^{*})}$
$\delta_{\mathcal{B}(N^{*})}(\Omega(V)(n))=\delta_{\mathcal{B}(N^{*})}(V(-n))\subseteq\bigoplus_{i\in\mathbb{Z}}\mathcal{B}(N^{*})(i)\otimes V(i-n)=\bigoplus_{i\in\mathbb{Z}}\mathcal{B}(N^{*})(i)\otimes\Omega(V(n-i)).$
The above two equations imply that $\Omega(V)$ is a well-defined $\mathbb{Z}$-graded object in $^{\mathcal{B}(N^{*})}_{\mathcal{B}(N^{*})}\mathcal{YD}(\mathcal{C})$}. ∎
4 Projections of Nichols algebras
Let $(\mathbbm{k}G,\Phi)$ be a pointed coquasi-Hopf algebra, where $G$ is a finite group, $\Phi$ is a $3$-cocycle on $G$. In this section, we study projection of Nichols algebras in $^{G}_{G}\mathcal{YD}^{\Phi}$}, which will play an important role when considering reflection of simple Yetter-Drinfeld modules. The proofs of the some lemmas and theorems in this section, while are conceptually parallel to those in the Hopf algebra setting HSdualpair, takes a different path due to the non-strict monoidal structure.
4.1 The structure of the space of coinvariant $K$
To study the projection properties of Nichols algebras, we must first examine their structure in the cateogry $^{G}_{G}\mathcal{YD}^{\Phi}$}.
For any $M,N\in{{}^{G}_{G}\mathcal{YD}^{\Phi}}$, there is a canonical surjection of braided Hopf algebra in $^{G}_{G}\mathcal{YD}^{\Phi}$}
$\pi_{\mathcal{B}(N)}:\mathcal{B}(M\oplus N)\rightarrow\mathcal{B}(N),\ \ \ \pi_{\mathcal{B}(N)}\mid_{N}=\operatorname{id},\ \ \ \pi_{\mathcal{B}(N)}\mid_{M}=0.$
This surjection naturally induces a canonical projection of coquasi-Hopf algebras:
$\pi=\pi_{\mathcal{B}(N)}\#\operatorname{id}_{\mathbbm{k}G}:\mathcal{B}(M\oplus N)\#\mathbbm{k}G\rightarrow\mathcal{B}(N)\#\mathbbm{k}G.$
To simplify our notation, we introduce the following conventions:
$\mathcal{A}(M\oplus N):=\mathcal{B}(M\oplus N)\#\mathbbm{k}G,\ \mathcal{A}(N):=\mathcal{B}(N)\#\mathbbm{k}G.$
Building upon this projection structure, we can now construct the space of coinvariants. There is a natural injection $\iota:\mathcal{A}(N)\rightarrow\mathcal{A}(M\oplus N)$ such that $\pi\circ\iota=\operatorname{id}_{\mathcal{A}(N)}$.
Let
$K=\mathcal{A}(M\oplus N)^{\operatorname{co}\mathcal{A}(N)}$
be the space of right $\mathcal{A}(N)$-coinvariant elements with respect to the projection $\pi$. By virtue of Lemma 2.9, $K$ inherits the structure of a braided Hopf algebra in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$} with $\mathcal{A}(N)$-coaction:
$\delta:K\rightarrow\mathcal{A}(N)\otimes K,\ \ \ X_{-1}\otimes X_{0}=:\delta(X):=\pi(X_{1})\otimes X_{2},$
and $\mathcal{A}(N)$-action:
$\operatorname{ad}:\mathcal{A}(N)\otimes K\rightarrow K,\ \ \ \ A\otimes X\mapsto\operatorname{ad}(A)(X)=\Phi(A_{1}X_{-1},\mathbb{S}(A_{3})_{1},A_{4})(A_{2}X_{0})\mathbb{S}(A_{3})_{2}.$ (4.1)
Here, $\mathbb{S}$ denotes preantipode of $\mathcal{A}(N)$.
Denote $\pi_{\mathbbm{k}G}:\mathcal{A}(M\oplus N)\rightarrow\mathbbm{k}G$, and $\pi^{\prime}_{\mathbbm{k}G}:\mathcal{A}(N)\rightarrow\mathbbm{k}G$ be the natural projection. These maps satisfy a compatibility condition, namely, there exists a commutative diagram relating these projections:
$\mathcal{A}(M\oplus N)$$\mathcal{A}(N)$$\mathbbm{k}G$$\mathcal{A}(N)$$\pi$$\pi^{\prime}_{\mathbbm{k}G}$$\pi_{\mathbbm{k}G}$$\iota$==
Lemma 4.1.
The space $K$ coincides with $\mathcal{B}(M\oplus N)^{\text{co}\mathcal{B}(N)}$, which is precisely the space of right $\mathcal{B}(N)$-coinvariant elements with respect to $\pi_{\mathcal{B}(N)}$.
Proof.
Let $X\in K$, then $X_{1}\otimes\pi_{\mathbbm{k}G}(X_{2})=X_{1}\otimes\pi^{\prime}_{\mathbbm{k}G}\circ\pi(X_{2})=X\otimes 1$. Hence $X\in\mathcal{A}(M\oplus N)^{\operatorname{co}\mathbbm{k}G}=\mathcal{B}(M\oplus N)$. Now, we assume $X$ is homogeneous. Under this assumption, we have
$X\otimes 1=(\operatorname{id}\otimes\pi)\Delta_{\mathcal{A}(M\oplus N)}(X)=X^{1}\#x^{2}\otimes\pi(X^{2})=X^{1}\#x^{2}\otimes\pi_{\mathcal{B}(N)}(X^{2}).$
On the other hand, consider the projection $\gamma=\operatorname{id}\#\varepsilon:\mathcal{A}(M\oplus N)\rightarrow\mathcal{B}(M\oplus N)$. Applying this projection yields
$X\otimes 1=(\gamma\otimes\operatorname{id})(X\otimes 1)=X^{1}\otimes\pi_{\mathcal{B}(N)}(X^{2})=(\operatorname{id}\otimes\pi_{\mathcal{B}(N)})(\Delta_{\mathcal{B}(M\oplus N)}(X)).$
This demonstrates that $X\in\mathcal{B}(M\oplus N)^{\operatorname{co}\mathcal{B}(N)}$, completing one direction of the proof.
Conversely, let $X\in\mathcal{B}(M\oplus N)$ be a homogeneous element with
$(\operatorname{id}\otimes\pi_{\mathcal{B}(N)})\Delta_{\mathcal{B}(M\oplus N)}(X)=X^{1}\otimes\pi_{\mathcal{B}(N)}(X^{2})=X\otimes 1.$
Since $\pi_{\mathcal{B}(N)}$ is $\mathbbm{k}G$-colinear, we can deduce that
$X\otimes 1\otimes 1=X^{1}\otimes x^{2}\otimes\pi_{\mathcal{B}(N)}(X^{2}).$
which implies $X\otimes 1=X^{1}\#x^{2}\otimes\pi_{\mathcal{B}(N)}(X^{2}).$ Furthermore, observe that
$(\operatorname{id}\otimes\pi)\Delta_{\mathcal{A}(M\oplus N)}(X\#1)=X^{1}\#x^{2}\otimes\pi(X^{2})=X^{1}\#x^{2}\otimes\pi_{\mathcal{B}(N)}(X^{2})=X\otimes 1.$
Therefore, we conclude that $X\in K$, establishing the reverse inclusion.
Having characterized the coinvariant space $K$, we now turn our attention to its primitive elements. Indeed, the space of primitive elements of $K$ remains an object in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}.
Lemma 4.2.
The space of primitive elements $P(K)=\{x\in K\mid\Delta_{K}(x)=x\otimes 1+1\otimes x\}$ is a subobject of $K$ in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}.
Proof.
Let $\Delta_{K}:K\rightarrow K\otimes K$ be its comultiplication map and $\widetilde{\Delta}:K\rightarrow K\otimes K,\ x\mapsto 1\otimes x+x\otimes 1$. We consider the following map
$\widehat{\Delta}:K\rightarrow K\otimes K,\ \ \ x\mapsto\Delta_{K}(x)-\widetilde{\Delta}(x).$
$\widehat{\Delta}$ is a comodule map since both $\Delta_{K}$ and $\widetilde{\Delta}$ is colinear. Furthermore, for all $A\in{{}^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}}$ and $X\in K$,
$\widehat{\Delta}(\operatorname{ad}(A)(X))$ $=\Delta_{K}(\operatorname{ad}(A)(X))-\operatorname{ad}(A)(X)\otimes 1-1\otimes\operatorname{ad}(A)(X)$
$=\operatorname{ad}(A)(\Delta_{K}(X))-\operatorname{ad}(A)(1\otimes X+X\otimes 1)$
$=\operatorname{ad}(A)(\Delta_{K}(X)-\widetilde{\Delta}(X)).$
Therefore $\widehat{\Delta}$ is a morphism in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}. We conclude that $P(K)$ is a subobject of $K$ in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}, because $P(K)$ is precisely the kernel of the map $\widehat{\Delta}$. ∎
With the structural properties of $P(K)$ established, we now investigate the explicit form of $K$ through its relation to the adjoint map. Since $K=\mathcal{B}(M\oplus N)^{\text{co}\mathcal{B}(N)}$, it is natural to consider $\text{ad}(\mathcal{B}(N))(M)$.
Lemma 4.3.
Let $M,N\in{{}^{G}_{G}\mathcal{YD}^{\Phi}}$, $K=(\mathcal{A}(M\oplus N))^{\operatorname{co}\mathcal{A}(N)}$. The standard $\mathbb{N}_{0}$-grading of $\mathcal{B}(M\oplus N)$ induces an $\mathbb{N}_{0}$-grading on
$L:=\operatorname{ad}(\mathcal{B}(N))(M)=\bigoplus_{n\in\mathbb{N}}\operatorname{ad}(N)^{n}(M)$
with degree $\operatorname{deg}(\operatorname{ad}(N)^{n}(M))=n+1$. Then $L$ is a $\mathbb{N}_{0}$-graded object in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}. Moreover, $L\subseteq K$ is a subobject of $K$ in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}.
Proof.
Let $A\in N$ and $X\in\mathcal{B}(M\oplus N)$ be homogeneous elements. By equation (2.30), we have
$\operatorname{ad}(A)(X)=AX-\operatorname{ad}(a)(X)A.$
The element $\operatorname{ad}(A)(X)$ is of degree $\operatorname{deg}(X)+1$ in the Nichols algebra $\mathcal{B}(M\oplus N)$. Moreover, we have $\operatorname{ad}(N)^{n}(M)\cap\operatorname{ad}(N)^{m}(M)=0$ for $n\neq m$. This is implies the direct sum decomposition of $L$.
Since $M\subseteq K$ and $K\in{{}^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}}$, we conclude that $L=\operatorname{ad}(\mathcal{B}(N))(M)\subseteq K$. Note that
$\operatorname{ad}(\mathcal{B}(N))(M)\subseteq\operatorname{ad}(\mathcal{A}(N))(M).$ Conversely, since $M\in{{}^{G}_{G}\mathcal{YD}^{\Phi}}$, we have
$\operatorname{ad}(\mathcal{A}(N))(M)\subseteq\operatorname{ad}(\mathcal{B}(N))(\mathbbm{k}G\rhd M)\subseteq\operatorname{ad}(\mathcal{B}(N))(M).$
Hence $L=\operatorname{ad}(\mathcal{A}(N))(M)$.
It is clear that the operation
$\operatorname{ad}:\mathcal{A}(N)\otimes L\rightarrow L$
is well-defined since $\operatorname{ad}$ satisfies the first axiom of Yetter-Drinfeld modules. Furthermore, the operation $\operatorname{ad}:\mathcal{A}(N)\otimes L\rightarrow L$ is $\mathbb{N}_{0}$-graded. Next, we want to show $L$ is a $\mathcal{A}(N)$-comodule. The comodule structure is given by $\delta_{K}$. Now we assume $A\in\mathcal{B}(N)$ and $X\in M$ are homogeneous elements. A direct computation shows:
$\delta_{K}(\operatorname{ad}(A)(X))$ $=p((\operatorname{ad}(A_{3})(X))_{-1},A_{4})q(A_{1}x,\mathcal{S}(A_{6}))(A_{2}x)\mathcal{S}(A_{5})\otimes(\operatorname{ad}(A_{3})(X))_{0}.$
Note that the condition
$\operatorname{ad}(A_{3})(X)_{-1}\notin G\Longrightarrow p((\operatorname{ad}(A_{3})(X))_{-1},A_{3})=0.$
Therefore, since $\operatorname{ad}(A_{3})(X)$ is homogeneous, we have
$\delta_{K}(\operatorname{ad}(A)(X))=p(\operatorname{deg}(\operatorname{ad}(A_{3})(X)),A_{4})q(A_{1}x,\mathcal{S}(A_{6}))(A_{2}x)\mathcal{S}(A_{5})\otimes(\operatorname{ad}(A_{3})(X)),$
where $\operatorname{deg}(\operatorname{ad}(A_{2})(X))\in G$ represents its associated degree. This shows $\delta_{K}:L\rightarrow\mathcal{A}(N)\otimes L$ is well-defined. The third axiom holds automatically since $L\subseteq K$. Thus $L$ is a $\mathbb{N}_{0}$-graded object in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$} and is a subobject of $K$.
Lemma 4.4.
For all $X\in L$, we have $\Delta_{\mathcal{B}(M\oplus N)}(X)-X\otimes 1\in\mathcal{B}(N)\otimes L$.
Proof.
Recall $L=\bigoplus_{n\in\mathbb{N}}\operatorname{ad}(N)^{n}(M)$. We proceed by induction on $n$. Now let $A\in N$ and $X\in M$ be homogeneous elements. A direct computation yields:
$\Delta_{\mathcal{B}(M\oplus N)}(\operatorname{ad}(A)(X))$ $=\Delta_{\mathcal{B}(M\oplus N)}(AX-(\operatorname{ad}(a)(X)A))$
$=AX\otimes 1+1\otimes AX+A\otimes X+\operatorname{ad}(a)(X)\otimes A$
$-\operatorname{ad}(a)(X)A\otimes 1-1\otimes\operatorname{ad}(a)(X)A-\operatorname{ad}(a)(X)\otimes A-\operatorname{ad}(axa^{-1})(A)\otimes\operatorname{ad}(a)(X)$
$=\operatorname{ad}(A)(X)\otimes 1+1\otimes\operatorname{ad}(A)(X)+A\otimes X-\operatorname{ad}(axa^{-1})(A)\otimes\operatorname{ad}(a)(X).$
Note that $X=\operatorname{ad}(1)(X)\in L$, therefore $\Delta_{\mathcal{B}(M\oplus N)}(\operatorname{ad}(A)(X))-\operatorname{ad}(A)(X)\otimes 1\in\mathcal{B}(N)\otimes L$.
For a fixed $m\in\mathbb{N}_{0}$. We assume that for all $Y\in\operatorname{ad}(N)^{m^{\prime}}(M)$, $1\leq m^{\prime}\leq m$, we have
$\Delta_{\mathcal{B}(M\oplus N)}(Y)-Y\otimes 1\in\mathcal{B}(N)\otimes L.$
Now let $A\in N$ and $X\in\operatorname{ad}(N)^{m}(M)$ be homogeneous, we have
$\Delta_{\mathcal{B}(M\oplus N)}(\operatorname{ad}(A)(X))=\Delta_{\mathcal{B}(M\oplus N)}(AX-(\operatorname{ad}(a)(X))A)$
$=(A\otimes 1+1\otimes A)(X^{1}\otimes X^{2})$
$-\frac{\Phi(a,x^{1},x^{2})\Phi(ax^{1}a^{-1},ax^{2}a^{-1},a)}{\Phi(ax^{1}a^{-1},a,x^{2})}(\operatorname{ad}(a)(X^{1})\otimes\operatorname{ad}(a)(X^{2}))(1\otimes A+A\otimes 1)$
$=\operatorname{ad}(A)(X)\otimes 1+\sum_{X^{1}\otimes X^{2}\neq X\otimes 1}\Phi(a,x^{1},x^{2})AX^{1}\otimes X^{2}+\frac{\Phi(a,x^{1},x^{2})}{\Phi(ax^{1}a^{-1},a,x^{2})}\operatorname{ad}(a)(X^{1})\otimes AX^{2}$
$-\frac{\Phi(a,x^{1},x^{2})}{\Phi(ax^{1}a^{-1},a,x^{2})}\operatorname{ad}(a)(X^{1})\otimes\operatorname{ad}(a)(X^{2})A$
$-\frac{\Phi(a,x^{1},x^{2})\Phi(ax^{1}a^{-1},ax^{2}a(ax^{2})^{-1},a)}{\Phi(ax^{1}a^{-1},a,x^{2})}\operatorname{ad}(a)(X^{1})\operatorname{ad}(ax^{2}a^{-1})(A)\otimes\operatorname{ad}(a)(X^{2})$
$=\operatorname{ad}(A)(X)\otimes 1+\sum_{X^{1}\otimes X^{2}\neq X\otimes 1}\Phi(a,x^{1},x^{2})AX^{1}\otimes X^{2}+\frac{\Phi(a,x^{1},x^{2})}{\Phi(ax^{1}a^{-1},a,x^{2})}\operatorname{ad}(a)(X^{1})\otimes\operatorname{ad}(A)(X^{2})$
$-\frac{\Phi(a,x^{1},x^{2})\Phi(ax^{1}a^{-1},ax^{2}a(ax^{2})^{-1},a)}{\Phi(ax^{1}a^{-1},a,x^{2})}\operatorname{ad}(a)(X^{1})\operatorname{ad}(ax^{2}a^{-1})(A)\otimes\operatorname{ad}(a)(X^{2}).$
By assumptions,
$\Delta_{\mathcal{B}(M\oplus N)}(X)-X\otimes 1\in\mathcal{B}(N)\otimes L.$
Therefore the second term lies in $\mathcal{B}(N)\otimes{L}$. Since $A\in N$, $\operatorname{ad}(a)(X^{1})\in\mathcal{B}(N)$ and
$\operatorname{ad}(A)(X^{2})\in\operatorname{ad}(N)(L)=\operatorname{ad}(N)(\bigoplus_{n\in\mathbb{N}_{0}}\operatorname{ad}(N)^{n}(M))\in L,$
the third term lies in $\mathcal{B}(N)\otimes{L}$. For the last term, by Lemma 4.3, we have $L\in{{}^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}}$. Hence it belongs to $\mathcal{B}(N)\otimes{L}$ as well, which completes the induction and proves the lemma.
4.2 The semisimplicity of the object $L$
We now turn to the structural analysis of the object $L$, whose properties are useful to our subsequent arguments regarding Nichols algebras.
Recall that a graded coalgebra is a coalgebra $C$ provided with a grading $C=\bigoplus_{m\in\mathbb{N}_{0}}C^{m}$ such that $\Delta(C^{m})\subseteq\bigoplus_{i+j=m}C^{i}\otimes C^{j}$. Let $\Delta_{i,j}:C^{m}\rightarrow C^{i}\otimes C^{j}$ denote the map $\text{pr}_{i,j}\Delta$, where $m=i+j$ and $\operatorname{pr}_{i,j}:C\otimes C\rightarrow C^{i}\otimes C^{j}$. More generally, if $i_{1},\ldots,i_{n}\in\mathbb{N}_{0}$ and $i_{1}+\cdots+i_{n}=m$, then $\Delta_{i_{1},\ldots,i_{n}}$ is defined by the commutative diagram
$C^{m}$$\bigoplus\limits_{j_{1}+j_{2}+\cdots+j_{n}=m}C^{j_{1}}\otimes C^{j_{2}}\otimes\cdots\otimes C^{j_{n}}$$C^{i_{1}}\otimes C^{i_{2}}\otimes\cdots\otimes C^{i_{n}}$$\Delta^{n-1}$$\operatorname{pr}_{i_{1},..,i_{n}}$$\Delta_{i_{1},..,i_{n}}$
This framework naturally leads us to consider the maximal coideal in a graded coalgebra. For notational convenience, we write $\Delta_{1,1,...,1}$ by $\Delta_{1^{n}}$. Let $I_{C}(n)=\ker(\Delta_{1^{n}})$ for all $n\geq 1$, and
$I_{C}=\bigoplus_{n\geq 1}I_{C}(n)=\bigoplus_{n\geq 2}I_{C}(n).$
Note that $I_{C}(1)=0$ since $\Delta_{1}=\text{id}$.
Lemma 4.5.
[rootsys, Lemma 1.3.13, Proposition 1.3.14]
$(1)$ Assume that $C$ is connected, that is, $C^{0}$ is one dimensional. Then $I_{C}\subseteq C$ is a coideal of $C$.
$(2)$ Assume that $C$ is connected. Then the following are equivalent.
$(a)$ $C$ is strictly graded, that is, $P(C)=C^{1}$.
$(b)$ For all $n\geq 2$, $\Delta_{1^{n}}:C^{n}\to(C^{1})^{\otimes n}$ is injective.
$(c)$ For all $n\geq 2$, $\Delta_{1,n-1}:C^{n}\to C^{1}\otimes C^{n-1}$ is injective.
$(d)$ $I_{C}=0$.
The following technical result provides the necessary machinery to analyze the semisimplicity properties of $L$.
Lemma 4.6.
[rootsys,Proposition 13.2.3 Let $C$ be an $\mathbb{N}_{0}$-graded coalgebra, $X$ an $\mathbb{N}_{0}$-graded left $C$-comodule, and $Y$ a $C$-subcomodule of $X$. Let $k>0$ be an integer. Assume that the map $\delta_{n-k,k}^{X}:X(n)\to C(n-k)\otimes X(k)$ is injective for all $n\geq k$, and $Y$ is not contained in $\bigoplus_{i=0}^{k-1}X(i)$. Then $Y\cap\bigoplus_{i=0}^{k}X(i)\neq 0$.
Lemma 4.7.
If $L^{\prime}\subseteq L$ is a non-zero $\mathcal{A}(N)$-subcomodule, then $L^{\prime}\cap M\neq 0$.
Proof.
We begin by observing that $\mathcal{B}(M\oplus N)$ carries a natural left $\mathcal{B}(N)$-comodule structure in the category $^{G}_{G}\mathcal{YD}^{\Phi}$}, defined via the composition:
$\delta:\mathcal{B}(M\oplus N)\xrightarrow{\Delta_{\mathcal{B}(M\oplus N)}}\mathcal{B}(M\oplus N)\otimes\mathcal{B}(M\oplus N)\xrightarrow{\pi_{\mathcal{B}(N)}\otimes\operatorname{id}}\mathcal{B}(N)\otimes\mathcal{B}(M\oplus N).$
Now let $X\in L$, then $\delta(X)\in\mathcal{B}(N)\otimes L$ by Lemma 4.4, which implies that L inherits an $\mathbb{N}_{0}$-graded $\mathcal{B}(N)$-comodule via the restriction of $\delta:L\rightarrow\mathcal{B}(N)\otimes L$.
Suppose $X\in L(n)$ for $n\geq 1$, consider the component map $\delta_{n-1,1}:L(n)\rightarrow\mathcal{B}(N)(n-1)\otimes L(1)=\mathcal{B}(N)(n-1)\otimes M.$ By Lemma 4.4, $\Delta_{\mathcal{B}(M\oplus N)}(X)=X\otimes 1+\mathcal{B}(N)\otimes L$, it follows that
$\delta_{n-1,1}(X)=(\Delta_{\mathcal{B}(M\oplus N)})_{n-1,1}(X).$
As Nichols algebras are strictly graded connected coalgebras, the map $(\Delta_{\mathcal{B}(M\oplus N)})_{n-1,1}$ is injective, and hence so is $\delta_{n-1,1}$. We now apply Lemma 4.6 with $C={{}^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}}$, $X=L$, $Y=L^{\prime}$ and $k=1$. Since $L(0)=0$ by its gradation. Therefore, we conclude that $L^{\prime}\cap L(1)=L^{\prime}\cap M\neq 0$, which completes the proof. ∎
Now we are ready to show the semisimplicity of $L$.
Lemma 4.8.
$(1)$ Assume $M=\bigoplus_{i\in I}M_{i}$ is a direct sum in $^{G}_{G}\mathcal{YD}^{\Phi}$}. Let $L_{i}=\operatorname{ad}\mathcal{B}(N)(M_{i})$ for all $i\in I$. Then we have such a decomposition in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}
$L=\bigoplus_{i\in I}L_{i}.$
$(2)$ If $M$ is irreducible in $^{G}_{G}\mathcal{YD}^{\Phi}$}, then $L=\operatorname{ad}\left(\mathcal{B}(N)\right)(M)$ is irreducible in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}.
Proof.
(1) It suffices to show that the sum is direct. Suppose, for contradiction, that the sum is not a direct sum. Then there exists an index $k\in I$ such that $L_{k}\cap\sum_{i\neq k}L_{i}\neq 0$. By Lemma 4.7, this implies
$(L_{k}\cap\sum_{i\neq k}L_{i})\cap M_{k}\neq 0,$
hence $\sum_{i\neq k}L_{i}\cap M_{k}\neq 0$. But since $M_{k}$ lies in degree one in $L_{k}$, we obtain
$\sum_{i\neq k}M_{i}\cap M_{k}\neq 0,$
contradicting the directness of the sum $M=\bigoplus_{i\in I}M_{i}.$
(2) Now let $0\neq L^{\prime}\subseteq L$ be a subobject in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}, then $L^{\prime}\cap M\neq 0$ by Lemma 4.7. Furthermore, $L^{\prime}\cap M$ is a subobject of $M$ in $^{G}_{G}\mathcal{YD}^{\Phi}$}. But $M$ is irreducible in $^{G}_{G}\mathcal{YD}^{\Phi}$}, then $L^{\prime}\cap M=M$. Hence
$L=\operatorname{ad}(\mathcal{A}(N))(M)=\operatorname{ad}(\mathcal{A}(N))(L^{\prime}\cap M)\subseteq L^{\prime},$
which implies $L^{\prime}=L$. We conclude that $L$ is irreducible. ∎
4.3 Realizing $K$ as a Nichols algebra
We now proceed to prove that $K$ possesses the structure of a Nichols algebra in an appropriate Yetter-Drinfeld module category.
Theorem 4.9.
There is an isomorphism
$K\cong\mathcal{B}(L)$ (4.2)
of Hopf algebras in the category $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}. In particular, $P(K)=L$.
To prepare for the proof of this theorem, we first establish a generation property:
Lemma 4.10.
The algebra $K$ is generated by $L$ in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}.
Proof.
We have previously established that $L\subseteq K$ is a subobject of $K$ in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}. Let $K^{\prime}$ be the subalgebra of $K$ generated by $L$ in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}. Obviously, $K^{\prime}$ inherits the structure of an object in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$} with structures induced by tensor product of $L$.
Recall from Lemma 4.3 that we have an algebra isomorphism $K\#\mathcal{B}(N)\cong\mathcal{B}(M\oplus N)$ in $^{G}_{G}\mathcal{YD}^{\Phi}$} where the multiplication is induced by that of $K\#(\mathcal{B}(N)\#\mathbbm{k}G)$. Now we let $W$ be the image of $K^{\prime}\#\mathcal{B}(N)$ under the isomorphism $K\#\mathcal{B}(N)\cong\mathcal{B}(M\oplus N)$. To prove the lemma, it suffices to show that $W=\mathcal{B}(M\oplus N)$. Note that $M\oplus N\subseteq W$ since $M,N\subseteq W$. Thus, we need only verify that $W$ forms a subalgebra of $\mathcal{B}(M\oplus N)$.
A crucial observation is that $K^{\prime}$ remains stable under the operation induced by $\mathcal{B}(N)$. Indeed, $L$ itself is stable under the operation of $\mathcal{B}(N)$. Now consider arbitrary homogeneous elements $Y,Z\in L$ and $A\in\mathcal{A}(N)$, by (2.12),
$\operatorname{ad}(A)(Y\otimes Z)$
$=\frac{\Phi(A_{1},Y_{-1},Z_{-2})\Phi((\operatorname{ad}(A_{2})(Y_{0}))_{-1},(\operatorname{ad}(A_{4})(Z_{0}))_{-1},A_{5}))}{\Phi((\operatorname{ad(A_{2})(Y_{0})})_{-2},A_{3},Z_{-1})}(\operatorname{ad}(A_{2})(Y_{0}))_{0}\otimes(\operatorname{ad}(A_{4})(Z_{0}))_{0}$
$=\frac{\Phi(A_{1},Y_{-1},Z_{-2})\Phi((\operatorname{deg}(\operatorname{ad}(A_{2})(Y_{0})),\operatorname{deg}(\operatorname{ad}(A_{4})(Z_{0})),A_{5}))}{\Phi((\operatorname{deg}(\operatorname{ad}(A_{2})(Y_{0})),A_{3},Z_{-1})})\operatorname{ad}(A_{2})(Y_{0})\otimes\operatorname{ad}(A_{4})(Z_{0}).$
Since the associator vanishes if any component does not lie in $G$. Note that $Y_{0},Z_{0}\in L$, it follows that $\operatorname{ad}(A)(Y\otimes Z)\in L\otimes L$. It follows that $K^{\prime}$ remains stable under the operation induced by $\mathcal{B}(N)$ by induction.
As both $K^{\prime}$ and $\mathcal{B}(N)$ are subalgebras of $\mathcal{B}(M\oplus N)$, we only need to examine the case when $A\in N$ and $X\in K^{\prime}$ are homogeneous. In this situation:
$\operatorname{ad}(A)(X)=AX-\operatorname{ad}(a)(X)A.$
Since $\operatorname{ad}(A)(X)\in K^{\prime}$ and $\operatorname{ad}(a)(X)A\in K^{\prime}\#N$, we conclude that $AX\in W$. Thus $W$ is a subalgebra of $\mathcal{B}(M\oplus N)$, which establishes that $K$ is generated by $L$ in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}. ∎
To complete our analysis, we require an additional result concerning coalgebra filtrations. Let $C$ be a coalgebra. Recall that an $\mathbb{N}_{0}$-filtration of $C$ is a family of subspaces $C_{n}$, $n\geq 0$, of $C$ satisfying
  • $C_{n}$ is a subspace of $C_{m}$ for all $m,n\in\mathbb{N}_{0}$ with $n\leq m$
  • $C=\bigcup_{n\in\mathbb{N}_{0}}C_{n}$,
  • $\Delta_{C}(x)\in\sum_{i=0}^{n}C_{i}\otimes C_{n-i}$ for all $x\in C_{n}$, $n\in\mathbb{N}_{0}$.
Lemma 4.11.
Let $C$ be a coalgebra having an $\mathbb{N}_{0}$-filtration $\{C_{n}\}_{n\geq 0}$. Let $U$ be a non-zero comodule of $C$. Then there exists $u\in U\setminus\{0\}$ such that $\delta(u)\in C_{0}\otimes U$.
Proof.
Let $x\in U\setminus\{0\}$. Then there exists $n\in\mathbb{N}_{0}$ with $\delta(x)\in C_{n}\otimes U$. If $n=0$, we are done. Assume now that $n\geq 1$ and let $\pi_{0}:C\to C/C_{0}$ be the canonical linear map. Since $C=\bigcup_{n\in\mathbb{N}_{0}}C_{n}$ is a coalgebra filtration, there is a maximal $m\in\mathbb{N}_{0}$ such that
$\pi_{0}(x_{-m})\otimes\cdots\otimes\pi_{0}(x_{-1})\otimes x_{0}\neq 0.$
Let $f_{1},\ldots,f_{m}\in C^{*}$ with $f_{i}|_{C_{0}}=0$ for all $i\in\{1,\ldots,m\}$ such that
$y:=f_{1}(x_{-m})\cdots f_{m}(x_{-1})x_{0}\neq 0.$
Then $\delta(y)=f_{1}(x_{-m-1})\cdots f_{m}(x_{-2})x_{-1}\otimes x_{0}\in C_{0}\otimes U$ by the maximality of $m$. ∎
Proof of Theorem 4.9.
We endow a non-standard grading on $\mathcal{B}(M\oplus N)$ via setting $\operatorname{deg}(M)=1$ and $\operatorname{deg}(N)=0$. Under this grading, $\mathcal{B}(M\oplus N)$ remains a $\mathbb{N}_{0}$-graded Hopf algebra in $^{G}_{G}\mathcal{YD}^{\Phi}$}. Furthermore, $\mathcal{B}(M\oplus N)\#\mathbbm{k}G$ becomes a $\mathbb{N}_{0}$-graded coquasi-Hopf algebra with $\operatorname{deg}(\mathbbm{k}G)=0$. It follows direct that $K$ inherits an $\mathbb{N}_{0}$-graded Hopf algebra in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$} with $K(n)=K\cap(\mathcal{B}(M\oplus N)\#\mathbbm{k}G)(n)$. In particular, we have $K(1)=K\cap(\mathcal{B}(M\oplus N)\#\mathbbm{k}G)(1)=K\cap(\mathcal{B}(M\oplus N)(1))=\operatorname{ad}(\mathcal{B}(N))(M)=L$. Hence $K$ is generated its degree one component $K(1)$.
It remains to show that $P(K)=K(1)$. That is, there is no primitive element in $K(n)$, $n\geq 2$. Suppose, for contradiction, that there exists a nonzero subspace $U\subseteq P(K(n))$ for some positive integer $n$. By Lemma 4.2, $U$ is an object in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}, since $\operatorname{deg}(N)=\operatorname{deg}(\mathbbm{k}G)=0$. Now, consider the coalgebra filtration on $\mathcal{B}(N)\#\mathbbm{k}G$ given by $(\mathcal{B}(N)\#\mathbbm{k}G)_{0}=\mathbbm{k}G$, $(\mathcal{B}(N)\#\mathbbm{k}G)_{1}=\mathbbm{k}G+N$. Applying Lemma 4.11, we find a nonzero element $u\in U$ such that $\delta_{K}(u)=u_{-1}\otimes u_{0}\in\mathbbm{k}G\otimes U$. Note that
$\delta_{K}(u)=(\pi\otimes\operatorname{id})\Delta_{\mathcal{A}(M\oplus N)}(u).$
Therefore
$\Delta_{\mathcal{A}(M\oplus N)}(u)=u\otimes 1+u_{-1}\otimes u_{0}.$
Applying the map $\mathrm{id}\#\varepsilon\otimes\mathrm{id}$, we obtain
$\Delta_{\mathcal{B}(M\oplus N)}(u)=(\operatorname{id}\#\varepsilon\otimes\operatorname{id})\Delta(u)=1\otimes u+u\otimes 1,$
which implies that $u$ is a primitive element in $\mathcal{B}(M\oplus N)$. However, since $K(n)$ is generated by $L=\operatorname{ad}(\mathcal{B}(N))(M)$, the element $u$ must have degree at least $n$ in the standard grading of $\mathcal{B}(M\oplus N)$. This leads to a contradiction, as primitive elements in a Nichols algebra lie in degree one. ∎
4.4 From the Nichols algebra back to the space of coinvariants
We now turn to a converse of the preceding result under an additional restriction.
Let $C$ be a coalgebra and $D\subseteq C$ be a subcoalgebra. Let $V$ be a left comodule of $C$ with comodule structure $\delta:V\rightarrow C\otimes V$. We denote the largest $D$-subcomodule of $V$ by
$V(D)=\{v\in V\mid\delta(v)\in D\otimes V\}.$
Lemma 4.12.
Let $N\in{{}^{G}_{G}\mathcal{YD}^{\Phi}}$ and $W\in{{}^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}}$. Assume that $W=\bigoplus_{i\in I}W_{i}$ is a decomposition into irreducible objects in ${{}^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}}(\mathbb{Z}\operatorname{-Gr}\mathcal{M}_{\mathbbm{k}})$. Let $M=W(\mathbbm{k}G)$, and $M_{i}=M\cap W_{i}$ for all $i\in I$.
$(1)$ $M=\bigoplus_{i\in I}M_{i}$ is a decomposition into irreducible objects in ${{}^{G}_{G}\mathcal{YD}^{\Phi}}(\mathbb{Z}\operatorname{-Gr}\mathcal{M}_{\mathbbm{k}})$.
$(2)$ For all $i\in I$, $M_{i}$ is the $\mathbb{Z}$-homogeneous component of $W_{i}$ of minimal degree, and $W_{i}=\mathcal{B}(N)\rhd M_{i}=\bigoplus_{n\geq 0}N^{\otimes n}\rhd M_{i}$.
Proof.
Let $W=\bigoplus_{n\in\mathbb{Z}}W(n)$ be the $\mathbb{Z}$-grading of $W$ in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}. Then $M$ is a $\mathbb{Z}$-graded object in $^{G}_{G}\mathcal{YD}^{\Phi}$} with homogeneous components $M(n)=M\cap W(n)$ for all $n\in\mathbb{Z}$. It is clear that $M=\bigoplus_{n\in\mathbb{Z}}\bigoplus_{i\in I}M_{i}(n)=\bigoplus_{i\in I}M_{i}$, where $M_{i}=M\cap W_{i}=W_{i}(\mathbbm{k}G)$.
We now show that each $M_{i}$ is an irreducible object in ${{}^{G}_{G}\mathcal{YD}^{\Phi}}(\mathbb{Z}\operatorname{-Gr}\mathcal{M}_{\mathbbm{k}})$ for each $i\in I$. First, by Lemma 4.11, we have $M_{i}\neq 0$. Suppose $0\neq M_{i}^{\prime}\subseteq M_{i}$ is a homogeneous subobject in $^{G}_{G}\mathcal{YD}^{\Phi}$}, and let $m_{i}$ be the degree of $M_{i}^{\prime}$. Define
$W_{i}^{\prime}:=\mathcal{B}(N)\rhd M_{i}^{\prime}.$
Using a method similar to that in Lemma 4.3, one verifies that $W_{i}^{\prime}$ is an $\mathbb{Z}$-graded $\mathcal{A}(N)$-subcomodule. Moreover, $\mathcal{A}(N)\rhd(\mathcal{B}(N)\rhd M_{i}^{\prime})\subseteq\mathcal{A}(N)\rhd M_{i}^{\prime}=\mathcal{B}(N)\rhd M_{i}^{\prime}$. Therefore $W_{i}^{\prime}$ is a $\mathbb{Z}$-graded subobject of $W_{i}$ in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}. The object $W_{i}^{\prime}$ is concentrated in degrees $\geq m_{i}$, and its degree $m_{i}$’s component is exactly $M_{i}^{\prime}$. Since $W_{i}$ is irreducible, we must have $W_{i}^{\prime}=W_{i}$. Therefore, $M_{i}^{\prime}=W_{i}^{\prime}\cap M=W_{i}\cap M=M_{i}$, which establishes the irreducibility of $M_{i}$. Moreover, $M_{i}$ is indeed the homogeneous component of $W_{i}$ of minimal degree. Finally, for each $i\in I$ and $n\in\mathbb{N}_{0}$,
$\operatorname{deg}(N^{\otimes n}\rhd M_{i})=n+\operatorname{deg}(M_{i}),$
since the map $\mathcal{B}(N)\#\mathbbm{k}G\otimes W_{i}\rightarrow W_{i}$ is $\mathbb{Z}$-graded. It follows that $\mathcal{B}(N)\rhd M_{i}=\bigoplus\limits_{n\geq 0}N^{\otimes n}\rhd M_{i}$ as required. ∎
Theorem 4.13.
Let $N\in{{}^{G}_{G}\mathcal{YD}^{\Phi}}$ and $W$ be a semisimple object in the category ${{}^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}}(\mathbb{Z}\operatorname{-Gr}\mathcal{M}_{\mathbbm{k}})$. Here $\mathcal{A}(N)$ is equipped with the standard $\mathbb{N}_{0}$-grading. Let $K=\mathcal{B}(W)$ be the Nichols algebra of $W$ in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$} and define $M=W(\mathbbm{k}G)$. Then there is a unique isomorphism
$K\#\mathcal{B}(N)\cong\mathcal{B}(M\oplus N)$ (4.3)
of braided Hopf algebras in $^{G}_{G}\mathcal{YD}^{\Phi}$} which is the identity on $M\oplus N$.
Proof.
Let $W=\bigoplus_{i\in I}W_{i}$ be the decomposition of $W$ into irreducible objects in the category
${{}^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}}(\mathbb{Z}\operatorname{-Gr}\mathcal{M}_{\mathbbm{k}})$. For each $i\in I$, define $M_{i}=M\cap W_{i}$. By Lemma 4.12, $W_{i}=\bigoplus_{n\in I}N^{\otimes n}\rhd M_{i}$. Let $\operatorname{deg}(M_{i})=m_{i}$ for each $i\in I$. We endow $W_{i}$ a new grading by $\widetilde{W_{i}}=W_{i}$ with
$\widetilde{W_{i}}(n)=W(n+m_{i}-1)=N^{n-1}\rhd M_{i},\ \text{for all}\ n\in\mathbb{N}_{0}.$
[…]
$\mathcal{B}(\widetilde{W})(0)=\mathbbm{k},\ \ \mathcal{B}(\widetilde{W})(1)=M.$
Moreover, $K\#(\mathcal{B}(N)\#\mathbbm{k}G)$ is a $\mathbb{N}_{0}$-graded coquasi-Hopf algebra, and
$R:=K\#\mathcal{B}(N)=(K\#(\mathcal{B}(N)\#\mathbbm{k}G))^{\operatorname{co}\mathbbm{k}G}$
is a $\mathbb{N}_{0}$-graded Hopf algebra in $^{G}_{G}\mathcal{YD}^{\Phi}$}. Its degree 0’s part is $\mathbbm{k}$, and its degree $1$’s part is $M\oplus N$.
We now show that $R$ is generated by $M\oplus N$ as an algebra in $^{G}_{G}\mathcal{YD}^{\Phi}$}. It is direct to see that $R$ is generated by $K(1)=\mathcal{B}(N)\rhd M$ and $N$. It therefore suffices to prove that $\mathcal{B}(N)\rhd M=\bigoplus_{n\geq 0}N^{\otimes n}\rhd M$ is contained in the subalgebra generated by $\mathcal{B}(N)$ and $M$. To see this, we proceed by induction on $n$. For the base case, take homogeneous elements $X\in M$ and $Y\in N$. Inside the coquasi-Hopf algebra $K\#(\mathcal{B}(N)\#\mathbbm{k}G)$, we have
$YX=(Y_{1}\rhd X)Y_{2}=(y\rhd X)Y+(Y\rhd X).$
Since $M\in{{}^{G}_{G}\mathcal{YD}^{\Phi}}$ and $y\in G$, the term $Y\rhd X$ is contained in the subalgebra generated by $\mathcal{B}(N)$ and $M$. Now assume for some $k\geq 0$, that for all homogeneous $Z\in\mathcal{B}(N)(k)$ and $X\in M$, the element $Z\rhd X$ lies in the subalgebra generated by $\mathcal{B}(N)$ and $M$, then for homogeneous $Y\in N$,
$(YZ)\rhd X$ $=\frac{\Phi(y,(Z_{2}\rhd X)_{-1},Z_{3}))}{\Phi(y,Z_{1},x)}Y\rhd(Z_{2}\rhd X)_{0}$
$=\frac{\Phi(y,(Z_{2}\rhd X)_{-1},Z_{3}))}{\Phi(y,Z_{1},x)}(Y(Z_{2}\rhd X)_{0}-(y\rhd(Z_{2}\rhd X)_{0}))$
$=\frac{\Phi(y,\operatorname{deg}(Z_{2}\rhd X),Z_{3}))}{\Phi(y,Z_{1},x)}(Y(Z_{2}\rhd X)-y\rhd(Z_{2}\rhd X)).$
By the induction hypothesis, the first term lies in the subalgebra generated by $\mathcal{B}(N)$ and $M$. Since $y$ stabilizes both $\mathcal{B}(N)$ and $M$, the second term also lies in that subalgebra. This completes the induction. Therefore, $R$ is generated by $M\oplus N$, and hence a pre-Nichols algebra of $M\oplus N$.
By the universal property of $\mathcal{B}(M\oplus N)$, there exists a surjective morphism of $\mathbb{N}_{0}$-graded Hopf algebras in $^{G}_{G}\mathcal{YD}^{\Phi}$}:
$\rho:R\rightarrow\mathcal{B}(M\oplus N),\rho\mid_{M\oplus N}=\operatorname{id}.$
This induces a surjective coquasi-Hopf algebra map:
$\rho\#\operatorname{id}:R\#\mathbbm{k}G\rightarrow\mathcal{B}(M\oplus N)\#\mathbbm{k}G.$
Let $K^{\prime}=(\mathcal{B}(M\oplus N)\#\mathbbm{k}G)^{\operatorname{co}\mathcal{B}(N)\#\mathbbm{k}G}$. Then we obtain two bijective coquasi-Hopf algebra maps:
$R\#\mathbbm{k}G\rightarrow K\#(\mathcal{B}(N)\#\mathbbm{k}G),\ \ K^{\prime}\#(\mathcal{B}(N)\#\mathbbm{k}G)\rightarrow\mathcal{B}(M\oplus N)\#\mathbbm{k}G.$
Then the map $\rho\#\operatorname{id}$ thus induces a surjective map of coquasi-Hopf algebras
$\rho^{\prime}:K\#(\mathcal{B}(N)\#\mathbbm{k}G)\rightarrow K^{\prime}\#(\mathcal{B}(N)\#\mathbbm{k}G),\ \ \rho^{\prime}\mid_{(M\oplus N)}=\operatorname{id}.$
Moreover, the following diagram commutes:
$K\#(\mathcal{B}(N)\#\mathbbm{k}G)$$K^{\prime}\#(\mathcal{B}(N)\#\mathbbm{k}G)$$\mathcal{B}(N)\#\mathbbm{k}G$$\mathcal{B}(N)\#\mathbbm{k}G$$\rho^{\prime}$$\varepsilon_{K}\#\operatorname{id}_{\mathcal{B}(N)\#\mathbbm{k}G}$$\varepsilon_{K^{\prime}}\#\operatorname{id}_{\mathcal{B}(N)\#\mathbbm{k}G}$$\operatorname{id}$
since ${\rho\mid}_{M\oplus N}=\operatorname{id}$. As $\mathcal{B}(N)\#\mathbbm{k}G$ acts on $K^{\prime}$ via adjoint operation. Theorem 4.9 implies that $K^{\prime}\cong\mathcal{B}(\operatorname{ad}(\mathcal{B}(N))(M))$. Therefore $\rho^{\prime}$ induces a surjective map
$\phi:K\rightarrow\mathcal{B}(\operatorname{ad}(\mathcal{B}(N))(M))$
in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$} between the right coinvariant subspaces of $\varepsilon_{K}\#\operatorname{id}_{\mathcal{B}(N)\#\mathbbm{k}G}$ and $\varepsilon_{K^{\prime}}\#\operatorname{id}_{\mathcal{B}(N)\#\mathbbm{k}G}$, satisfying $\phi\mid_{M}=\operatorname{id}$. Furthermore, there is a surjective map in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}:
$\phi_{1}:\mathcal{B}(N)\rhd M\rightarrow\operatorname{ad}(\mathcal{B}(N))(M),\ \ \phi_{1}\mid_{M}=\operatorname{id}.$
Since $M=\bigoplus_{i\in I}M_{i}$ is a decomposition into irreducible objects in $^{G}_{G}\mathcal{YD}^{\Phi}$}, each $\operatorname{ad}(\mathcal{B}(N))(M_{i})$ is irreducible in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}. On the other hand, each $\mathcal{B}(N)\rhd M_{i}$ is irreducible as well. Thus $\phi_{1}$ is an isomorphism in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}, and it follows that $\phi$ is an isomorphism of Hopf algebra in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}. Consequently, $\rho\#\operatorname{id}_{\mathbbm{k}G}=\phi\#\operatorname{id}_{\mathcal{A}(N)}$ is an isomorphism of coquasi-Hopf algebras. Therefore,
$\rho=(\rho\#\operatorname{id}_{\mathbbm{k}G})^{\operatorname{co}\mathbbm{k}G}:K\#\mathcal{B}(N)\rightarrow\mathcal{B}(M\oplus N)$
is an isomorphism of Hopf algebras in $^{G}_{G}\mathcal{YD}^{\Phi}$}.
5 Reflection of Nichols algebras
In this section, we develop reflections of Nichols algebras over pointed cosemisimple coquasi-Hopf algebras, which will serve as the cornerstone for our subsequent construction of semi-Cartan graphs. The reflection allows us to systematically relate different Nichols algebra realizations through a well-defined transformation procedure.
5.1 Definition of reflection and basic properties
Let us begin by establishing the basic framework. We still assume $G$ is a finite group, $\Phi$ is a $3$-cocycle on $G$. Fix a positive integer $\theta$ and denote the index set $\mathbb{I}=\{1,2,...,\theta\}$.
Definition 5.1.
Let $\mathcal{F}_{\theta}$ denote the class of all $\theta$-tuples $M=(M_{1},\ldots,M_{\theta})$, where $M_{1},\ldots,M_{\theta}\in{{}^{G}_{G}\mathcal{YD}^{\Phi}}$ are finite-dimensional Yetter-Drinfeld modules. If $M\in\mathcal{F}_{\theta}$, we define
$\mathcal{B}(M):=\mathcal{B}(M_{1}\oplus\cdots\oplus M_{\theta}).$
For tuples $M,M^{\prime}\in\mathcal{F}_{\theta}$, we write $M\cong M^{\prime}$, if $M_{j}\cong M_{j}^{\prime}$ in $^{G}_{G}\mathcal{YD}^{\Phi}$} for each $j$. The isomorphism class of $M\in\mathcal{F}_{\theta}$ is denote by $[M]$.
For $1\leq i\leq\theta$ and $M\in\mathcal{F}_{\theta}$, we say the tuple $M$ admits the $i$-th reflection $R_{i}(M)$ if for all $j\neq i$ there is a natural number $m_{ij}^{M}\geq 0$ such that $(\mathrm{ad}\,M_{i})^{m_{ij}^{M}}(M_{j})$ is a non-zero finite-dimensional subspace of $\mathcal{B}(M)$, and $(\mathrm{ad}\,M_{i})^{m_{ij}^{M}+1}(M_{j})=0$. Assume $M$ admits the $i$-th reflection. Then we set $R_{i}(M)=(V_{1},\ldots,V_{\theta})$, where
$V_{j}=\begin{cases}M_{i}^{*},&\text{if }j=i,\\ (\mathrm{ad}\,M_{i})^{m_{ij}^{M}}(M_{j}),&\text{if }j\neq i.\end{cases}$
Having defined individual reflections, we now extend this notion to sequences of reflections, which will be important for our study of repeated reflections of tuples.
Definition 5.2.
Suppose $M\in\mathcal{F}_{\theta}$, such that $M_{i}$ is irreducible for each $i$. Let $l\in\mathbb{N}_{0}$ and $i_{1},i_{2},...,i_{l}\in\mathbb{I}$,
$(1)$ We say $M$ admits the reflection sequence $(i_{1},i_{2},...,i_{l})$ if $l=0$ or $M$ amdits the $i_{1}$-th reflection and $R_{i_{1}}(M)$ satisfies the reflection sequence $(i_{2},i_{3},...,i_{l})$.
$(2)$ We say $M$ admits all reflection sequence if $M$ admits all reflection sequence $(i_{1},i_{2},...,i_{l})$ for all $l\in\mathbb{N}_{0}$ and $i_{1},i_{2},...,i_{l}\in\mathbb{I}$.
For $M\in\mathcal{F}_{\theta}$ admitting all reflections, we denote
$\mathcal{F}_{\theta}(M)=\{R_{i_{1}}(...(R_{i_{l}}(M))...)\mid l\in\mathbb{N}_{0},\ i_{1},..,i_{l}\in\mathbb{I}\}.$
Such a question arises naturally: how do the irreducibility properties behave under reflections? The following result provides a reassuring answer.
Lemma 5.3.
Suppose $M\in\mathcal{F}_{\theta}$ admits the $i$-th reflection for some $i\in\mathbb{I}$, and $M_{j}$ is irreducible for each $j\in\mathbb{I}$. Then each $R_{i}(M)_{j}$ is irreducible in $^{G}_{G}\mathcal{YD}^{\Phi}$} for $1\leq j\leq\theta$.
Proof.
By definition, $R_{i}(M)$ is defined, thus ${R_{i}(M)}_{i}\cong M_{i}^{*}$ is irreducible since $M_{i}$ is. For $j\neq i$, we observe that $\operatorname{ad}\mathcal{B}(M_{i})(M_{j})=\bigoplus_{n=0}^{m_{ij}}\operatorname{ad}(M_{i})^{n}(M_{j})$. Consider the embedding $\mathbbm{k}G\rightarrow\mathcal{A}(M_{i})$ and the projection $\mathcal{A}(M_{i})\rightarrow\mathbbm{k}G$. The object $\operatorname{ad}(M_{i})^{m_{ij}}(M_{j})$ belongs to $^{G}_{G}\mathcal{YD}^{\Phi}$}. Now suppose $0\neq L\subset\operatorname{ad}(M_{i})^{m_{ij}}(M_{j})$ is a Yetter-Drinfeld submodule over $^{G}_{G}\mathcal{YD}^{\Phi}$}. Let $\langle L\rangle$ be the $\mathcal{A}(M_{i})$-subcomodule of $\operatorname{ad}\mathcal{B}(M_{i})(M_{j})$ generated by $L$, defined explicitly as
$\langle L\rangle:=\{\langle f,X_{-1}\rangle X_{0}\mid X\in L,f\in\operatorname{Hom}(\mathcal{A}(M_{i}),\mathbbm{k})\}.$
We want to prove $\langle L\rangle$ is a Yetter-Drinfeld submodule of $\operatorname{ad}\mathcal{B}(M_{i})(M_{j})$ in $^{A(M_{i})}_{A(M_{i})}\mathcal{YD}$}. To see this, take any homogeneous $A\in\mathcal{A}(N)$. Then for all homogeneous $X\in L$, by Lemma 2.6,
$\operatorname{ad}(A)$ $(\langle f,X_{-1}\rangle X_{0})=\langle f,X_{-1}\rangle\operatorname{ad}(A)(X_{0})$
$=\langle f,s(\mathcal{S}(A_{1}),(\operatorname{ad}(A_{4})(X_{0}))_{-1}A_{6})t(A_{3},X_{-1})\mathcal{S}(A_{2})(\operatorname{ad}(A_{4})(X_{0}))_{-2}A_{5})\rangle(\operatorname{ad}(A_{4})(X_{0}))_{0}.$
Note that $A$ and $X$ are homogeneous, thus if $(\operatorname{ad}(A_{4})(X_{0}))_{-1}\notin G$, then this term equals zero. Therefore
$\operatorname{ad}(A)(\langle f,X_{-1}\rangle X_{0})$
$=\langle f,s(\mathcal{S}(A_{1}),\operatorname{deg}(\operatorname{ad}(A_{4})(X_{0}))A_{6})t(A_{3},X_{-1})\mathcal{S}(A_{2})((\operatorname{ad}(A_{4})(X_{0}))_{-1}A_{5})\rangle(\operatorname{ad}(A_{4})(X_{0}))_{0}.$
Hence $\operatorname{ad}(A)(\langle f,X_{-1}\rangle X_{0})\in\langle L\rangle.$ This ensures the first axiom of a Yetter-Drinfeld module, and the third axiom holds automatically. Since $\operatorname{ad}\mathcal{B}(M_{i})(M_{j})$ is generated by $\operatorname{ad}(M_{i})^{m_{ij}}(M_{j})$ as a Yetter-Drinfeld module and is irreducible in $^{A(M_{i})}_{A(M_{i})}\mathcal{YD}$} by Lemma 4.8 (2). Then $L$ must generate the entire space. This forces $L=\operatorname{ad}(M_{i})^{m_{ij}}(M_{j}),$ establishing the irreducibility.
Lemma 5.4.
Suppose $M\in\mathcal{F}_{\theta}$ and $M$ admits $i$-th reflection for each $i\in\mathbb{I}$. We define $a_{ii}^{M}=2$ for all $1\leq i\leq\theta$ and define $a_{ij}^{M}=-m_{ij}^{M}$. Then $(a_{ij}^{M})_{i,j\in\mathbb{I}}$ is a generalized Cartan matrix.
Proof.
To establish this result, we need only verify: if $1\leq i<j\leq\theta$ such that $a_{ij}^{M}=0$, then $a_{ji}^{M}=0$.
Let $X\in M_{i},Y\in M_{j}$ be homogeneous. A direct computation yields:
$\Delta_{\mathcal{B}(M)}(\operatorname{ad}(X)(Y))$ $=\Delta_{\mathcal{B}(M)}(XY-(\operatorname{ad}(x)(Y))X)$
$=(1\otimes X+X\otimes 1)(1\otimes Y+Y\otimes 1)-(\operatorname{ad}(x)(Y)\otimes 1+1\otimes\operatorname{ad}(x)(Y))(1\otimes X+X\otimes 1)$
$=XY\otimes 1+X\otimes Y+1\otimes XY+\operatorname{ad}(x)(Y)\otimes X$
$-\operatorname{ad}(x)(Y)\otimes X-\operatorname{ad}(x)(Y)X\otimes 1-1\otimes\operatorname{ad}(x)(Y)X-\operatorname{ad}(xyx^{-1})(X)\otimes\operatorname{ad}(x)(Y)$
$=\operatorname{ad}(X)(Y)\otimes 1+1\otimes\operatorname{ad}(X)(Y)+X\otimes Y-c^{2}(X\otimes Y).$
Note that $a_{ij}^{M}=0$ implies $\operatorname{ad}(X)(Y)=0$, from which we deduce $X\otimes Y-c^{2}(X\otimes Y)=0$. This establishes $(\operatorname{id}-c^{2})(M_{i}\otimes M_{j})=0$. Furthermore, we have $(\operatorname{id}-c^{2})c(M_{j}\otimes M_{i})=0$. Since the braiding $c$ is invertible, it follows that $(\operatorname{id}-c^{2})(M_{j}\otimes M_{i})=0.$ Consequently, $\Delta_{\mathcal{B}(M)}(\operatorname{ad}(Y)(X))=0$ in $\mathcal{B}(M_{i}\oplus M_{j})$, for all $X\in M_{i},Y\in M_{j}$, which implies $\operatorname{ad}(Y)(X)=0$. We conclude that $a_{ji}^{M}=0$, as required. ∎
Lemma 5.5.
Suppose $M\cong N$ in $\mathcal{F}_{\theta}$, if $M$ admits the $i$-th reflection for some $i\in\mathbb{I}$, so is $N$. Furthermore, $R_{i}(M)\cong R_{i}(N)$ and $a_{ij}^{M}=a_{ij}^{N}$ for each $j\in\mathbb{I}$.
Proof.
We denote $\phi:M\rightarrow N$ be isomorphism of tuples with each component $\phi_{j}:M_{j}\rightarrow N_{j}$ an isomorphism of Yetter-Drinfeld modules. We first demonstrate that $\phi_{i}$ inducing a unique morphism $\mathcal{B}(\phi_{i}):\mathcal{B}(M_{i})\rightarrow\mathcal{B}(N_{i})$.
The tensor algebra map $T(\phi_{i}):T(M_{i})\rightarrow T(N_{i})$ is a morphism of bialgebras in $^{G}_{G}\mathcal{YD}^{\Phi}$}. Since $T(\phi_{i})$ is a $\mathbb{N}_{0}$-graded coalgebra map, it sends the coideal $I(M_{i})=\bigoplus_{n\geq 2}\operatorname{ker}(\delta_{1^{n}})$ to $I(N_{i})$. Hence $\phi_{i}$ descends to a well-defined morphism $\mathcal{B}(\phi_{i}):\mathcal{B}(M_{i})\rightarrow\mathcal{B}(N_{i})$. This morphism is $\mathbb{N}_{0}$-graded, as its restriction to each homogeneous component $\mathcal{B}(V)(n)$ is induced by $\phi_{i}^{\otimes n}$. Explicitly, for any $n\in\mathbb{N}_{0}$ and $\gamma,\gamma_{1},\ldots,\gamma_{n}\in\mathbb{N}_{0}$ with $\gamma=\gamma_{1}+\cdots+\gamma_{n}$, we have
$\mathcal{B}(\phi_{i})(V(\gamma_{1})\cdots V(\gamma_{n}))=\phi_{i}(V(\gamma_{1}))\cdots\phi_{i}(V(\gamma_{n}))\subseteq\mathcal{B}(W)(\gamma).$
The claim on the surjectivity of $\mathcal{B}(\phi_{i})$ is immediate. While injectivity follows from the equation
$\Delta_{1^{n}}\phi_{i}^{\otimes n}=\phi_{i}^{\otimes n}\Delta_{1^{n}}$
for all $n\in\mathbb{N}_{0}$. We thus obtain an equivalence of categories: ${{}^{A(M_{i})}_{A(M_{i})}\mathcal{YD}}\cong{{}^{A(N_{i})}_{A(N_{i})}\mathcal{YD}}$. For each $j\neq i\in\mathbb{I}$,
$\operatorname{ad}(\mathcal{B}(\phi_{i}))(\phi_{j}):\operatorname{ad}(\mathcal{B}(M_{i}))(M_{j})\cong\operatorname{ad}(\mathcal{B}(N_{i}))(N_{j})$
is an isomorphism of $\mathbb{N}_{0}$-graded objects in $^{A(M_{i})}_{A(M_{i})}\mathcal{YD}$}. In particular, $\operatorname{ad}(M_{i})^{m_{ij}^{M}}(M_{j})\cong\operatorname{ad}(N_{i})^{m_{ij}^{M}}(N_{j})\in{{}^{G}_{G}\mathcal{YD}^{\Phi}}$ and $\operatorname{ad}(N_{i})^{m_{ij}^{M}+1}(N_{j})=0$. This establishes both $a_{ij}^{M}=a_{ij}^{N}$ and $R_{i}(M)\cong R_{i}(N)$. ∎
5.2 Main Theorem
In this subsection, we always fix a tuple $M=(M_{1},\ldots,M_{\theta})\in\mathcal{F}_{\theta}$, with each component $M_{i}$ is irredcuible, $i\in\mathbb{I}$. Furhtermore, we assume $\mathcal{B}(M_{i})$ is finite-dimensional for each $i\in\mathbb{I}$. By Corollary 3.8, we have such a braided tensor equivalence for each $i\in\mathbb{I}$.
$\Omega_{i}:{{}^{A(M_{i})}_{A(M_{i})}\mathcal{YD}}\cong{{}^{A(M_{i}^{*})}_{A(M_{i}^{*})}\mathcal{YD}}.$ (5.1)
The main result of this paper proves that the reflection of Nichols algebra in $^{G}_{G}\mathcal{YD}^{\Phi}$} gives rise to a semi-Cartan graph. The first lemma provides some relations between reflection of $M$ and the braided tensor equivalence $\Omega_{i}$.
Lemma 5.6.
With above assumptions on $M$, and suppose $M$ admits the $i$-th reflection with $R_{i}(M)=(V_{1},...,V_{\theta})$. Then
$\operatorname{ad}(\mathcal{B}(M_{i})(M_{j}))=\bigoplus_{n=0}^{m_{ij}}\operatorname{ad}(M_{i})^{n}(M_{j})$ $,\ \ \ \ V_{j}=\mathcal{F}_{0}(\operatorname{ad}\mathcal{B}(M_{i})(M_{j})),$
$\Omega_{i}(\operatorname{ad}(\mathcal{B}(M_{i})(M_{j})))=\bigoplus_{n=0}^{m_{ij}}\operatorname{ad}(M_{i}^{*})^{n}(V_{j})$ $,\ \ \ \ M_{j}\cong\operatorname{ad}(M_{i}^{*})^{m_{ij}}(V_{j}).$
Proof.
The first equality follows directly from the decomposition:
$\operatorname{ad}\left(\mathcal{B}(M_{i})(M_{j})\right)=\bigoplus_{n\geq 0}\operatorname{ad}(M_{i})^{n}(M_{j}).$
For the second statement, the inclusion $V_{j}\subseteq\mathcal{F}_{0}(\operatorname{ad}\mathcal{B}(M_{i})(M_{j}))$ is immediate. Suppose, for contradiction, that the containment is proper. Then there exists a nonzero $X\in\operatorname{ad}(M_{i})^{l}(M_{j})$ for $0\leq l\leq m_{ij}-1$ such that $X\in\mathcal{F}_{0}(\operatorname{ad}\mathcal{B}(M_{i})(M_{j}))$. The Yetter-Drinfeld submodule generated by $X$ is contained in $\bigoplus_{k=0}^{l}\operatorname{ad}(M_{i})^{k}(M_{j})$. However, since $V_{j}\neq 0$, and $\operatorname{ad}(\mathcal{B}(M_{i})(M_{j}))$ is irreducible in $^{A(M_{i})}_{A(M_{i})}\mathcal{YD}$} by Lemma 4.8(ii), this leads to a contradiction. Hence, $V_{j}=\mathcal{F}_{0}(\operatorname{ad}\mathcal{B}(M_{i})(M_{j}))$.
For the third statement, we consider the category ${{}^{A(M_{i})}_{A(M_{i})}\mathcal{YD}}(\mathbb{Z}\operatorname{-Gr}\mathcal{M}_{\mathbbm{k}})$ by setting $\operatorname{deg}(M_{k})=1$ for $1\leq k\leq n$. Then $\operatorname{ad}(M_{i})^{l}(M_{j})$ has degree $l+1$ for $1\leq l\leq m_{ij}$. By Lemma 3.11, $\Omega_{i}(\operatorname{ad}(M_{i})^{l}(M_{j}))=\operatorname{ad}(M_{i})^{l}(M_{j})$ is $\mathbb{Z}$-graded with $\operatorname{deg}(\Omega_{i}(\operatorname{ad}(M_{i})^{l}(M_{j})))=-1-l$. Moreover,
$\mathcal{F}^{0}\Omega_{i}(\operatorname{ad}\mathcal{B}(M_{i})(M_{j}))=\mathcal{F}_{0}(\operatorname{ad}\mathcal{B}(M_{i})(M_{j}))=\operatorname{ad}(M_{i})^{m_{ij}}(M_{j})=V_{j}.$
Since $\Omega_{i}(\operatorname{ad}(\mathcal{B}(M_{i})(M_{j})))$ is irreducible in $^{A(M_{i}^{*})}_{A(M_{i}^{*})}\mathcal{YD}$}, with the minimal degree part $V_{j}$. By Lemma 4.12,
$\Omega_{i}\left(\operatorname{ad}\left(\mathcal{B}(M_{i})(M_{j})\right)\right)=\operatorname{ad}\left(\mathcal{B}(M_{i}^{*})(V_{j})\right)=\bigoplus_{n=0}^{m_{ij}}\operatorname{ad}(M_{i}^{*})^{n}(V_{j}).$
In particular, we have
$M_{j}=\mathcal{F}^{0}\operatorname{ad}\mathcal{B}(M_{i})(M_{j})=\mathcal{F}_{0}\operatorname{ad}\mathcal{B}(M_{i}^{*})(V_{j})=\operatorname{ad}(M_{i}^{*})^{m_{ij}}(V_{j}).$
The next theorem gives a natural explanation of reflections of tuples of Yetter-Drinfeld modules. Although the proof strategy parallels the Hopf algebra case [rootsys, Theorem 13.4.9], we include the details here for the sake of completeness and the reader’s convenience.
Theorem 5.7.
With above assumptions on $M$. For each $i\in\mathbb{I}$, suppose $M$ admits the $i$-th reflection, then there is an isomorphism of Hopf algebras in $^{G}_{G}\mathcal{YD}^{\Phi}$}:
$\Theta_{i}:\mathcal{B}(R_{i}(M))\cong\Omega_{i}\left(\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}\right)\#\mathcal{B}(M_{i}^{*}).$ (5.2)
Proof.
Let $N=\bigoplus_{j\neq i}M_{j}$, and define $Q=\operatorname{ad}\mathcal{B}(M_{i})(N)$. By Lemma 4.8, we have the decomposition $Q=\bigoplus_{j\neq i}Q_{j}$, where each $Q_{j}=\operatorname{ad}\mathcal{B}(M_{i})(M_{j})$ is irreducible in $^{A(M_{i})}_{A(M_{i})}\mathcal{YD}$}. Applying Theorem 4.9 yields an isomorphism of Hopf algebras:
$\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}\cong\mathcal{B}(Q).$
Now observe that the functor $\Omega_{i}$ sends Nichols algebras to Nichols algebras by Lemma 2.16. Therefore,
$\Omega_{i}\left(\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}\right)\cong\Omega_{i}(\mathcal{B}(Q))\cong\mathcal{B}(\Omega_{i}(Q)).$
Hence, in the category $^{G}_{G}\mathcal{YD}^{\Phi}$},
$\Omega_{i}\left(\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}\right)\#\mathcal{B}(M_{i}^{*})\cong\mathcal{B}(\Omega_{i}(Q))\#\mathcal{B}(M_{i}^{*}).$
Since $Q$ is a direct sum of irreducible objects, it is semisimple. The equivalence $\Omega_{i}$ preserves semisimplicity, so $\Omega_{i}(Q)$ is also semisimple in ${{}^{A(M_{i}^{*})}_{A(M_{i}^{*})}\mathcal{YD}}(\mathbb{Z}\operatorname{-Gr}\mathcal{M}_{\mathbbm{k}})$. By Proposition 4.13, we obtain
$\mathcal{B}\left(\Omega_{i}\left(Q\right)\right)\#\mathcal{B}(M_{i}^{*})\cong\mathcal{B}\left(\mathcal{F}^{0}(\Omega_{i}(Q))\oplus M_{i}^{*}\right).$
We have already shown that $\mathcal{F}^{0}(\Omega_{i}(Q))=\bigoplus_{j\neq i}V_{j}$. Thus
$\mathcal{B}(\Omega_{i}(Q))\#\mathcal{B}(M_{i}^{*})\cong\mathcal{B}\left(\bigoplus_{j\neq i}V_{j}\oplus M_{i}^{*}\right)\cong\mathcal{B}(R_{i}(M)).$
Combining the above isomorphisms, we conclude that in $^{G}_{G}\mathcal{YD}^{\Phi}$}, there is an isomorphism of Hopf algebras:
$\Theta_{i}:\mathcal{B}(R_{i}(M))\cong\Omega_{i}\left(\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}\right)\#\mathcal{B}(M_{i}^{*}).$
Corollary 5.8.
With above assumptions on $M$. For each $i\in\mathbb{I}$, suppose $M$ admits $i$-th reflection, then $R_{i}(M)$ admits $i$-th reflection. Furthermore, we have:
$R_{i}^{2}(M)\cong M,$ (5.3)
and
$a_{ij}^{M}=a_{ij}^{R_{i}(M)},$ (5.4)
for all $1\leq j\leq\theta$.
Proof.
By Theorem 5.7, $R_{i}(M)=(V_{1},...,M_{i}^{*},...,V_{\theta}).$ Note that
$\operatorname{ad}\mathcal{B}(M_{i}^{*})(V_{j})=\Omega_{i}\left(\operatorname{ad}\mathcal{B}(M_{i})(M_{j})\right).$
It follows that
$\operatorname{ad}\mathcal{B}(M_{i}^{*})^{1-a_{ij}^{M}}(V_{j})=0,$
which implies that $R_{i}(M)$ admits $i$-th reflection and $a_{ij}^{M}=a_{ij}^{R_{i}(M)}$ for all $j\neq i$. It is clear that the $i$-th position of $R_{i}^{2}(M)$ is isomorphic to $M_{i}$. For $j\neq i$,
$\mathcal{F}_{0}\left(\operatorname{ad}\mathcal{B}(M_{i}^{*})(V_{j})\right)=\mathcal{F}_{0}\left(\Omega_{i}\left(\operatorname{ad}\left(\mathcal{B}\left(M_{i}\right)(M_{j})\right)\right)\right)=\mathcal{F}^{0}(\operatorname{ad}(\mathcal{B}(M_{i}))(M_{j}))=M_{j}.$
Therefore $R_{i}^{2}(M)_{j}\cong M_{j}$ for all $j\neq i$, and hence $R_{i}^{2}(M)\cong M$. ∎
Theorem 5.9.
Let $M=(M_{1},\ldots,M_{\theta})\in\mathcal{F}_{\theta}$ such that each component $M_{i}$ is irreducible. Suppose $M$ admits all reflections. We define the set
$\mathcal{X}=\{[P]\mid P\in\mathcal{F}_{\theta}(M)\},$
and the map
$r:\mathbb{I}\times\mathcal{X}\rightarrow\mathcal{X},\ i\times[X]\mapsto[R_{i}(X)].$
Furthermore, we assume each $\mathcal{B}(P_{i})$ is finite-dimensional for $P\in\mathcal{X}$ and $i\in\mathbb{I}$, then
$\mathcal{G}(M)=(\mathbb{I},\mathcal{X},r,(A^{X})_{X\in\mathcal{X}}),$
where $A^{[X]}=(a_{ij}^{X})_{i,j\in\mathbb{I}}$ for all $X]\in\mathcal{X}$, is a semi-Cartan graph.
Proof.
It follows by definition of semi-Cartan graph and Corollary 5.8. ∎
6 Applications
Building upon the framework developed in the preceding sections, we now present some applications of the semi-Cartan graph theory for Nichols algebras over certain coquasi-Hopf algebras. We establish several criteria that connect the finite-dimensionality of a Nichols algebra to the finiteness and structure of its associated semi-Cartan graph and Weyl groupoid. Then we apply this theory to provide a new proof for the key result in [huang2024classification].
6.1 A criterion for finite-dimensional Nichols algebras
We still work on the setting of $^{G}_{G}\mathcal{YD}^{\Phi}$}, where $G$ is a finite group and $\Phi$ is a $3$-cocycle on $G$. Now fix a positive integer $\theta$ and denote the index set $\mathbb{I}=\{1,2,...,\theta\}$. Let $M=(M_{1},\ldots,M_{\theta})\in\mathcal{F}_{\theta}$. We endow a $\mathbb{Z}^{\theta}$-grading on $\mathcal{B}(M)$ by setting
$\operatorname{deg}(M_{i})=\alpha_{i},\ 1\leq i\leq\theta,$
where $\alpha_{i}$, $1\leq i\leq\theta$ are the standard basis of $\mathbb{Z}^{\theta}$. For each $\beta\in\mathbb{Z}^{\theta}$, define the homogeneous component $\mathcal{B}(M)_{\beta}=\{x\in\mathcal{B}(M)\mid\operatorname{deg}(x)=\beta\}$, and the support of $\mathcal{B}(M)$ as
$\operatorname{Supp}\mathcal{B}(M):=\{\beta\in\mathbb{Z}^{\theta}\mid\mathcal{B}(M)_{\beta}\neq 0\}.$
Lemma 6.1.
For each $i\in\mathbb{I}$, suppose $M$ admits $i$-th reflection,
$\operatorname{Supp}\mathcal{B}(M)\cup\operatorname{Supp}\mathcal{B}(M^{*})$ $=\operatorname{Supp}\mathcal{B}(R_{i}(M))\cup\operatorname{Supp}\mathcal{B}(R_{i}(M)^{*}),$ (6.1)
$s_{i}^{M}(\operatorname{Supp}\mathcal{B}(M)\cup\operatorname{Supp}\mathcal{B}(M^{*}))$ $=\operatorname{Supp}\mathcal{B}(R_{i}(M))\cup\operatorname{Supp}\mathcal{B}(R_{i}(M)^{*}).$ (6.2)
Proof.
If we regard $\mathcal{B}(M)$ as an $\mathbb{N}_{0}$-graded object in $^{G}_{G}\mathcal{YD}^{\Phi}$}. By [defofnichols, Lemma 2.6] the graded dual of $\mathcal{B}(M)$ is isomorphic to $\mathcal{B}(M^{*})$. We thus obtain an isomorphism of $\mathbb{N}_{0}$-graded objects in $^{G}_{G}\mathcal{YD}^{\Phi}$}:
$\mathcal{B}(M)\otimes\mathcal{B}(M^{*})\cong K_{i}\otimes\mathcal{B}(M_{i})\otimes K_{i}^{\operatorname{gr-dual}}\otimes\mathcal{B}(M_{i}^{*}).$
Similarly, for the reflected tuple, we have
$\mathcal{B}(R_{i}(M))\otimes\mathcal{B}(R_{i}(M)^{*})\cong K_{i}\otimes\mathcal{B}(M_{i}^{*})\otimes K_{i}^{\operatorname{gr-dual}}\otimes\mathcal{B}(M_{i}).$
Since $A\otimes B\cong B\otimes A$ as $\mathbb{Z}^{\theta}$-graded vector spaces. the supports on both sides coincide, which establishes the first claim.
The second claim follows directly from the definition of $s_{i}^{M}$. ∎
Next, we show that reflections preserve the dimension of Nichols algebras.
Lemma 6.2.
Suppose $\operatorname{dim}\mathcal{B}(M)<\infty$, then $M$ gives rise to a semi-Cartan graph, and $\operatorname{dim}\mathcal{B}(R_{i}(M))=\operatorname{dim}\mathcal{B}(M)$ for each $i\in\mathbb{I}$.
Proof.
For each $i\in\mathbb{I}$, as vector spaces, we have the isomorphism $\mathcal{B}(M)\cong K_{i}\otimes\mathcal{B}(M_{i})$. Therefore $K_{i}$ is finite-dimensional, and $\operatorname{ad}(\mathcal{B}(M_{i}))(M_{j})$ is finite-dimensional for each $j\in\mathbb{I}$. It follows by definition that $M$ admits the $i$-th reflection. Applying the equivalence $\Omega_{i}$, we obtain $\Omega_{i}(K_{i})=K_{i}\in{{}^{A(M_{i}^{*})}_{A(M_{i}^{*})}\mathcal{YD}}$ and $\mathcal{B}(R_{i}(M))\cong K_{i}\otimes\mathcal{B}(M_{i}^{*})$ as vector space. Since $\mathcal{B}(M_{i})$ is finite-dimensional, we have $\operatorname{dim}\mathcal{B}(M_{i})=\operatorname{dim}\mathcal{B}(M_{i}^{*})$. Therefore $\operatorname{dim}\mathcal{B}(R_{i}(M))=\operatorname{dim}\mathcal{B}(M)$ for each $i\in\mathbb{I}$. By iterated reflections, $M$ admits all reflections and gives rise to a semi-Cartan graph. ∎
Before we state the next lemma, we recall the definition of a finite semi-Cartan graph.
Definition 6.3.
Let $\mathcal{G}=(\mathbb{I},\mathcal{X},r,(A^{X})_{X\in\mathcal{X}})$ be a semi-Cartan graph. For all $X\in\mathcal{X}$, the set
$\Delta^{X,\operatorname{re}}=\{\omega(\alpha_{i})\in\mathbb{Z}^{\mathbb{I}}\mid\omega\in\operatorname{Hom}(\mathcal{W}(\mathcal{G}),X),i\in\mathbb{I}\}$
is called the set of real roots of $\mathcal{G}$ at $X$. We call the semi-Cartan graph $\mathcal{G}$ is finite, if $\Delta^{X,\operatorname{re}}$ is finite for all $X\in\mathcal{X}$.
We now relate the finiteness of the Nichols algebra to the structure of the associated semi-Cartan graph.
Lemma 6.4.
Suppose $\mathcal{B}(M)$ is finite-dimensional, then the semi-Cartan graph $\mathcal{G}(M)$ is finite.
Proof.
Let $W=M_{1}\oplus\cdots\oplus M_{\theta}$. Clearly, the homogeneous degrees of elements in $W$ lie in the union $W\subseteq\operatorname{Supp}\mathcal{B}(M)\cup\operatorname{Supp}\mathcal{B}(M^{*})$. For $1\leq j\leq\theta$, we apply equations (6.1),(6.2), together with the fact that $R_{i}^{2}(M)\cong M$ to obtain
$s_{i}^{R_{i}(M)}(\alpha_{j})$ $\subseteq s_{i}^{R_{i}(M)}(\operatorname{Supp}\mathcal{B}(M)\cup\operatorname{Supp}\mathcal{B}(M^{*}))$
$=s_{i}^{R_{i}(M)}(\operatorname{Supp}\mathcal{B}(R_{i}(M))\cup\operatorname{Supp}\mathcal{B}(R_{i}(M^{*})))$
$=\operatorname{Supp}\mathcal{B}(M)\cup\operatorname{Supp}\mathcal{B}(M^{*}).$
We recall that every element $\omega\in\operatorname{Hom}(\mathcal{W}(\mathcal{G}),M)$ can be expressed as an iteration of simple reflections. It follows that Thus $\Delta^{M\operatorname{re}}\subseteq\operatorname{Supp}\mathcal{B}(M)\cup\operatorname{Supp}\mathcal{B}(M^{*})$. Now, since $\mathcal{B}(M)$ is finite-dimensional, the set $\operatorname{Supp}\mathcal{B}(M)$ is finite. Moreover, as vector spaces, $\mathcal{B}(M)\cong\mathcal{B}(M^{*})^{\mathrm{gr\text{-}dual}}$, so $\operatorname{Supp}\mathcal{B}(M^{*})$ is also finite. Combining these observations, we conclude that $\Delta^{M,\operatorname{re}}$ is finite. The same reasoning applies to any $N\in X$: we have
$\Delta^{N,\operatorname{re}}\subseteq\operatorname{Supp}\mathcal{B}(N)\cup\operatorname{Supp}\mathcal{B}(N^{*}).$
Then $\Delta^{N,\operatorname{re}}$ is again finite by Lemma 6.2. Therefore, $\mathcal{G}(M)$ is a finite semi-Cartan graph. ∎
Lemma 6.5.
[HS10, Lemma 5.1] Suppose $\mathcal{G}=(\mathbb{I},\mathcal{X},r,(A^{X})_{X\in\mathcal{X}})$ is a semi-Cartan graph. Then $\mathcal{G}$ is finite semi-Cartan graph if and only if the Weyl groupoid $\mathcal{W}(\mathcal{G})$ is finite.
Definition 6.6.
Suppose $\mathcal{G}=(\mathbb{I},\mathcal{X},r,(A^{X})_{X\in\mathcal{X}})$ is a semi-Cartan graph. If $A^{X}=A^{Y}$ for all $X,Y\in\mathcal{X}$, then $\mathcal{G}$ is called a standard semi-Cartan graph.
The proof of next theorem is similar to the Hopf algebra situations, we write down here for completeness.
Theorem 6.7.
Let $M=(M_{1},\ldots,M_{\theta})\in\mathcal{F}_{\theta}$, where each $M_{i}$ is irreducible for all $i\in\mathbb{I}$. If $\operatorname{dim}\mathcal{B}(M)<\infty$ and $\mathcal{G}(M)$ is standard, then $A^{M}$ must be a finite Cartan matrix.
Proof.
Since $\operatorname{dim}\mathcal{B}(M)<\infty$, Lemma 6.4 implies that $\mathcal{G}(M)$ is a finite Semi-Cartan graph. By Lemma 6.5, $\mathcal{W}(\mathcal{G}(M))$ is finite, in particular, $\operatorname{Hom}(\mathcal{W}(\mathcal{G}(M)))$ is finite.
Let $W(A^{M})$ be the Weyl group of the Kac-Moody Lie algebra of the Cartan matrix $A^{M}$. Note that
$W(A^{M})=\langle s_{i}^{M}\in\operatorname{Aut}(\mathbb{Z}^{\theta})\mid 1\leq i\leq\theta\rangle.$
Since $\mathcal{G}(M)$ is standard, we know that
$\operatorname{Hom}(\mathcal{W}(\mathcal{G}(M)))\longrightarrow W(A^{M}),\ \ (Y,s,X)\mapsto s$
is well-defined and surjective. Therefore $W(A^{M})$ is finite, which implies that $A^{M}$ is of finite type. ∎
6.2 A class of infinite-dimensional Nichols algebras
The purpose of this section is to give a new proof of Proposition 4.1 in huang2024classification through applying our previous observations.
Let $G=\mathbb{Z}_{m_{1}}\times\mathbb{Z}_{m_{2}}\times\mathbb{Z}_{m_{3}}=\langle g_{1}\rangle\times\langle g_{2}\rangle\times\langle g_{3}\rangle$ with $m_{1}\mid m_{2}$, $m_{2}\mid m_{3}$, and
$\Phi\left(g_{1}^{i_{1}}g_{2}^{i_{2}}g_{3}^{i_{3}},g_{1}^{j_{1}}g_{2}^{j_{2}}g_{3}^{j_{3}},g_{1}^{k_{1}}g_{2}^{k_{2}}g_{3}^{k_{3}}\right)=(-1)^{k_{1}j_{2}i_{3}}$ (6.3)
be a $3$-cocycle on $G$, where $0\leq i_{l}<m_{1}$, $0\leq j_{l}<m_{2}$, $0\leq k_{l}<m_{3}$, $1\leq l\leq 3$.
Let $V_{1},V_{2},V_{3}\in{{}^{G}_{G}\mathcal{YD}^{\Phi}}$ be simple modules such that $\operatorname{deg}(V_{1})=g_{1}$, $\operatorname{deg}(V_{2})=g_{2}$, $\operatorname{deg}(V_{3})=g_{3}$ and $\operatorname{dim}(V_{1})=\operatorname{dim}(V_{2})=\operatorname{dim}(V_{3})=2$. Moreover, $V_{1}=\left\{X_{1}^{\prime},X_{2}^{\prime}\right\},V_{2}=\left\{Y_{1}^{\prime},Y_{2}^{\prime}\right\},V_{3}=\left\{Z_{1}^{\prime},Z_{2}^{\prime}\right\}$ can be assumed satisfying the following equations:
[…]
Remark 6.8.
(i) Note that $m_{1},m_{2},m_{3}$ must be even. Indeed, from the relation $X_{i}=e^{m_{1}}\rhd X_{i}=(-1)^{m_{1}}X_{i}$, it follows that $m_{1}$ is even. Similarly $m_{2},m_{3}$ are even.
(ii) $\mathcal{B}(V_{1})$, $\mathcal{B}(V_{2})$, $\mathcal{B}(V_{3})$ are finite-dimensional Nichols algebras by [huang2024classification, Proposition 2.12]. Moreover, $\mathcal{B}(V_{1}\oplus V_{2})$, $\mathcal{B}(V_{1}\oplus V_{3})$, $\mathcal{B}(V_{2}\oplus V_{3})$ are finite-dimensional by [nondiagonal Proposition 5.1].
(iii) The requirement $\operatorname{dim}(V_{1})=\operatorname{dim}(V_{2})=\operatorname{dim}(V_{3})=2$ is in fact necessary, otherwise the corresponding Nichols algebra must be infinite-dimensional.
Theorem 6.9.
[huang2024classification, Proposition 4.1] With above notations, $\mathcal{B}(V_{1}\oplus V_{2}\oplus V_{3})$ is infinite-dimensional.
It should be emphasized that the above theorem plays the key role in that paper and the original proof relays on heavy computations. The purpose of the following context is to give a new proof of this theorem through applying reflection theory of Nichols algebras.
To simplify the proof, we proceed with the following reduction. Let $H=\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}=\langle h_{1}\rangle\times\langle h_{2}\rangle\times\langle h_{3}\rangle$, and
$\phi\left(h_{1}^{i_{1}}h_{2}^{i_{2}}h_{3}^{i_{3}},h_{1}^{j_{1}}h_{2}^{j_{2}}h_{3}^{j_{3}},h_{1}^{k_{1}}h_{2}^{k_{2}}h_{3}^{k_{3}}\right)=(-1)^{k_{1}j_{2}i_{3}},$
be a $3$-cocycle on $H$, where $0\leq i_{l},j_{l},k_{l}\leq 1$, $1\leq l\leq 3$. By huang2024classification Lemma 3.5, we may give a complete list of irreducible Yetter-Drinfeld modules. In particular, restricting to those irreducible Yetter–Drinfeld modules that generate finite-dimensional Nichols algebras, the isomorphism classes are given by the set
$S=\{[^{h}W]\mid\ ^{h}W\ \text{is irreducible},\ h\in\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}-\{1,h_{1}h_{2}h_{3}\}\}.$ (6.4)
There are exactly six isomorphism classes. Here $^{h}W$ represents the simple Yetter–Drinfeld module has comodule structure
$^{h}W:=\{w\in W\mid\delta_{W}(w)=h\otimes w\}.$
We now choose explicit representatives of $S$. For $1\leq i\leq 6$, let $W_{i}\in S$ be pairwise non-isomorphic simple modules such that $\operatorname{dim}(W_{i})=2$, and $\operatorname{deg}(W_{1})=h_{1}$, $\operatorname{deg}(W_{2})=h_{2}$, $\operatorname{deg}(W_{3})=h_{3}$, $\operatorname{deg}(W_{4})=h_{1}h_{2}$, $\operatorname{deg}(W_{5})=h_{1}h_{3}$, $\operatorname{deg}(W_{6})=h_{2}h_{3}$. Moreover, we may take
$W_{1}$ $=\{X_{1},X_{2}\},$ $W_{2}$ $=\{Y_{1},Y_{2}\},$ $W_{3}$ $=\{Z_{1},Z_{2}\},$
$W_{4}$ $=\{R_{1},R_{2}\},$ $W_{5}$ $=\{S_{1},S_{2}\},$ $W_{6}$ $=\{T_{1},T_{2}\},$
satisfying the following relations:
[…]
Remark 6.10.
The category ${{}^{H}_{H}\mathcal{YD}^{\phi}}_{\operatorname{fd}}$ is rigid. It is straightforward to verify that for each $1\leq i\leq 6$, $W_{i}^{*}$ does exist with $W_{i}^{*}\cong W_{i}$. This implies $W_{i}^{*}\in S$ as well. Moreover, $\mathcal{B}(W_{i})\cong\mathcal{B}(W_{i}^{*})$ as Hopf algebras in $^{H}_{H}\mathcal{YD}$}.
The application of semi-Cartan graph theory is the following result.
Theorem 6.11.
Let $W=(W_{1},W_{2},W_{3})$ be the $3$-tuple, then $W$ admits all reflections and $\mathcal{G}(W)$ is a standard semi-Cartan graph.
To establish this result, we first prove the following key lemma:
Lemma 6.12.
For all $1\leq i\neq j\leq 6$, if $\operatorname{deg}(W_{i})\cdot\operatorname{deg}(W_{j})\neq 1,h_{1}h_{2}h_{3}$, then $\operatorname{ad}(W_{i})^{2}(W_{j})=0.$ Moreover, $\operatorname{ad}(W_{i})(W_{j})\in S$, and as Yetter-Drinfeld module in $^{H}_{H}\mathcal{YD}^{\phi}$}, we have
$\operatorname{ad}(W_{i})(W_{j})\cong\operatorname{ad}(W_{j})(W_{i}).$ (6.5)
We list the following isomorphism for further use
$\operatorname{ad}(W_{1})(W_{2})\cong W_{4},\operatorname{ad}(W_{1})(W_{3})\cong W_{5},\operatorname{ad}(W_{1})(W_{4})\cong W_{2},$ (6.6)
$\operatorname{ad}(W_{1})(W_{5})\cong W_{3},\operatorname{ad}(W_{2})(W_{3})\cong W_{6},\operatorname{ad}(W_{2})(W_{4})\cong W_{1},$
$\operatorname{ad}(W_{2})(W_{6})\cong W_{3},\operatorname{ad}(W_{3})(W_{5})\cong W_{2},\operatorname{ad}(W_{3})(W_{6})\cong W_{2}$
$\operatorname{ad}(W_{4})(W_{5})\cong W_{6},\operatorname{ad}(W_{4})(W_{6})\cong W_{5},\operatorname{ad}(W_{5})(W_{6})\cong W_{4}.$
Proof.
This proof involves extensive but analogous computations. We begin by establishing that $\operatorname{ad}(W_{1})^{2}(W_{2})=0.$ To this end, we compute the coproducts:
$\Delta\left(\operatorname{ad}(X_{1})\left(Y_{1}\right)\right)=1\otimes\operatorname{ad}({X_{1}})\left(Y_{1}\right)+\operatorname{ad}({X_{1}})\left(Y_{1}\right)\otimes 1+X_{1}\otimes\left(Y_{1}+Y_{2}\right),$
$\Delta\left(\operatorname{ad}({X_{1}})\left(Y_{2}\right)\right)=1\otimes\operatorname{ad}({X_{1}})\left(Y_{2}\right)+\operatorname{ad}({X_{1}})\left(Y_{2}\right)\otimes 1+X_{1}\otimes\left(Y_{1}+Y_{2}\right),$
$\Delta\left(\operatorname{ad}({X_{2}})\left(Y_{1}\right)\right)=1\otimes\operatorname{ad}({X_{2}})\left(Y_{1}\right)+\operatorname{ad}({X_{2}})\left(Y_{1}\right)\otimes 1+X_{2}\otimes\left(Y_{1}-Y_{2}\right),$
$\Delta\left(\operatorname{ad}({X_{2}})\left(Y_{2}\right)\right)=1\otimes\operatorname{ad}({X_{2}})\left(Y_{2}\right)+\operatorname{ad}({X_{2}})\left(Y_{2}\right)\otimes 1+X_{2}\otimes\left(Y_{2}-Y_{1}\right).$
From these we deduce the relations:
$\operatorname{ad}({X_{1}})\left(Y_{1}\right)-\operatorname{ad}({X_{1}})\left(Y_{2}\right)=0,$ (6.7)
$\operatorname{ad}({X_{2}})\left(Y_{1}\right)+\operatorname{ad}({X_{2}})\left(Y_{2}\right)=0.$
Therefore $\operatorname{ad}(W_{1})(W_{2})=\operatorname{span}\{\operatorname{ad}(X_{1})(Y_{1}),\operatorname{ad}(X_{2})(Y_{1})\}$. Now since $X_{1}^{2}=0$, we have
$\operatorname{ad}(X_{1})^{2}\left(Y_{1}\right)=X_{1}(X_{1}Y_{1}-Y_{2}X_{1})-h_{1}\rhd((X_{1}Y_{1}-Y_{2}X_{1}))X_{1}=0$
which implies $\operatorname{ad}(X_{1})^{2}\left(Y_{1}\right)=\operatorname{ad}(X_{1})^{2}\left(Y_{2}\right)=0$. Meanwhile,
$\operatorname{ad}({X_{2}})\left(\operatorname{ad}({X_{1}})\left(Y_{1}\right)\right)=\operatorname{ad}({X_{2}})\left(\operatorname{ad}({X_{1}})\left(Y_{2}\right)\right)=0.$
Since
$\begin{aligned} \Delta(\operatorname{ad}({X_{2}})\left(\operatorname{ad}({X_{1}})\left(Y_{1}\right)\right))&=(1\otimes X_{2}+X_{2}\otimes 1)(1\otimes\operatorname{ad}({X_{1}})\left(Y_{1}\right)+\operatorname{ad}({X_{1}})\left(Y_{1}\right)\otimes 1+X_{1}\otimes\left(Y_{1}+Y_{2}\right)\\ &-(1\otimes\operatorname{ad}({X_{1}})\left(Y_{1}\right)+\operatorname{ad}({X_{1}})\left(Y_{1}\right)\otimes 1+X_{1}\otimes\left(Y_{1}+Y_{2}\right))(1\otimes X_{2}+X_{2}\otimes 1)\\ &=1\otimes\operatorname{ad}({X_{2}})\left(\operatorname{ad}({X_{1}})\left(Y_{1}\right)\right)+\operatorname{ad}({X_{2}})\left(\operatorname{ad}({X_{1}})\left(Y_{1}\right)\right)\otimes 1\end{aligned},$
where we have used equation (6.7). A similar argument shows $\operatorname{ad}(X_{2})^{2}\left(Y_{1}\right)=0$ and
$\operatorname{ad}({X_{1}})\left(\operatorname{ad}({X_{2}})\left(Y_{1}\right)\right)=0$. We conclude that $\operatorname{ad}(W_{1})^{2}(W_{2})=0$.
By Lemma 5.3, the object $\operatorname{ad}(W_{1})(W_{2})\in{{}^{H}_{H}\mathcal{YD}^{\phi}}$ is simple. Since $\operatorname{deg}(\operatorname{ad}(W_{1})(W_{2}))=h_{1}h_{2}$, $\operatorname{ad}(W_{1})(W_{2})\cong W_{4}$. Furthermore, by proof of Lemma 5.5, we have an isomorphism of $\mathbb{N}_{0}$-graded objects $\operatorname{ad}\mathcal{B}(W_{1})(W_{4})\cong\operatorname{ad}\mathcal{B}(W_{1})(\operatorname{ad}(W_{1})(W_{2}))$. Since $\operatorname{ad}(W_{1})^{2}(W_{2})=0$, we have $\operatorname{ad}(W_{1})^{2}(W_{4})=0$ since it has same degree as $\operatorname{ad}(W_{1})^{2}(W_{2})$ in the standard $\mathbb{N}_{0}$-grading. Therefore, $\operatorname{ad}(W_{1})(W_{4})$ is irreducible and one readily verifies that $\operatorname{ad}(W_{1})(W_{4})\cong W_{2}$ in $^{H}_{H}\mathcal{YD}^{\phi}$}.
Analogous computations yield:
$\operatorname{ad}(W_{2})^{2}(W_{1})=\operatorname{ad}(W_{4})^{2}(W_{1})=0.$
These identities further imply:
$\operatorname{ad}(W_{2})(W_{1})\cong W_{4},\ \operatorname{ad}(W_{4})(W_{1})\cong W_{2},$
and
$\operatorname{ad}(W_{2})^{2}(W_{4})=\operatorname{ad}(W_{4})^{2}(W_{2})=0,$
$\operatorname{ad}(W_{2})(W_{4})\cong W_{1},\ \operatorname{ad}(W_{4})(W_{2})\cong W_{1}.$
We now proceed to prove $\operatorname{ad}(W_{1})^{2}(W_{3})=0$:
$\Delta\left(\operatorname{ad}({X_{1}})\left(Z_{1}\right)\right)=1\otimes\operatorname{ad}({X_{1}})\left(Z_{1}\right)+\operatorname{ad}({X_{1}})\left(Z_{1}\right)\otimes 1+X_{1}\otimes Z_{1}+X_{2}\otimes Z_{2},$
$\Delta\left(\operatorname{ad}({X_{1}})\left(Z_{2}\right)\right)=1\otimes\operatorname{ad}({X_{1}})\left(Z_{2}\right)+\operatorname{ad}({X_{1}})\left(Z_{2}\right)\otimes 1+X_{1}\otimes Z_{2}+X_{2}\otimes Z_{1},$
$\Delta\left(\operatorname{ad}({X_{2}})\left(Z_{1}\right)\right)=1\otimes\operatorname{ad}({X_{2}})\left(Z_{1}\right)+\operatorname{ad}({X_{2}})\left(Z_{1}\right)\otimes 1+X_{2}\otimes Z_{1}+X_{1}\otimes Z_{2},$
$\Delta\left(\operatorname{ad}({X_{2}})\left(Z_{2}\right)\right)=1\otimes\operatorname{ad}({X_{2}})\left(Z_{2}\right)+\operatorname{ad}({X_{2}})\left(Z_{2}\right)\otimes 1+X_{2}\otimes Z_{2}+X_{1}\otimes Z_{1}.$
It is straightforward to observe that:
$\operatorname{ad}({X_{1}})\left(Z_{1}\right)-\operatorname{ad}({X_{2}})\left(Z_{2}\right)=0,$ (6.8)
$\operatorname{ad}({X_{2}})\left(Z_{1}\right)+\operatorname{ad}({X_{1}})\left(Z_{2}\right)=0.$
Hence $\operatorname{ad}(W_{1})(W_{3})=\operatorname{span}\{\operatorname{ad}(X_{1})(Z_{1}),\operatorname{ad}(X_{2})(Z_{1})\}$. By $X_{1}^{2}=X_{2}^{2}=0$,
$\operatorname{ad}(X_{1})^{2}\left(Z_{1}\right)$ $=X_{1}(X_{1}Z_{1}-Z_{2}X_{1})-h_{1}\rhd(X_{1}Z_{1}-Z_{2}X_{1})X_{1}=0,$
$\operatorname{ad}(X_{1})^{2}(Z_{2})$ $=X_{1}(X_{1}Z_{2}-Z_{1}X_{1})-h_{1}\rhd(X_{1}Z_{2}-Z_{1}X_{1})X_{1}=0,$
$\operatorname{ad}(X_{2})^{2}\left(Z_{1}\right)$ $=X_{2}(X_{2}Z_{1}-Z_{2}X_{2})-h_{1}\rhd(X_{2}Z_{1}-Z_{2}X_{2})X_{2}=0,$
$\operatorname{ad}(X_{2})^{2}(Z_{2})$ $=X_{2}(X_{2}Z_{2}-Z_{1}X_{2})-h_{1}\rhd(X_{2}Z_{2}-Z_{1}X_{2})X_{2}=0.$
This combines (6.8) implies $\operatorname{ad}(W_{1})^{2}(W_{3})=0$. As in the first case, we can derive the following identities:
$\operatorname{ad}(W_{3})^{2}(W_{1})=$ $\operatorname{ad}(W_{1})^{2}(W_{5})=\operatorname{ad}(W_{3})^{2}(W_{5})=\operatorname{ad}(W_{5})^{2}(W_{1})=\operatorname{ad}(W_{5})^{2}(W_{3})=0.$
$\operatorname{ad}(W_{1})(W_{3})\cong\operatorname{ad}(W_{3})(W_{1})\cong W_{5},$
$\operatorname{ad}(W_{1})(W_{5})\cong\operatorname{ad}(W_{5})(W_{1})\cong W_{3},$
$\operatorname{ad}(W_{3})(W_{5})\cong\operatorname{ad}(W_{5})(W_{3})\cong W_{1}.$
Now we turn to proving $\operatorname{ad}(W_{2})^{2}(W_{3})=0$:
$\Delta\left(\operatorname{ad}({Y_{1}})\left(Z_{1}\right)\right)=1\otimes\operatorname{ad}({Y_{1}})\left(Z_{1}\right)+\operatorname{ad}({Y_{1}})\left(Z_{1}\right)\otimes 1,$
$\Delta\left(\operatorname{ad}({Y_{1}})\left(Z_{2}\right)\right)=1\otimes\operatorname{ad}({Y_{1}})\left(Z_{2}\right)+\operatorname{ad}({Y_{1}})\left(Z_{2}\right)\otimes 1+2Y_{1}\otimes Z_{2},$
$\Delta\left(\operatorname{ad}({Y_{2}})\left(Z_{1}\right)\right)=1\otimes\operatorname{ad}({Y_{2}})\left(Z_{1}\right)+\operatorname{ad}({Y_{2}})\left(Z_{1}\right)\otimes 1+2Y_{2}\otimes Z_{1},$
$\Delta\left(\operatorname{ad}({Y_{2}})\left(Z_{2}\right)\right)=1\otimes\operatorname{ad}({Y_{2}})\left(Z_{2}\right)+\operatorname{ad}({Y_{2}})\left(Z_{2}\right)\otimes 1.$
Then $\operatorname{ad}(Y_{1})(Z_{1})=\operatorname{ad}(Y_{2})(Z_{2})=0$ and $\operatorname{ad}(W_{2})(W_{3})=\{\operatorname{ad}(Y_{1})(Z_{2}),\operatorname{ad}(Y_{2})(Z_{1})\}.$ By $Y_{1}^{2}=Y_{2}^{2}=0$,
$\operatorname{ad}(Y_{1})^{2}(Z_{2})$ $=Y_{1}(Y_{1}Z_{2}+Z_{2}Y_{1})-h_{2}\rhd(Y_{1}Z_{2}+Z_{2}Y_{1})Y_{1}=0,$
$\operatorname{ad}(Y_{2})^{2}\left(Z_{1}\right)$ $=Y_{2}(Y_{2}Z_{1}-Z_{1}Y_{2})-h_{2}\rhd(Y_{2}Z_{1}-Z_{1}Y_{2})Y_{2}=0.$
On the other hand, since $\operatorname{ad}Y_{2}(Z_{2})=0,$ and $Y_{1}Y_{2}+Y_{2}Y_{1}=0$. The following identity holds.
$\Delta(\operatorname{ad}({Y_{2}})\left(\operatorname{ad}({Y_{1}})\left(Z_{2}\right)\right))$ $=(1\otimes Y_{2}+Y_{2}\otimes 1)(1\otimes\operatorname{ad}({Y_{1}})\left(Z_{2}\right)+\operatorname{ad}({Y_{1}})\left(Z_{2}\right)\otimes 1+2Y_{1}\otimes Z_{2})$
$-(1\otimes\operatorname{ad}({Y_{1}})\left(Z_{2}\right)+\operatorname{ad}({Y_{1}})\left(Z_{2}\right)\otimes 1+2Y_{1}\otimes Z_{2})(1\otimes Y_{2}+Y_{2}\otimes 1)$
$=1\otimes\operatorname{ad}({X_{2}})\left(\operatorname{ad}({X_{1}})\left(Y_{1}\right)\right)+\operatorname{ad}({X_{2}})\left(\operatorname{ad}({X_{1}})\left(Y_{1}\right)\right)\otimes 1.$
Similarly, $\operatorname{ad}(Y_{1})(\operatorname{ad}(Y_{2})(Z_{1}))=0$. Hence $\operatorname{ad}(W_{2})^{2}(W_{3})=0$. We list all identities which can be obtained in this case.
$\operatorname{ad}(W_{3})^{2}(W_{2})=$ $\operatorname{ad}(W_{2})^{2}(W_{6})=\operatorname{ad}(W_{6})^{2}(W_{2})=\operatorname{ad}(W_{3})^{2}(W_{6})=\operatorname{ad}(W_{6})^{2}(W_{3})=0.$
$\operatorname{ad}(W_{2})(W_{3})\cong\operatorname{ad}(W_{3})(W_{2})\cong W_{6},$
$\operatorname{ad}(W_{2})(W_{6})\cong\operatorname{ad}(W_{6})(W_{2})\cong W_{3},$
$\operatorname{ad}(W_{3})(W_{6})\cong\operatorname{ad}(W_{6})(W_{3})\cong W_{2}.$
The final case to confirm is $\operatorname{ad}(W_{4})^{2}(W_{5})=0$:
$\Delta\left(\operatorname{ad}({R_{1}})\left(T_{1}\right)\right)=1\otimes\operatorname{ad}({R_{1}})\left(T_{1}\right)+\operatorname{ad}({R_{1}})\left(T_{1}\right)\otimes 1+R_{1}\otimes T_{1}-R_{2}\otimes T_{2},$
$\Delta\left(\operatorname{ad}({R_{1}})\left(T_{2}\right)\right)=1\otimes\operatorname{ad}({R_{1}})\left(T_{2}\right)+\operatorname{ad}({R_{1}})\left(T_{2}\right)\otimes 1+R_{1}\otimes T_{2}+R_{2}\otimes T_{1},$
$\Delta\left(\operatorname{ad}({R_{2}})\left(T_{1}\right)\right)=1\otimes\operatorname{ad}({R_{2}})\left(T_{1}\right)+\operatorname{ad}({R_{2}})\left(T_{1}\right)\otimes 1+R_{2}\otimes T_{1}+R_{1}\otimes T_{2},$
$\Delta\left(\operatorname{ad}({R_{2}})\left(T_{2}\right)\right)=1\otimes\operatorname{ad}({R_{2}})\left(T_{2}\right)+\operatorname{ad}({R_{2}})\left(T_{2}\right)\otimes 1+R_{2}\otimes T_{2}-R_{1}\otimes T_{1}.$
Then it is direct to see that
$\operatorname{ad}({R_{1}})\left(T_{1}\right)+\operatorname{ad}({T_{2}})\left(R_{2}\right)=0,$ (6.9)
$\operatorname{ad}({R_{2}})\left(T_{1}\right)-\operatorname{ad}({R_{1}})\left(T_{2}\right)=0.$
Hence $\operatorname{ad}(W_{4})(W_{5})=\operatorname{span}\{\operatorname{ad}(R_{1})(T_{1}),\operatorname{ad}(R_{2})(T_{1})\}$. By $R_{1}^{2}=R_{2}^{2}=0$,
$\operatorname{ad}(R_{1})^{2}\left(T_{1}\right)$ $=R_{1}(R_{1}T_{1}+T_{2}R_{1})-(h_{1}h_{2})\rhd(R_{1}T_{1}+T_{2}R_{1})R_{1}=0,$
$\operatorname{ad}(R_{1})^{2}(T_{2})$ $=R_{1}(R_{1}T_{2}-T_{1}R_{1})-(h_{1}h_{2})\rhd(R_{1}T_{2}-T_{1}R_{1})R_{1}=0,$
$\operatorname{ad}(R_{2})^{2}\left(T_{1}\right)$ $=R_{2}(R_{2}T_{1}+T_{2}R_{2})-(h_{1}h_{2})\rhd(R_{2}T_{1}+T_{2}R_{2})R_{2}=0,$
$\operatorname{ad}(R_{2})^{2}(T_{2})$ $=R_{2}(R_{2}T_{2}-T_{1}R_{2})-(h_{1}h_{2})\rhd(R_{2}T_{2}-T_{1}R_{2})R_{2}=0.$
This combines (6.9) implies $\operatorname{ad}(W_{4})^{2}(W_{5})=0$.
The following identities can be obtained in this case.
$\operatorname{ad}(W_{4})^{2}(W_{6})=$ $\operatorname{ad}(W_{6})^{2}(W_{4})=\operatorname{ad}(W_{6})^{2}(W_{5})=\operatorname{ad}(W_{5})^{2}(W_{6})=\operatorname{ad}(W_{5})^{2}(W_{4})=0.$
$\operatorname{ad}(W_{4})(W_{5})\cong\operatorname{ad}(W_{5})(W_{4})\cong W_{6},$
$\operatorname{ad}(W_{4})(W_{6})\cong\operatorname{ad}(W_{6})(W_{4})\cong W_{5},$
$\operatorname{ad}(W_{5})(W_{6})\cong\operatorname{ad}(W_{6})(W_{5})\cong W_{4}.$
We have listed all cases, which proves this lemma. ∎
To establish that the tuple $W$ admits all reflections, the following structural observation is essential.
Lemma 6.13.
Suppose $W$ admits the reflection sequence $(i_{1},i_{2},..,i_{l})$, where $i_{1},..,i_{l}\in\{1,2,3\}$. Then for $j\in\{1,2,3\}$,
$\operatorname{deg}(R_{i_{l}}(...(R_{i_{1}}(W))...)_{j})\neq 1,h_{1}h_{2}h_{3}.$ (6.10)
and
$\prod_{j=1}^{3}\operatorname{deg}(R_{i_{l}}(...(R_{i_{1}}(W))...)_{j})=h_{1}h_{2}h_{3}.$
Proof.
We prove it by induction. For the base case $l=1$, Lemma 5.5 gives,
$R_{1}(W)\cong(W_{1},W_{4},W_{5}),\ R_{2}(W)\cong(W_{4},W_{2},W_{6}),\ R_{3}(W)\cong(W_{5},W_{6},W_{3}).$
It is straightforward to verify that for each $1\leq i,j\leq 3$, $\prod_{j=1}^{3}R_{i}(W)=h_{1}h_{2}h_{3}$.
Now assume, for some $l\geq 2$, that $W$ admits the reflection sequence $(i_{1},i_{2},..,i_{l-1})$, and
$\prod_{j=1}^{3}\operatorname{deg}(R_{i_{l-1}}(...(R_{i_{1}}(W))...)_{j})=h_{1}h_{2}h_{3}.$
Let us denote $\operatorname{deg}(R_{i_{l-1}}(...(R_{i_{1}}(W))...)_{j})=h_{j}^{\prime}$ for each $1\leq j\leq 3$, where $h_{j}^{\prime}\in\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}$. Then for any $1\leq i_{l}\leq 3$, we compute
$\prod_{j=1}^{3}\operatorname{deg}(R_{i_{l}}(...(R_{i_{1}}(W))...)_{j})$ $=(\prod_{k\neq i_{l}}h_{k}^{\prime}h_{i_{l}})h_{i_{l}}^{-1}$
$=h_{1}h_{2}h_{3},$
which establishes the second claim.
Without loss of generality, Suppose for contradiction that $\operatorname{deg}(R_{i_{l}}(...(R_{i_{1}}(W))...)_{1})=h_{1}h_{2}h_{3}$. We must have $\operatorname{deg}(R_{i_{l}}(...(R_{i_{1}}(W))...)_{2})=\operatorname{deg}(R_{i_{l}}(...(R_{i_{1}}(W))...)_{3})=h^{\prime}$ by the second claim. If $i_{l}=1$, then
$R_{i_{l-1}}(...(R_{i_{1}}(W))...)=(h_{1}h_{2}h_{3},h_{1}h_{2}h_{3}h^{\prime},h_{1}h_{2}h_{3}h^{\prime}),$
which contradicts the assumption. If $i_{1}=2$,the degrees become
$R_{i_{l-1}}(...(R_{i_{1}}(W))...)=(h_{1}h_{2}h_{3}h^{\prime},h^{\prime},1),$
and if $i_{1}=3$
$R_{i_{l-1}}(...(R_{i_{1}}(W))...)=(h_{1}h_{2}h_{3}h^{\prime},1,h^{\prime}).$
Both of which again lead to contradictions.
A similar contradiction arises if we assume $\operatorname{deg}(R_{i_{l}}(...(R_{i_{1}}(W))...)_{1})=1$,
$\operatorname{deg}(R_{i_{l}}(...(R_{i_{1}}(W))...)_{2})=k^{\prime}$, $\operatorname{deg}(R_{i_{l}}(...(R_{i_{1}}(W))...)_{3})=k^{\prime\prime}$, where $k^{\prime}\cdot k^{\prime\prime}=h_{1}h_{2}h_{3}$. Now if $i_{l}=1$, we have
$R_{i_{l-1}}(...(R_{i_{1}}(W))...)=(1,k^{\prime},k^{\prime\prime}).$
If $i_{l}=2$, we obtain
$R_{i_{l-1}}(...(R_{i_{1}}(W))...)=(k^{\prime},k^{\prime},k^{\prime}k^{\prime\prime}),$
and if $i_{l}=3$
$R_{i_{l-1}}(...(R_{i_{1}}(W))...)=(k^{\prime\prime},k^{\prime}k^{\prime\prime},k^{\prime\prime}).$
All cases are leading to contradictions. This completes the induction. ∎
Proof of Theorem 6.11.
We proceed by induction to show that $W$ admits all reflection sequences. Without loss of generality, we first consider the 11st-reflection. By proof of Lemma 6.12, for all $1\leq j\leq 3$, we have $R_{1}(W)_{j}\in S$. Recall that two tuples $W$ and $W^{\prime}$ are isomorphic if and only if their corresponding components are isomorphic. $R_{1}(W)\cong(W_{1},W_{4},W_{5})$. By Lemma 5.5, $R_{i}(R_{1}(W))\cong R_{i}(W_{1},W_{4},W_{5})$ for $1\leq i\leq 3$. Applying Lemma 6.12 again, we conclude that $(W_{1},W_{4},W_{5})$ admits the $i$-th reflection by analogous reasoning.
Now, suppose W admits the reflection sequence $(i_{1},i_{2},\ldots,i_{l-1})$, where $1\leq i_{1},...,i_{l-1}\leq 3$. Then we have $\operatorname{deg}(R_{i_{l-1}}(...(R_{i_{1}}(W))...)_{j})\neq 1,h_{1}h_{2}h_{3}$ for $1\leq j\leq 3$ and $\prod_{j=1}^{3}\operatorname{deg}(R_{i_{l-1}}(...(R_{i_{1}}(W))...)_{j})=h_{1}h_{2}h_{3}.$ Consequently, for any $1\leq m<n\leq 3$, the product
$\operatorname{deg}(R_{i_{l-1}}(...(R_{i_{1}}(W))...)_{m})\cdot\operatorname{deg}(R_{i_{l-1}}(...(R_{i_{1}}(W))...)_{n})$
is neither $1$ nor $h_{1}h_{2}h_{3}$. Therefore $W$ admits the $(i_{1},i_{2},...,i_{l-1},i_{l})$-th reflection by Lemma 6.12. This completes the induction and shows that $W$ admits all reflection sequences.
For any $N\in\mathcal{F}_{3}(W)$, we have $\operatorname{deg}(N_{i})\neq 1,h_{1}h_{2}h_{3}$ for $1\leq i\leq 3$ by Lemma 6.13. Therefore, $\mathcal{B}(N_{i})$ is finite-dimensional since $N_{i}$ belongs to $S$. Thus $W$ will give rise to a semi-Cartan graph $\mathcal{G}(W)$.
Now for all $N\in\mathcal{F}_{3}(W)$, Lemma 6.12 implies that the generalized Cartan matrix $A^{N}$ is given by
$\begin{pmatrix}2&-1&-1\\ -1&2&-1\\ -1&-1&2\end{pmatrix}$
Thus $\mathcal{G}(W)$ is a standard semi-Cartan graph. ∎
Corollary 6.14.
The Nichols algebra $\mathcal{B}(W)$ is infinite-dimensional.
Proof.
$\mathcal{G}(W)$ is a standard semi-Cartan graph by Theorem 6.11. Suppose, for contradiction, that $\mathcal{B}(W)$ is finite-dimensional, then $A^{W}$ is a finite Cartan matrix by Theorem 6.7. However
$A^{W}=\begin{pmatrix}2&-1&-1\\ -1&2&-1\\ -1&-1&2\end{pmatrix}$
is not a finite Cartan matrix. This contradiction implies that $\mathcal{B}(W)$ must be infinite-dimensional. ∎
Proof of Theorem 6.9.
By Remark 6.8(1), there is a group surjection satisfying
$\pi:G\longrightarrow H,\ g_{i}\mapsto h_{i},\ 1\leq i\leq 3.$
Furthermore, $\pi^{*}(\phi)=\Phi$.
We construct such a linear map for $1\leq j\leq 2$,
$F:V_{1}\oplus V_{2}\oplus V_{3}\rightarrow W_{1}\oplus W_{2}\oplus W_{3},\ X_{j}^{\prime}\mapsto X_{j},Y_{j}^{\prime}\mapsto Y_{j},Z_{j}^{\prime}\mapsto Z_{j}$
A direct verification shows that for each $1\leq i\leq 3$, $v_{i}\in V_{i}$ and $g\in G$, the following compatibilities hold:
$(\pi\otimes F)\delta_{V_{i}}(v_{i})$ $=\delta_{W_{i}}(F(v_{i})),$
$F(g\rhd v_{i})$ $=\pi(g)\rhd F(v_{i}).$
By [QQG, Lemma 4.4], we deduce that
$\mathcal{B}(V)\cong\mathcal{B}(W)$
in the sense of [QQG, Definition 4.3]. It follows that $\mathcal{B}(V)$ is infinite-dimensional. ∎
Remark 6.15.
For the sake of simplicity, this article studies only the cosemisimple pointed coquasi-Hopf algebras under minor conditions. We hope to extend this theory to all coquasi-Hopf algebras with bijective antipodes.
Acknowledgment This work is supported by the National Key R&D Program of China
2024YFA1013802 and NSFC 12271243.

### v2

@@
-Reflection of Nichols Algebras over Coquasi-Hopf Algebras
+arXiv:2512.04560v2[math.QA] 27 Feb 2026
 Abstract
-This paper extends the foundational reflection theory of Nichols algebras to the setting of some certain coquasi-Hopf algebras.
+This paper extends the foundational reflection theory of Nichols algebras to the setting of certain coquasi-Hopf algebras.
 Our primary motivation arises from the classification of pointed finite-dimensional coquasi-Hopf algebras.
 We develop a reflection theory for tuples of simple Yetter-Drinfeld modules in the category $^{G}_{G}\mathcal{YD}^{\Phi}$}, where $G$ is a finite group and $\Phi$ is a 3-cocycle on $G$.
-We prove that such a tuple gives rise to a semi-Cartan graph if admitting all reflections.
+We prove that such a tuple gives rise to a semi-Cartan graph if it admits all reflections.
 Consequently, its Weyl groupoid is well-defined.
 We further establish several criteria for the finite-dimensionality of Nichols algebras in terms of the associated semi-Cartan graph.
-As an application, we provide a new proof for the infinite-dimensionality of a specific class of Nichols algebras previously studied in huang2024classification, bypassing extensive computational arguments.
+As an application, we provide a new proof for the infinite-dimensionality of a specific class of Nichols algebras previously studied in Huang et al.
+, bypassing extensive computational arguments.
+Mathematics Subject Classification: 16T05; 18M15; 20F55
 1 Introduction
@@
 A crucial breakthrough in this direction was achieved through the systematic study of Nichols algebras.
-Heckenberger introduced the revolutionary concepts of semi-Cartan graphs and Weyl groupoids [Hec06, Hec09, Hec10], which provide powerful combinatorial tools for understanding the structure and classification of Nichols algebras.
+Heckenberger introduced the revolutionary concepts of semi-Cartan graphs and Weyl groupoids [Heckenberger, 2006,Heckenberger, 2009,Heckenberger, 2010], which provide powerful combinatorial tools for understanding the structure and classification of Nichols algebras.
 The reflection theory of Nichols algebras represents one of the most important developments in the field.
-The fundamental theorem states that a tuple of simple Yetter-Drinfeld modules over Hopf algebra with bijective antipode will give rise to a semi-Cartan graph provided that it admits all reflections [AHS10,HS10].
-This result was proved again using a categorical approach in the setting of Hopf algebras in HSdualpair.
+The fundamental theorem states that a tuple of simple Yetter-Drinfeld modules over a Hopf algebra with bijective antipode will give rise to a semi-Cartan graph provided that it admits all reflections [Andruskiewitsch et al., 2010,Heckenberger and Schneider, 2010b].
+This result was proved again using a categorical approach in the setting of Hopf algebras in Heckenberger and Schneider (2013a).
 We recall that a semi-Cartan graph consists of the following data,
@@
 Definition 1.2.
-We denote by $\mathcal{D}(\mathcal{X},\operatorname{End}(\mathbb{Z}^{\mathbb{I}}))$ to be the category with objects $\operatorname{Ob}\mathcal{D}(\mathcal{X},\operatorname{End}(\mathbb{Z}^{\mathbb{I}}))=\mathcal{X}$, and morphisms
+Let $\mathcal{D}(\mathcal{X},\operatorname{End}(\mathbb{Z}^{\mathbb{I}}))$ denote the category with objects $\operatorname{Ob}\mathcal{D}(\mathcal{X},\operatorname{End}(\mathbb{Z}^{\mathbb{I}}))=\mathcal{X}$, and morphisms
 $\operatorname{Hom}(X,Y)=\{(Y,f,X)\mid f\in\operatorname{End}(\mathbb{Z}^{\mathbb{I}})\},$
@@
 We call the smallest subcategory of $\mathcal{D}(\mathcal{X},\operatorname{End}(\mathbb{Z}^{\mathbb{I}}))$ which contains all morphisms $(r_{i}(X),s_{i}^{X},X)$ with $i\in\mathbb{I}$, $X\in\mathcal{X}$ the Weyl groupoid of $\mathcal{G}$, denoted by $\mathcal{W}(\mathcal{G})$.
-In the context of Nichols algebras, the Weyl groupoid captures the symmetries and reflection properties of the underlying braided vector space, establishing a bridge between the algebraic structure of Nichols algebras and the combinatorics of root systems [AHS10].
+In the context of Nichols algebras, the Weyl groupoid captures the symmetries and reflection properties of the underlying braided vector space, establishing a bridge between the algebraic structure of Nichols algebras and the combinatorics of root systems [Andruskiewitsch et al., 2010].
 The interplay between Weyl groupoid theory and reflection theory has led to spectacular progress in the classification of finite-dimensional Nichols algebras.
-The classification program, initiated by Andruskiewitsch and Schneider in a series of papers [AS98,AS00,AS02,AS10], aims to determine all finite-dimensional pointed Hopf algebras over algebraically closed fields of characteristic zero whose group of group-like elements is abelian.
-A cornerstone of this classification is Heckenberger’s complete classification of finite-dimensional Nichols algebras of diagonal type [Hec09].
-Building on this classification, Angiono made a breakthrough contribution by providing explicit presentations by generators and relations for all Nichols algebras of diagonal type [A13, A15].
-The lifting problem of Nichols algebra of diagonal type has been addressed through cocycle deformation [AA+14, AG19, AA17].
-The classification of finite-dimensional Nichols algebras over non-abelian groups has achieved excellent progress as well [AGn99,AGn03,AHS10,MS00,FK99,Gn00,Gn+11,HLV12,HS10,HS10b,HS10c,HS14,Hs15,HV17].
+The classification program, initiated by Andruskiewitsch and Schneider in a series of papers [Andruskiewitsch and Schneider, 1998,Andruskiewitsch and Schneider, 2000,Andruskiewitsch and Schneider, 2002,Andruskiewitsch and Schneider, 2010], aims to determine all finite-dimensional pointed Hopf algebras over algebraically closed fields of characteristic zero whose group of group-like elements is abelian.
+A cornerstone of this classification is Heckenberger’s complete classification of finite-dimensional Nichols algebras of diagonal type [Heckenberger, 2009].
+Building on this classification, Angiono made a breakthrough contribution by providing explicit presentations by generators and relations for all Nichols algebras of diagonal type [Angiono, 2013,Angiono, 2015].
+The lifting problem of Nichols algebras of diagonal type has been addressed through cocycle deformation [Andruskiewitsch et al., 2014,Angiono and García Iglesias, 2019,Andruskiewitsch and Angiono, 2017].
+The classification of finite-dimensional Nichols algebras over non-abelian groups has achieved excellent progress as well [Andruskiewitsch and Graña, 1999,Andruskiewitsch and Graña, 2003,Andruskiewitsch et al., 2010,Milinski and Schneider, 2000,Fomin and Kirillov, 1999,Graña, 2000,Graña et al., 2011,Dello and Zhang, 2016,Heckenberger and Schneider, 2010b,Heckenberger and Schneider, 2013b,Heckenberger and Schneider, 2010a,Heckenberger and Vendramin, 2014,Heckenberger and Vendramin, 2015,Heckenberger and Vendramin, 2017].
 In this paper, we extend the reflection theory to the setting of some special coquasi-Hopf algebras.
-Our motivation comes from the classification of coradically graded pointed finite-dimensional coquasi-Hopf algebras, which has been studied in [rank1,coquasitriangular,tame-rep,FQQR2,QQG,huang2024classification,LiLiu1].
+Our motivation comes from the classification of coradically graded pointed finite-dimensional coquasi-Hopf algebras, which has been studied in [Huang et al., 2011,Huang and Liu, 2012,Huang et al., 2015,Huang et al., 2021,Huang et al., 2020,Huang et al., ,Li and Liu, 2025].
 In the classification of pointed Hopf algebras over abelian groups, Nichols algebras of non-diagonal type do not appear.
 However, in the coquasi-Hopf algebra setting, Nichols algebras of non-diagonal type do occur and present greater challenges.
-The authors in huang2024classification used extensive computations to establish that a certain class of such Nichols algebras is infinite-dimensional.
-We later provided an alternative proof of their main result via categorical Morita equivalence LiLiu1.
+The authors in Huang et al.
+used extensive computations to establish that a certain class of such Nichols algebras is infinite-dimensional.
+We later provided an alternative proof of their main result via categorical Morita equivalence Li and Liu (2025).
 Nonetheless, these arguments are not entirely natural, and we hope that the theory of semi-Cartan graphs can be extended to coquasi-Hopf algebras.
 If such a generalization holds, it could be applied to the classification of finite-dimensional coquasi-Hopf algebras over non-abelian groups.
-Our approach combines the methods of Simon and HSdualpair together.
-The next theorem generalizes [HSdualpair, Theorem 7.1] to the case of coquasi-Hopf algebras with bijective antipodes under some minor conditions, which is crucial to description of reflection of Nichols algebras.
+Our approach combines the methods of Barvels et al.
+(2015) and Heckenberger and Schneider (2013a).
+The next theorem generalizes [Heckenberger and Schneider, 2013a, Theorem 7.1] to the case of coquasi-Hopf algebras with bijective antipodes under some minor conditions, which is crucial to the description of the reflection of Nichols algebras.
 Theorem A.
@@
 Now we restrict to the case of pointed cosemisimple coquasi-Hopf algebras.
-Let $G$ be a finite group, and $\Phi$ be a $3$-cocycle on $G$, which gives rise to the Yetter-Drinfeld module category $^{G}_{G}\mathcal{YD}^{\Phi}$} QQG.
+Let $G$ be a finite group, and $\Phi$ be a $3$-cocycle on $G$, which gives rise to the Yetter-Drinfeld module category $^{G}_{G}\mathcal{YD}^{\Phi}$} Huang et al.
+(2020).
 Let $\theta\geq 1$ be a positive integer, $\mathbb{I}=\{1,2,..,\theta\}$ and $M=(M_{1},\ldots,M_{\theta})$, where $M_{1},\ldots,M_{\theta}\in{{}^{G}_{G}\mathcal{YD}^{\Phi}}$ are irreducible Yetter-Drinfeld modules.
-We always denote $\mathcal{B}(M)$ to be the Nichols algebra of $M_{1}\oplus\cdots\oplus M_{\theta}$ in the following context.
-Suppose $M$ admits $i$-th reflection for some $1\leq i\leq\theta$ (see Definition 5.1).
+We always denote $\mathcal{B}(M)$ as the Nichols algebra of $M_{1}\oplus\cdots\oplus M_{\theta}$ in the following context.
+Suppose $M$ admits the $i$-th reflection for some $1\leq i\leq\theta$ (see Definition 5.1).
 Then the $i$-th reflection of $M$ is denoted by $R_{i}(M)=(V_{1},\ldots,V_{\theta})$, which is a tuple consisting of finite-dimensional irreducible Yetter-Drinfeld modules.
@@
 Theorem B.
-(Theorem 5.7) With above assumptions on $M$.
-Suppose $M$ admits the $i$-th reflection for $1\leq i\leq\theta$, and $\mathcal{B}(M_{i})$ is finite-dimensional.
+(Theorem 5.7) Under the above assumptions on M, suppose $M$ admits the $i$-th reflection for $1\leq i\leq\theta$, and $\mathcal{B}(M_{i})$ is finite-dimensional.
 Then there is an isomorphism of Hopf algebras in $^{G}_{G}\mathcal{YD}^{\Phi}$}:
@@
 Readers who are familiar with reflection theory in the Hopf algebra context will recognize this formula, we emphasize that our contribution lies in establishing this result in the framework of coquasi-Hopf algebras.
-The extension from Hopf algebras to coquasi-Hopf algebras introduces technical challenges, particularly in sections 3 and 4 where the non-strict monoidal structure necessitates a reworking of the reflection theory.
-Now with the help of above theorems, our main result follows naturally.
+The extension from Hopf algebras to coquasi-Hopf algebras introduces technical challenges, particularly in Sections 3 and 4 where the non-strict monoidal structure necessitates a reworking of the reflection theory.with the help of the above theorems, our main result follows naturally.
 Theorem C.
@@
 We show that if $\mathcal{B}(M)$ is finite-dimensional, then the semi-Cartan graph $\mathcal{G}(M)$ is a finite semi-Cartan graph.
-Furthermore, if $\mathcal{G}(M)$ is moreover a standard finite semi-Cartan graph, that is $A^{X}=A^{Y}$, for all $X,Y\in\mathcal{X}$, then $A^{M}$ must be a finite Cartan matrix.
-As an application of these results, we provide a third proof of [huang2024classification, Proposition 4.1].
+Furthermore, if $\mathcal{G}(M)$ is a standard finite semi-Cartan graph, that is $A^{X}=A^{Y}$, for all $X,Y\in\mathcal{X}$, then $A^{M}$ must be a finite Cartan matrix.
+As an application of these results, we provide a third proof of [Huang et al., , Proposition 4.1].
 The paper is organized as follows.
@@
 2.1 Coquasi-Hopf algebras
-By definition, coquasi-Hopf algebras are exactly the dual of Drinfeld’s quasi-Hopf algebras Drinfeld. Their formal definition can be given as follows.
+By definition, coquasi-Hopf algebras are exactly the dual of Drinfeld’s quasi-Hopf algebras Drinfeld (1989). Their formal definition can be given as follows.
 Definition 2.1.
@@
     $\mu:\mathbbm{k}\rightarrow H,\qquad\lambda\mapsto\lambda 1_{H},$
-  • A convolution-invertible map $\Phi:H^{\otimes 3}\rightarrow\mathbbm{k}$ called associator,
+  • A convolution-invertible map $\Phi:H^{\otimes 3}\rightarrow\mathbbm{k}$ called an associator,
   • A coalgebra antimorphism $\mathcal{S}:H\rightarrow H$,
@@
 Furthermore, for homogeneous elements $a,b,c\in\operatorname{gr}(H)$, $\operatorname{gr}(\Phi)(a,b,c)=0$ unless they all lie in $H_{0}$.
-Following the definition in [preantipode], we now recall the definition of a preantipode, which plays an important role in further calculations.
+Following the definition in [Ardizzoni and Pavarin, 2012], we now recall the definition of a preantipode, plays an important role in subsequent calculations.
 Definition 2.2.
-[preantipode, Definition 3.6] Let $(H,\Phi)$ be a coquasi-bialgebra, a preantipode for $H$ is a $\mathbbm{k}$-linear map $\mathbb{S}:H\rightarrow H$ such that, for all $h\in H$,
+[Ardizzoni and Pavarin, 2012, Definition 3.6] Let $(H,\Phi)$ be a coquasi-bialgebra, a preantipode for $H$ is a $\mathbbm{k}$-linear map $\mathbb{S}:H\rightarrow H$ such that, for all $h\in H$,
 $\mathbb{S}(h_{1})_{1}h_{2}\otimes\mathbb{S}(h_{1})_{2}$ $=1_{H}\otimes\mathbb{S}(h),$
@@
 $\Phi(h_{1},\mathbb{S}(h_{2}),h_{3})$ $=\varepsilon(h).$
-Furthermore, [preantipode, Remark 3.7] provides us a useful property:
+Furthermore, [Ardizzoni and Pavarin, 2012, Remark 3.7] provides us a useful property:
 $h_{1}\mathbb{S}(h_{2})=\varepsilon(\mathbb{S}(h))1_{H}=\mathbb{S}(h_{1})h_{2},\ \text{for all}\ h\in H.$ (2.7)
-Meanwhile, when $H$ is equipped with the structure of a coquasi-Hopf algebra with antipode $(\mathcal{S},\alpha,\beta)$, then preantipode for $H$ always exists.
+Meanwhile, when $H$ is equipped with the structure of a coquasi-Hopf algebra with an antipode $(\mathcal{S},\alpha,\beta)$, then preantipode for $H$ always exists.
 Lemma 2.3.
-[preantipode, Theorem 3.10] Let $H$ be a coquasi-Hopf algebra with antipode $(\mathcal{S},\alpha,\beta)$, then
+[Ardizzoni and Pavarin, 2012, Theorem 3.10] Let $H$ be a coquasi-Hopf algebra with an antipode $(\mathcal{S},\alpha,\beta)$, then
 $\mathbb{S}:=\beta*\mathcal{S}*\alpha$ (2.8)
@@
 We now turn our attention to the Yetter-Drinfeld module category structure.
-The definition of Yetter-Drinfeld module category $^{H}_{H}\mathcal{YD}$} over an arbitrary coquasi Hopf-algebra $H$ was already given in [YD-module].
+The definition of the Yetter-Drinfeld module category $^{H}_{H}\mathcal{YD}$} over an arbitrary coquasi Hopf-algebra $H$ was already given in [Ardizzoni and Pavarin, 2013].
 It is defined as the center $\mathcal{Z}({{{}^{H}}\mathscr{M}})$ of the comodule category $^{H}\mathscr{M}$}.
@@
 Definition 2.5.
-[YD-module, Definition 3.1] Let $H$ be a coquasi-Hopf algebra with associator $\Phi$. A left-left Yetter–Drinfeld module over $H$ is a triple $(V,\delta_{V},\rhd)$ such that:
+[Ardizzoni and Pavarin, 2013, Definition 3.1] Let $H$ be a coquasi-Hopf algebra with associator $\Phi$. A left-left Yetter–Drinfeld module over $H$ is a triple $(V,\delta_{V},\rhd)$ such that:
   • $(V,\delta_{V})$ is a left comodule of $H$ and we denote $\delta_{V}(v)$ by $v_{-1}\otimes v_{0}$ as usual;
@@
 $v_{-1}\otimes(h\rhd v_{0})$ $=\varepsilon(v_{-1})\varepsilon(h_{1})v_{-2}\otimes(h_{2}\rhd v_{0})$
-$\overset{(\ref{2.18})}{=}s(\mathcal{S}(h_{1}),h_{3}v_{-1})t(h_{2},v_{-2})v_{-3}\otimes(h_{4}\rhd v_{0})$
-$\overset{(\ref{2.17})}{=}s(\mathcal{S}(h_{1}),h_{5}v_{-1})t(h_{3},v_{-3})\mathcal{S}(h_{2})(h_{4}v_{-2})\otimes h_{6}\rhd v_{0}$
+[…]
+[…]
 […]
@@
 For our purpose, we aim to describe Yetter-Drinfeld modules over coquasi-Hopf algebra of the form $(\mathbbm{k}G,\Phi)$, where $G$ is a finite group and $\Phi$ a $3$-cocycle on $G$.
-Assume that $V$ is a left $kG$-comodule with comodule structure map $\delta_{L}:V\rightarrow kG\otimes V$.
+Assume that $V$ is a left $kG$-comodule with a comodule structure map $\delta_{L}:V\rightarrow kG\otimes V$.
 Define
@@
 The category of Yetter-Drinfeld modules $^{A}_{A}\mathcal{YD}(\mathcal{C})$} is well-defined.
-The reader may refer to [rootsys, Section 3] for details.
+The reader may refer to [Heckenberger and Schneider, [2020] ©2020, Section 3] for details.
 2.3 Bosonization for coquasi-Hopf algebras
 Bosonization for Hopf algebras is well-known.
-In [quasibon], the authors introduced bosonization for quasi-Hopf algebras.
-Dually, bosonization for coquasi-Hopf algebras is given in YD-module.
+In [Bulacu and Nauwelaerts, 2002], the authors introduced bosonization for quasi-Hopf algebras.
+Dually, bosonization for coquasi-Hopf algebras is given in Ardizzoni and Pavarin (2013).
 Let $H$ be a coquasi-Hopf algebra with associator $\Phi$.
@@
 Lemma 2.8.
-[YD-module, Theorem 5.2] Let us consider on $N:=R\otimes H$ the following structures:
+[Ardizzoni and Pavarin, 2013, Theorem 5.2] Let us consider on $N:=R\otimes H$ the following structures:
 $(r\otimes h)(s\otimes k)$ $=\frac{\Phi(h_{2},s_{-1},k_{2})\Phi(r_{-1},(h_{3}\rhd s_{0})_{-1},h_{5}k_{4})}{\Phi(r_{-2},h_{1},s_{-2}k_{1})\Phi((h_{3}\rhd s_{0})_{-2},h_{4},k_{3})}r_{0}(h_{3}\rhd s_{0})_{0}\otimes h_{6}k_{5},$
@@
 where $r,s,t\in R$, $h,k,l\in H$. With above operations, $N$ is a coquasi-Hopf algebra.
-Now suppose $N$ are $H$ are both coquasi-Hopf algebras. Furthermore, assume there exist morphisms of coquasi-Hopf algebras
+Now suppose $N$ and $H$ are both coquasi-Hopf algebras. Furthermore, assume there exist morphisms of coquasi-Hopf algebras
 $\pi:N\rightarrow H\ \text{ and }\sigma:H\rightarrow N$
@@
 Lemma 2.9.
-[YD-module, Theorem 5.8] Under above assumptions, $L:=N^{\operatorname{co}H}$ is a Hopf algebra in $^{H}_{H}\mathcal{YD}$}. For all $a\in N$, we denote
+[Ardizzoni and Pavarin, 2013, Theorem 5.8] Under above assumptions, $L:=N^{\operatorname{co}H}$ is a Hopf algebra in $^{H}_{H}\mathcal{YD}$}. For all $a\in N$, we denote
 $\tau(a)=\Phi_{N}(a_{1},\sigma\mathbb{S}\pi(a_{3})_{1},a_{4})a_{2}\sigma\mathbb{S}\pi(a_{3})_{2}.$ (2.29)
@@
 $\phi(r\otimes h)=r\sigma(h),\ \ \phi^{-1}(a)=\tau(a_{1})\otimes\pi(a_{2}).$
-For reader’s convenience, we still give an example about $^{G}_{G}\mathcal{YD}^{\Phi}$}, where $G$ is a finite group and $\Phi$ is a $3$-cocycle on $G$.
+For the reader’s convenience, we provide an example about $^{G}_{G}\mathcal{YD}^{\Phi}$}, where $G$ is a finite group and $\Phi$ is a $3$-cocycle on $G$.
 Suppose $R$ is a Hopf algebra in $^{G}_{G}\mathcal{YD}^{\Phi}$}.
@@
 where $^{x}R:=\{X\in R\mid\rho(x)=x\otimes X\}$.
-For simplicity, in this paper, homogeneous elements in $^{G}_{G}\mathcal{YD}^{\Phi}$} are denoted by capital letters, say $X,Y,Z$..., and the associated degrees are denoted by their lower cases, say $x,y,z$
+For simplicity, in this paper, homogeneous elements in $^{G}_{G}\mathcal{YD}^{\Phi}$} are denoted by capital letters, say $X,Y,Z$..., and the associated degrees are denoted by their lowercase letters, say $x,y,z$
 Example 2.10.
@@
 Our primary focus now shifts to analyzing the linear map $\operatorname{ad}$. Given $g\in G$ and $X\in{{}^{x}R}$,
-$\operatorname{ad}(g)(X)=\frac{\Phi(gx,g^{-1},g)}{\Phi(g,g^{-1},g)}(gX)g^{-1}=g\rhd X.$
+$\operatorname{ad}(g)(X)=\frac{\Phi(gx,g^{-1},g)}{\Phi(g,g^{-1},g)}(gX)g^{-1}.$
 Now suppose $Y$ is a $(y,1)$-primitive element in $H$, we first need to compute what $\mathbb{S}(Y)$ is,
@@
 The definition of the bosonization naturally extends to the case that $R$ is a Hopf algebra in the braided monoidal category $^{A}_{A}\mathcal{YD}(\mathcal{C})$}, where $\mathcal{C}$ is now an arbitrary braided monoidal category.
-We collect several key results from B95 and AF00 that will be essential for the construction presented in Section 3.
+We collect several key results from Bespalov (1995) and Alonso Alvarez and Fernández Vilaboa (2000) that will be essential for the construction presented in Section 3.
 Lemma 2.12.
-[AF00, Theorem 3.2] Let $H$ and $A$ be Hopf algebras in a braided monoidal category $\mathcal{C}$.
+[Alonso Alvarez and Fernández Vilaboa, 2000, Theorem 3.2] Let $H$ and $A$ be Hopf algebras in a braided monoidal category $\mathcal{C}$.
 Let $\pi:H\to A$ and $\iota:A\to H$ be Hopf algebra morphisms such that $\pi\circ\iota=\mathrm{id}_{A}$.
@@
 Lemma 2.13.
-[B95, Proposition 4.2.3] Let $A$ be a Hopf algebra in $\mathcal{C}$ and $K$ a Hopf algebra in $^{A}_{A}\mathcal{YD}(\mathcal{C})$}. There is an obvious isomorphism of braided monoidal categories
+[Bespalov, 1995, Proposition 4.2.3] Let $A$ be a Hopf algebra in $\mathcal{C}$ and $K$ a Hopf algebra in $^{A}_{A}\mathcal{YD}(\mathcal{C})$}.
+There is an obvious isomorphism of braided monoidal categories
 ${{}^{K\#A}_{K\#A}\mathcal{YD}}(\mathcal{C})\cong{{}^{K}_{K}\mathcal{YD}}({{}^{A}_{A}\mathcal{YD}(\mathcal{C})}).$ (2.31)
-2.4 Nichols algebras in braided monoidal category
+2.4 Nichols algebras in a braided monoidal category
 Having established the necessary framework for bosonization, we now turn to the theory of Nichols algebras.
@@
 We present the definition of Nichols algebras in $\mathcal{C}$.
-Let $V$ be a nonzero object in $\mathcal{C}$, by $T(V)$ we denote the tensor algebra in $\mathcal{C}$ generated freely by $V$. Here we denote that
+Let $V$ be a nonzero object in $\mathcal{C}$; by $T(V)$ we denote the tensor algebra in $\mathcal{C}$ generated freely by $V$. Here we denote that
 $V^{\otimes n}:=\left(\cdots\left(\left(V\otimes V\right)\otimes V\right)\cdots\otimes V\right).$
@@
 Definition 2.15.
-[defofnichols, Definition 2.4] For a given object $V\in\mathcal{C}$ the Nichols algebra $\mathcal{B}(V)$ is the unique Hopf algebra in $\mathcal{C}$ that satisfies the following conditions:
+[Meir, 2022, Definition 2.4] For a given object $V\in\mathcal{C}$ the Nichols algebra $\mathcal{B}(V)$ is the unique Hopf algebra in $\mathcal{C}$ that satisfies the following conditions:
 $(1)$ The Hopf algebra $\mathcal{B}(V)$ is graded by the non-negative integers.
@@
 That is $F(T(V))\cong T(F(V))$.
-Recall that the Nichols algebra in $\mathcal{C}$ are of the form $T(X)/I$ where $X\in\mathcal{C}$ and $I$ is the unique maximal graded Hopf ideal in $T(V)$ generated by homogeneous elements of degree greater than or equal to $2$.
+Recall that the Nichols algebras in $\mathcal{C}$ are of the form $T(X)/I$ where $X\in\mathcal{C}$ and $I$ is the unique maximal graded Hopf ideal in $T(V)$ generated by homogeneous elements of degree greater than or equal to $2$.
 According to $F(T(V))\cong T(F(V))$, we have
@@
 Here $F(I)$ is a homogeneous Hopf ideal of degree greater than or equal to $2$ of $T(F(V))$.
-Note that Nichols algebra generated by $F(V)$ must be of the form $T(F(V))/J$, where $J$ is the unique maximal homogeneous graded Hopf ideal of $T(F(V))\in\mathcal{D}$ with degree greater than or equal to $2$.
+Note that the Nichols algebra generated by $F(V)$ must be of the form $T(F(V))/J$, where $J$ is the unique maximal homogeneous graded Hopf ideal of $T(F(V))\in\mathcal{D}$ with degree greater than or equal to $2$.
 Hence $F(I)\subset J$ and $T(F(V))/J\subset T(F(V))/F(I)$.
-On the other hand, $F^{-1}:\mathcal{D}\longrightarrow\mathcal{C}$ is inverse of $F$, which is an exact monoidal functor.
+On the other hand, $F^{-1}:\mathcal{D}\longrightarrow\mathcal{C}$ is the inverse of $F$, which is an exact monoidal functor.
 Hence
@@
 3 Monoidal equivalence between Yetter-Drinfeld module categories
-We build upon the theory of Hopf pairings and partial dualization in braided monoidal categories Simon.
+We build upon the theory of Hopf pairings and partial dualization in braided monoidal categories Barvels et al.
+(2015).
 We apply this theory to the setting of coquasi-Hopf algebras.
@@
 3.1 Hopf pairings in braided monoidal categories
-In this subsection, $\mathcal{C}$ always represents a braided monoidal category with associative isomorphism $a$ and braiding isomorphism $c$ except for special instructions.
+In this subsection, $\mathcal{C}$ always represents a braided monoidal category with associative isomorphism $a$ and braiding isomorphism $c$ unless stated otherwise.
 We begin by introducing the notion of Hopf pairings.
@@
 $\omega\circ(\operatorname{id}_{A}\otimes\eta_{B})$ $=\varepsilon_{A}.$
-A Hopf pairing $\omega:A\otimes B\to 1$ is called non-degenerate, if there is a morphism $\omega^{\prime}:1\rightarrow B\otimes A$, such that
+A Hopf pairing $\omega:A\otimes B\to 1$ is called non-degenerate if there is a morphism $\omega^{\prime}:1\rightarrow B\otimes A$, such that
 $(\omega\otimes\operatorname{id}_{A})\circ a^{-1}_{A,B,A}\circ(\operatorname{id}_{A}\otimes\omega^{\prime})=\operatorname{id}_{A},\ \ \text{and}\ \ (\operatorname{id}_{B}\otimes\omega)\circ a^{-1}_{B,A,B}\circ(\omega^{\prime}\otimes\operatorname{id}_{B})=\operatorname{id}_{B}.$
@@
 Remark 3.2.
-[Simon, Remark 2.13, 2.16]
+[Barvels et al., 2015, Remark 2.13, 2.16]
   • (1)
@@
 Theorem 3.3.
-[Simon, Theorem 3.20] Let $\omega:A\otimes B\to\mathbf{1}$ be a non-degenerate Hopf pairing.
+[Barvels et al., 2015, Theorem 3.20] Let $\omega:A\otimes B\to\mathbf{1}$ be a non-degenerate Hopf pairing.
 The categories $^{A}_{A}\mathcal{YD}(\mathcal{C})$} and $^{B}_{B}\mathcal{YD}(\mathcal{C})$} are braided monoidal equivalent via the monoidal functors $\Omega^{\omega}$ as follows.
@@
     $\Omega^{\omega^{-}}\circ\Omega^{\omega}:{{}^{A}_{A}\mathcal{YD}(\mathcal{C})}\rightarrow{{}^{A}_{A}\mathcal{YD}(\mathcal{C})}$
-    is isomorphic to the identity functor by [Simon, Proposition 3.22].
+    is isomorphic to the identity functor by [Barvels et al., 2015, Proposition 3.22].
   • (3)
-    For completeness, we recall the construction of the functor $\Omega^{\omega}$, which is investigated in [Simon, Theorem 3.16, Corollary 3.19].
+    For completeness, we recall the construction of the functor $\Omega^{\omega}$, which is investigated in [Barvels et al., 2015, Theorem 3.16, Corollary 3.19].
 Let $\omega:A\otimes B\to\mathbf{1}$ be a non-degenerate Hopf pairing.
@@
 We now turn our attention to the construction of partial dualization.
-This construction, originally developed in [Simon, Theorem 4.4], provides a powerful tool for studying Hopf algebras in braided monoidal categories.
+This construction, originally developed in [Barvels et al., 2015, Theorem 4.4], provides a powerful tool for studying Hopf algebras in braided monoidal categories.
 Let us begin by recalling the setup. Consider a braided monoidal category $\mathcal{C}$ and a Hopf algebra $R$ in this category. A partial dualization datum is defined as follows.
@@
 We establish this result in several steps.
-First, the object $\mathcal{B}(N)^{*\operatorname{op,cop}}$ with the following operations is a Hopf algebra in $^{H}_{H}\mathcal{YD}$} by [AGn99, Lemma 2.2.3]
+First, the object $\mathcal{B}(N)^{*\operatorname{op,cop}}$ with the following operations is a Hopf algebra in $^{H}_{H}\mathcal{YD}$} by [Andruskiewitsch and Graña, 1999, Lemma 2.2.3]
 $(\mathcal{B}(N)^{*},\mu_{{\mathcal{B}(N)^{*}}},c^{-1}_{{\mathcal{B}(N)^{*}},{\mathcal{B}(N)^{*}}},\eta_{{\mathcal{B}(N)^{*}}},c_{{\mathcal{B}(N)^{*}},{\mathcal{B}(N)^{*}}}\Delta_{{\mathcal{B}(N)^{*}}},\varepsilon_{{\mathcal{B}(N)^{*}}},\mathcal{S}_{{\mathcal{B}(N)^{*}}}).$
@@
 in $^{H}_{H}\mathcal{YD}$}.
-The object $\mathcal{B}(N^{*})$ is the graded dual of $\mathcal{B}(N)$ by [defofnichols, Lemma 2.6].
+The object $\mathcal{B}(N^{*})$ is the graded dual of $\mathcal{B}(N)$ by [Meir, 2022, Lemma 2.6].
 Therefore $\operatorname{dim}(\mathcal{B}(N^{*}))=\operatorname{dim}(\mathcal{B}(N)^{*})$ since $\operatorname{dim}(\mathcal{B}(N))$ is finite-dimensional.
@@
 Let $H$ be a coquasi-Hopf algebra with bijective antipode, and $\mathcal{C}={{}^{H}_{H}\mathcal{YD}}$.
-Suppose $N=\bigoplus_{i\in I}N_{i}\in{{}^{H}_{H}\mathcal{YD}}$, and $\mathcal{B}(N_{i})$ are all finite-dimensional.
+Suppose $N=\bigoplus_{i\in I}N_{i}\in{{}^{H}_{H}\mathcal{YD}}$, and the Nichols algebras $\mathcal{B}(N_{i})$ are all finite-dimensional.
 For each $i\in I$, we have a partial dualization datum $\mathcal{A}_{i}=(\mathcal{B}(N)\overset{\pi_{i}}{\rightarrow}\mathcal{B}(N_{i}),\mathcal{B}(N_{i}^{*}),\operatorname{ev}_{\mathcal{B}(N_{i})}^{-})$.
@@
 (1) The module and comodule maps of $K(1)$ are $\mathbb{Z}$-graded and hence the infinitesimal braiding $c\in\mathrm{Aut}(K(1)\otimes K(1))$, being determined by the module and comodule maps, is $\mathbb{Z}$-graded.
-Moreover, the structure map of $K(n)$ for $n\in\mathbb{N}$ are determined by $c$ and $K(1)$, since $K(1)$ generates $K$.
+Moreover, the structure maps of $K(n)$ for $n\in\mathbb{N}$ are determined by $c$ and $K(1)$, since $K(1)$ generates $K$.
 Therefore $K$ is a $\mathbb{Z}$-graded objects in ${{}^{H}_{H}\mathcal{YD}}(\mathbb{Z}\operatorname{-Gr}\mathcal{M}_{\mathbbm{k}})$.
@@
 Suppose $R$ is an $\mathbb{N}_{0}$-graded Hopf algebra in $^{H}_{H}\mathcal{YD}$}.
-Let $X$ be a $R$-module in $\mathcal{C}$ and $Y$ be a $R$-comodule in $\mathcal{C}$.
+Let $X$ be an $R$-module in $\mathcal{C}$ and $Y$ be an $R$-comodule in $\mathcal{C}$.
 For $n\geq 0$, we define:
@@
 $\mathcal{F}^{n}Y$ $=\{y\in Y\mid\delta_{Y}(y)\in\bigoplus_{i=0}^{n}R(i)\otimes Y\}.$
-With above notations, the functor $\Omega$ relates the operator $\mathcal{F}_{n}$ to $\mathcal{F}^{n}$.
+With the above notations, the functor $\Omega$ relates the operator $\mathcal{F}_{n}$ to $\mathcal{F}^{n}$.
 Lemma 3.11.
@@
 Define an auxiliary map $\rho^{\prime}:=(\operatorname{ev}\otimes\operatorname{id}_{V})\circ a^{-1}_{\mathcal{B}(N^{*}),\mathcal{B}(N),V}\circ(\operatorname{id}_{\mathcal{B}(N^{*})}\otimes\delta_{\mathcal{B}(N)})$, which is derived from $\rho_{\mathcal{B}(N^{*})}$.
-By [rootsys, Corollary 3.3.4, Lemma 12.2.11], $(V,\rho^{\prime})$ is still a $\mathcal{B}(N^{*})$-module in $\mathcal{C}$ and
+By [Heckenberger and Schneider, [2020] ©2020, Corollary 3.3.4, Lemma 12.2.11], $(V,\rho^{\prime})$ is still a $\mathcal{B}(N^{*})$-module in $\mathcal{C}$ and
 $\mathcal{F}_{n}\Omega(V)=\mathcal{F}_{n}(V,\rho^{\prime}).$
@@
 This means, if $v\in\mathcal{F}^{n}V$, then for each $i>n$, $\rho^{\prime}(\mathcal{B}(N^{*})(i)\otimes v)=0$ by (3.5).
-On the other hand, if $v\in\mathcal{F}_{n}(V,\rho^{\prime})$, this will imply $\delta(v)\in\bigoplus_{i=0}^{n}\mathcal{B}(N)(i)\otimes V$.
-which completes the proof of (1): $\mathcal{F}_{n}\Omega(V)=\mathcal{F}_{n}(V,\rho^{\prime})=\mathcal{F}^{n}V$.
+On the other hand, if $v\in\mathcal{F}_{n}(V,\rho^{\prime})$, this will imply $\delta(v)\in\bigoplus_{i=0}^{n}\mathcal{B}(N)(i)\otimes V$, which completes the proof of (1): $\mathcal{F}_{n}\Omega(V)=\mathcal{F}_{n}(V,\rho^{\prime})=\mathcal{F}^{n}V$.
 (2) For the second statement, recall $\delta_{\mathcal{B}(N^{*})}$ is given by
@@
 We denote $\delta^{\prime}=(\operatorname{id}_{\mathcal{B}(N^{*})}\otimes\rho_{\mathcal{B}(N)})\circ a_{{\mathcal{B}(N^{*})},{\mathcal{B}(N)},V}\circ((\operatorname{ev}^{-})^{\prime}\otimes\operatorname{id}_{V})$, which origins from $\delta_{\mathcal{B}(N)}$.
-Again by [rootsys, Corollary 3.3.6, Lemma 12.2.11]
+Again by [Heckenberger and Schneider, [2020] ©2020, Corollary 3.3.6, Lemma 12.2.11]
 $\mathcal{F}^{n}(V,\delta^{\prime})=\mathcal{F}^{n}\Omega(V).$
@@
 In this section, we study projection of Nichols algebras in $^{G}_{G}\mathcal{YD}^{\Phi}$}, which will play an important role when considering reflection of simple Yetter-Drinfeld modules.
-The proofs of the some lemmas and theorems in this section, while are conceptually parallel to those in the Hopf algebra setting HSdualpair, takes a different path due to the non-strict monoidal structure.
+The proofs of some lemmas and theorems in this section, while are conceptually parallel to those in the Hopf algebra setting Heckenberger and Schneider (2013a), take a different path due to the non-strict monoidal structure.
 4.1 The structure of the space of coinvariant $K$
-To study the projection properties of Nichols algebras, we must first examine their structure in the cateogry $^{G}_{G}\mathcal{YD}^{\Phi}$}.
+To study the projection properties of Nichols algebras, we must first examine their structure in the category $^{G}_{G}\mathcal{YD}^{\Phi}$}.
 For any $M,N\in{{}^{G}_{G}\mathcal{YD}^{\Phi}}$, there is a canonical surjection of braided Hopf algebra in $^{G}_{G}\mathcal{YD}^{\Phi}$}
@@
 Lemma 4.5.
-[rootsys, Lemma 1.3.13, Proposition 1.3.14]
+[Heckenberger and Schneider, [2020] ©2020, Lemma 1.3.13, Proposition 1.3.14]
 $(1)$ Assume that $C$ is connected, that is, $C^{0}$ is one dimensional. Then $I_{C}\subseteq C$ is a coideal of $C$.
@@
 Lemma 4.6.
-[rootsys,Proposition 13.2.3 Let $C$ be an $\mathbb{N}_{0}$-graded coalgebra, $X$ an $\mathbb{N}_{0}$-graded left $C$-comodule, and $Y$ a $C$-subcomodule of $X$.
+[Heckenberger and Schneider, [2020] ©2020,Proposition 13.2.3] Let $C$ be an $\mathbb{N}_{0}$-graded coalgebra, $X$ an $\mathbb{N}_{0}$-graded left $C$-comodule, and $Y$ a $C$-subcomodule of $X$.
 Let $k>0$ be an integer.
@@
 $\delta:\mathcal{B}(M\oplus N)\xrightarrow{\Delta_{\mathcal{B}(M\oplus N)}}\mathcal{B}(M\oplus N)\otimes\mathcal{B}(M\oplus N)\xrightarrow{\pi_{\mathcal{B}(N)}\otimes\operatorname{id}}\mathcal{B}(N)\otimes\mathcal{B}(M\oplus N).$
-Now let $X\in L$, then $\delta(X)\in\mathcal{B}(N)\otimes L$ by Lemma 4.4, which implies that L inherits an $\mathbb{N}_{0}$-graded $\mathcal{B}(N)$-comodule via the restriction of $\delta:L\rightarrow\mathcal{B}(N)\otimes L$.
+Now let $X\in L$, then $\delta(X)\in\mathcal{B}(N)\otimes L$ by Lemma 4.4, which implies that L inherits an $\mathbb{N}_{0}$-graded $\mathcal{B}(N)$-comodule structure via the restriction of $\delta:L\rightarrow\mathcal{B}(N)\otimes L$.
 Suppose $X\in L(n)$ for $n\geq 1$, consider the component map $\delta_{n-1,1}:L(n)\rightarrow\mathcal{B}(N)(n-1)\otimes L(1)=\mathcal{B}(N)(n-1)\otimes M.$ By Lemma 4.4, $\Delta_{\mathcal{B}(M\oplus N)}(X)=X\otimes 1+\mathcal{B}(N)\otimes L$, it follows that
@@
 In particular, we have $K(1)=K\cap(\mathcal{B}(M\oplus N)\#\mathbbm{k}G)(1)=K\cap(\mathcal{B}(M\oplus N)(1))=\operatorname{ad}(\mathcal{B}(N))(M)=L$.
-Hence $K$ is generated its degree one component $K(1)$.
+Hence $K$ is generated by its degree one component $K(1)$.
 It remains to show that $P(K)=K(1)$.
@@
 Applying the map $\mathrm{id}\#\varepsilon\otimes\mathrm{id}$, we obtain
-$\Delta_{\mathcal{B}(M\oplus N)}(u)=(\operatorname{id}\#\varepsilon\otimes\operatorname{id})\Delta(u)=1\otimes u+u\otimes 1,$
+$\Delta_{\mathcal{B}(M\oplus N)}(u)=(\operatorname{id}\#\varepsilon\otimes\operatorname{id})\Delta_{\mathcal{A}(M\oplus N)}(u)=1\otimes u+u\otimes 1,$
 which implies that $u$ is a primitive element in $\mathcal{B}(M\oplus N)$.
@@
 $W_{i}^{\prime}:=\mathcal{B}(N)\rhd M_{i}^{\prime}.$
-Using a method similar to that in Lemma 4.3, one verifies that $W_{i}^{\prime}$ is an $\mathbb{Z}$-graded $\mathcal{A}(N)$-subcomodule.
+Using a method similar to that in Lemma 4.3, one verifies that $W_{i}^{\prime}$ is a $\mathbb{Z}$-graded $\mathcal{A}(N)$-subcomodule.
 Moreover, $\mathcal{A}(N)\rhd(\mathcal{B}(N)\rhd M_{i}^{\prime})\subseteq\mathcal{A}(N)\rhd M_{i}^{\prime}=\mathcal{B}(N)\rhd M_{i}^{\prime}$.
@@
 $R:=K\#\mathcal{B}(N)=(K\#(\mathcal{B}(N)\#\mathbbm{k}G))^{\operatorname{co}\mathbbm{k}G}$
-is a $\mathbb{N}_{0}$-graded Hopf algebra in $^{G}_{G}\mathcal{YD}^{\Phi}$}. Its degree 0’s part is $\mathbbm{k}$, and its degree $1$’s part is $M\oplus N$.
+is a $\mathbb{N}_{0}$-graded Hopf algebra in $^{G}_{G}\mathcal{YD}^{\Phi}$}. Its degree 0 component is $\mathbbm{k}$, and its degree $1$ component is $M\oplus N$.
 We now show that $R$ is generated by $M\oplus N$ as an algebra in $^{G}_{G}\mathcal{YD}^{\Phi}$}.
@@
 $\mathcal{F}_{\theta}(M)=\{R_{i_{1}}(...(R_{i_{l}}(M))...)\mid l\in\mathbb{N}_{0},\ i_{1},..,i_{l}\in\mathbb{I}\}.$
-Such a question arises naturally: how do the irreducibility properties behave under reflections? The following result provides a reassuring answer.
+A natural question arises: how do the irreducibility properties behave under reflections? The following result provides a reassuring answer.
 Lemma 5.3.
@@
 We want to prove $\langle L\rangle$ is a Yetter-Drinfeld submodule of $\operatorname{ad}\mathcal{B}(M_{i})(M_{j})$ in $^{A(M_{i})}_{A(M_{i})}\mathcal{YD}$}.
-To see this, take any homogeneous $A\in\mathcal{A}(N)$.
+To see this, take any homogeneous $A\in\mathcal{A}(M_{i})$.
 Then for all homogeneous $X\in L$, by Lemma 2.6,
 $\operatorname{ad}(A)$ $(\langle f,X_{-1}\rangle X_{0})=\langle f,X_{-1}\rangle\operatorname{ad}(A)(X_{0})$
-$=\langle f,s(\mathcal{S}(A_{1}),(\operatorname{ad}(A_{4})(X_{0}))_{-1}A_{6})t(A_{3},X_{-1})\mathcal{S}(A_{2})(\operatorname{ad}(A_{4})(X_{0}))_{-2}A_{5})\rangle(\operatorname{ad}(A_{4})(X_{0}))_{0}.$
+$=\langle f,s(\mathcal{S}(A_{1}),(\operatorname{ad}(A_{4})(X_{0}))_{-1}A_{6})t(A_{3},X_{-1})\mathcal{S}(A_{2})((\operatorname{ad}(A_{4})(X_{0}))_{-2}A_{5})\rangle(\operatorname{ad}(A_{4})(X_{0}))_{0}.$
 Note that $A$ and $X$ are homogeneous, thus if $(\operatorname{ad}(A_{4})(X_{0}))_{-1}\notin G$, then this term equals zero. Therefore
@@
 Hence $\operatorname{ad}(A)(\langle f,X_{-1}\rangle X_{0})\in\langle L\rangle.$ This ensures the first axiom of a Yetter-Drinfeld module, and the third axiom holds automatically.
-Since $\operatorname{ad}\mathcal{B}(M_{i})(M_{j})$ is generated by $\operatorname{ad}(M_{i})^{m_{ij}}(M_{j})$ as a Yetter-Drinfeld module and is irreducible in $^{A(M_{i})}_{A(M_{i})}\mathcal{YD}$} by Lemma 4.8 (2).
-Then $L$ must generate the entire space.
-This forces $L=\operatorname{ad}(M_{i})^{m_{ij}}(M_{j}),$ establishing the irreducibility.
+Therefore $\langle P\rangle$ is a Yetter-Drinfeld submodule of $\operatorname{ad}\mathcal{B}(M_{i})(M_{j})$ in $^{A(M_{i})}_{A(M_{i})}\mathcal{YD}$}.
+Since $\operatorname{ad}(\mathcal{B}(M_{i}))(M_{j})$ is irreducible in $^{A(M_{i})}_{A(M_{i})}\mathcal{YD}$} by Lemma 4.8 (2), and $P\neq 0$, we have $\langle P\rangle=\operatorname{ad}(\mathcal{B}(M_{i}))(M_{j})$.
+It is direct to see that $\langle P\rangle=\bigoplus_{i=0}^{m_{ij}}\langle P\rangle\cap\mathcal{B}^{i}(M_{i}\oplus M_{j})$.
+Then
+$P=\langle P\rangle\cap\mathcal{B}^{m_{ij}}(M_{i}\oplus M_{j})=\operatorname{ad}(\mathcal{B}(M_{i}))(M_{j})\cap\mathcal{B}^{m_{ij}}(M_{i}\oplus M_{j})=\operatorname{ad}(M_{i})^{m_{ij}}(M_{j}).$
+This establishes the irreducibility. ∎
 Lemma 5.4.
-Suppose $M\in\mathcal{F}_{\theta}$ and $M$ admits $i$-th reflection for each $i\in\mathbb{I}$.
+Suppose $M\in\mathcal{F}_{\theta}$ and $M$ admits the $i$-th reflection for each $i\in\mathbb{I}$.
 We define $a_{ii}^{M}=2$ for all $1\leq i\leq\theta$ and define $a_{ij}^{M}=-m_{ij}^{M}$.
@@
 Lemma 5.5.
-Suppose $M\cong N$ in $\mathcal{F}_{\theta}$, if $M$ admits the $i$-th reflection for some $i\in\mathbb{I}$, so is $N$.
+Suppose $M\cong N$ in $\mathcal{F}_{\theta}$, if $M$ admits the $i$-th reflection for some $i\in\mathbb{I}$, so does $N$.
 Furthermore, $R_{i}(M)\cong R_{i}(N)$ and $a_{ij}^{M}=a_{ij}^{N}$ for each $j\in\mathbb{I}$.
 Proof.
-We denote $\phi:M\rightarrow N$ be isomorphism of tuples with each component $\phi_{j}:M_{j}\rightarrow N_{j}$ an isomorphism of Yetter-Drinfeld modules.
+Let $\phi:M\rightarrow N$ be an isomorphism of tuples with each component $\phi_{j}:M_{j}\rightarrow N_{j}$ an isomorphism of Yetter-Drinfeld modules.
 We first demonstrate that $\phi_{i}$ inducing a unique morphism $\mathcal{B}(\phi_{i}):\mathcal{B}(M_{i})\rightarrow\mathcal{B}(N_{i})$.
@@
 5.2 Main Theorem
-In this subsection, we always fix a tuple $M=(M_{1},\ldots,M_{\theta})\in\mathcal{F}_{\theta}$, with each component $M_{i}$ is irredcuible, $i\in\mathbb{I}$.
-Furhtermore, we assume $\mathcal{B}(M_{i})$ is finite-dimensional for each $i\in\mathbb{I}$.
+In this subsection, we always fix a tuple $M=(M_{1},\ldots,M_{\theta})\in\mathcal{F}_{\theta}$, with each component $M_{i}$ being irreducible, $i\in\mathbb{I}$.
+Furthermore, we assume $\mathcal{B}(M_{i})$ is finite-dimensional for each $i\in\mathbb{I}$.
 By Corollary 3.8, we have such a braided tensor equivalence for each $i\in\mathbb{I}$.
 $\Omega_{i}:{{}^{A(M_{i})}_{A(M_{i})}\mathcal{YD}}\cong{{}^{A(M_{i}^{*})}_{A(M_{i}^{*})}\mathcal{YD}}.$ (5.1)
-The main result of this paper proves that the reflection of Nichols algebra in $^{G}_{G}\mathcal{YD}^{\Phi}$} gives rise to a semi-Cartan graph.
+The main result of this paper proves that the reflection of Nichols algebras in $^{G}_{G}\mathcal{YD}^{\Phi}$} gives rise to a semi-Cartan graph.
 The first lemma provides some relations between reflection of $M$ and the braided tensor equivalence $\Omega_{i}$.
@@
 The next theorem gives a natural explanation of reflections of tuples of Yetter-Drinfeld modules.
-Although the proof strategy parallels the Hopf algebra case [rootsys, Theorem 13.4.9], we include the details here for the sake of completeness and the reader’s convenience.
+Although the proof strategy parallels that of the Hopf algebra case [Heckenberger and Schneider, [2020] ©2020, Theorem 13.4.9], we include the details here for the sake of completeness and the reader’s convenience.
 Theorem 5.7.
-With above assumptions on $M$. For each $i\in\mathbb{I}$, suppose $M$ admits the $i$-th reflection, then there is an isomorphism of Hopf algebras in $^{G}_{G}\mathcal{YD}^{\Phi}$}:
+With the above assumptions on $M$. For each $i\in\mathbb{I}$, suppose $M$ admits the $i$-th reflection, then there is an isomorphism of Hopf algebras in $^{G}_{G}\mathcal{YD}^{\Phi}$}:
 $\Theta_{i}:\mathcal{B}(R_{i}(M))\cong\Omega_{i}\left(\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}\right)\#\mathcal{B}(M_{i}^{*}).$ (5.2)
@@
 6 Applications
-Building upon the framework developed in the preceding sections, we now present some applications of the semi-Cartan graph theory for Nichols algebras over certain coquasi-Hopf algebras.
+Building upon the framework developed in the preceding sections, we now present some applications of the semi-Cartan graph theory to Nichols algebras over certain coquasi-Hopf algebras.
 We establish several criteria that connect the finite-dimensionality of a Nichols algebra to the finiteness and structure of its associated semi-Cartan graph and Weyl groupoid.
-Then we apply this theory to provide a new proof for the key result in [huang2024classification].
+Then we apply this theory to provide a new proof for the key result in [Huang et al., ].
 6.1 A criterion for finite-dimensional Nichols algebras
@@
 Proof.
-If we regard $\mathcal{B}(M)$ as an $\mathbb{N}_{0}$-graded object in $^{G}_{G}\mathcal{YD}^{\Phi}$}.
-By [defofnichols, Lemma 2.6] the graded dual of $\mathcal{B}(M)$ is isomorphic to $\mathcal{B}(M^{*})$.
+If we regard $\mathcal{B}(M)$ as an $\mathbb{N}_{0}$-graded object in $^{G}_{G}\mathcal{YD}^{\Phi}$}, by [Meir, 2022, Lemma 2.6] the graded dual of $\mathcal{B}(M)$ is isomorphic to $\mathcal{B}(M^{*})$.
 We thus obtain an isomorphism of $\mathbb{N}_{0}$-graded objects in $^{G}_{G}\mathcal{YD}^{\Phi}$}:
@@
 $\mathcal{B}(R_{i}(M))\otimes\mathcal{B}(R_{i}(M)^{*})\cong K_{i}\otimes\mathcal{B}(M_{i}^{*})\otimes K_{i}^{\operatorname{gr-dual}}\otimes\mathcal{B}(M_{i}).$
-Since $A\otimes B\cong B\otimes A$ as $\mathbb{Z}^{\theta}$-graded vector spaces. the supports on both sides coincide, which establishes the first claim.
+Since $A\otimes B\cong B\otimes A$ as $\mathbb{Z}^{\theta}$-graded vector spaces, the supports on both sides coincide, which establishes the first claim.
 The second claim follows directly from the definition of $s_{i}^{M}$. ∎
@@
 $\Delta^{X,\operatorname{re}}=\{\omega(\alpha_{i})\in\mathbb{Z}^{\mathbb{I}}\mid\omega\in\operatorname{Hom}(\mathcal{W}(\mathcal{G}),X),i\in\mathbb{I}\}$
-is called the set of real roots of $\mathcal{G}$ at $X$. We call the semi-Cartan graph $\mathcal{G}$ is finite, if $\Delta^{X,\operatorname{re}}$ is finite for all $X\in\mathcal{X}$.
+is called the set of real roots of $\mathcal{G}$ at $X$. We call the semi-Cartan graph $\mathcal{G}$ finite, if $\Delta^{X,\operatorname{re}}$ is finite for all $X\in\mathcal{X}$.
 We now relate the finiteness of the Nichols algebra to the structure of the associated semi-Cartan graph.
@@
 We recall that every element $\omega\in\operatorname{Hom}(\mathcal{W}(\mathcal{G}),M)$ can be expressed as an iteration of simple reflections.
-It follows that Thus $\Delta^{M\operatorname{re}}\subseteq\operatorname{Supp}\mathcal{B}(M)\cup\operatorname{Supp}\mathcal{B}(M^{*})$.
+It follows that $\Delta^{M\operatorname{re}}\subseteq\operatorname{Supp}\mathcal{B}(M)\cup\operatorname{Supp}\mathcal{B}(M^{*})$.
 Now, since $\mathcal{B}(M)$ is finite-dimensional, the set $\operatorname{Supp}\mathcal{B}(M)$ is finite.
@@
 Lemma 6.5.
-[HS10, Lemma 5.1] Suppose $\mathcal{G}=(\mathbb{I},\mathcal{X},r,(A^{X})_{X\in\mathcal{X}})$ is a semi-Cartan graph.
+[Heckenberger and Schneider, 2010b, Lemma 5.1] Suppose $\mathcal{G}=(\mathbb{I},\mathcal{X},r,(A^{X})_{X\in\mathcal{X}})$ is a semi-Cartan graph.
 Then $\mathcal{G}$ is finite semi-Cartan graph if and only if the Weyl groupoid $\mathcal{W}(\mathcal{G})$ is finite.
@@
 If $A^{X}=A^{Y}$ for all $X,Y\in\mathcal{X}$, then $\mathcal{G}$ is called a standard semi-Cartan graph.
-The proof of next theorem is similar to the Hopf algebra situations, we write down here for completeness.
+The proof of the next theorem is similar to the Hopf algebra situation; we include it here for completeness.
 Theorem 6.7.
@@
 6.2 A class of infinite-dimensional Nichols algebras
-The purpose of this section is to give a new proof of Proposition 4.1 in huang2024classification through applying our previous observations.
+The purpose of this section is to give a new proof of Proposition 4.1 in Huang et al. through applying our previous observations.
 Let $G=\mathbb{Z}_{m_{1}}\times\mathbb{Z}_{m_{2}}\times\mathbb{Z}_{m_{3}}=\langle g_{1}\rangle\times\langle g_{2}\rangle\times\langle g_{3}\rangle$ with $m_{1}\mid m_{2}$, $m_{2}\mid m_{3}$, and
@@
 (i) Note that $m_{1},m_{2},m_{3}$ must be even. Indeed, from the relation $X_{i}=e^{m_{1}}\rhd X_{i}=(-1)^{m_{1}}X_{i}$, it follows that $m_{1}$ is even. Similarly $m_{2},m_{3}$ are even.
-(ii) $\mathcal{B}(V_{1})$, $\mathcal{B}(V_{2})$, $\mathcal{B}(V_{3})$ are finite-dimensional Nichols algebras by [huang2024classification, Proposition 2.12].
-Moreover, $\mathcal{B}(V_{1}\oplus V_{2})$, $\mathcal{B}(V_{1}\oplus V_{3})$, $\mathcal{B}(V_{2}\oplus V_{3})$ are finite-dimensional by [nondiagonal Proposition 5.1].
+(ii) $\mathcal{B}(V_{1})$, $\mathcal{B}(V_{2})$, $\mathcal{B}(V_{3})$ are finite-dimensional Nichols algebras by [Huang et al., , Proposition 2.12].
+Moreover, $\mathcal{B}(V_{1}\oplus V_{2})$, $\mathcal{B}(V_{1}\oplus V_{3})$, $\mathcal{B}(V_{2}\oplus V_{3})$ are finite-dimensional by [Huang et al., 2018, Proposition 5.1].
 (iii) The requirement $\operatorname{dim}(V_{1})=\operatorname{dim}(V_{2})=\operatorname{dim}(V_{3})=2$ is in fact necessary, otherwise the corresponding Nichols algebra must be infinite-dimensional.
 Theorem 6.9.
-[huang2024classification, Proposition 4.1] With above notations, $\mathcal{B}(V_{1}\oplus V_{2}\oplus V_{3})$ is infinite-dimensional.
-It should be emphasized that the above theorem plays the key role in that paper and the original proof relays on heavy computations.
-The purpose of the following context is to give a new proof of this theorem through applying reflection theory of Nichols algebras.
+[Huang et al., , Proposition 4.1] With above notations, $\mathcal{B}(V_{1}\oplus V_{2}\oplus V_{3})$ is infinite-dimensional.
+It should be emphasized that the above theorem plays a key role in that paper and the original proof relies on heavy computations The purpose of the following discussion is to give a new proof of this theorem by applying the reflection theory of Nichols algebras.
 To simplify the proof, we proceed with the following reduction.
@@
 be a $3$-cocycle on $H$, where $0\leq i_{l},j_{l},k_{l}\leq 1$, $1\leq l\leq 3$.
-By huang2024classification Lemma 3.5, we may give a complete list of irreducible Yetter-Drinfeld modules.
+By [Huang et al., , Lemma 3.5], we may give a complete list of irreducible Yetter-Drinfeld modules.
 In particular, restricting to those irreducible Yetter–Drinfeld modules that generate finite-dimensional Nichols algebras, the isomorphism classes are given by the set
@@
 This implies $W_{i}^{*}\in S$ as well.
-Moreover, $\mathcal{B}(W_{i})\cong\mathcal{B}(W_{i}^{*})$ as Hopf algebras in $^{H}_{H}\mathcal{YD}$}.
+Moreover, $\mathcal{B}(W_{i})\cong\mathcal{B}(W_{i}^{*})$ as Hopf algebras in $^{H}_{H}\mathcal{YD}^{\phi}$}.
 The application of semi-Cartan graph theory is the following result.
@@
 Proof.
-This proof involves extensive but analogous computations. We begin by establishing that $\operatorname{ad}(W_{1})^{2}(W_{2})=0.$ To this end, we compute the coproducts:
+This proof involves extensive but analogous computations.
+Step1
+We begin by establishing that $\operatorname{ad}(W_{1})^{2}(W_{2})=0.$ To this end, we compute the coproducts:
 $\Delta\left(\operatorname{ad}(X_{1})\left(Y_{1}\right)\right)=1\otimes\operatorname{ad}({X_{1}})\left(Y_{1}\right)+\operatorname{ad}({X_{1}})\left(Y_{1}\right)\otimes 1+X_{1}\otimes\left(Y_{1}+Y_{2}\right),$
@@
 $\operatorname{ad}(W_{2})(W_{4})\cong W_{1},\ \operatorname{ad}(W_{4})(W_{2})\cong W_{1}.$
+Step2
 We now proceed to prove $\operatorname{ad}(W_{1})^{2}(W_{3})=0$:
@@
 $\operatorname{ad}(X_{2})^{2}(Z_{2})$ $=X_{2}(X_{2}Z_{2}-Z_{1}X_{2})-h_{1}\rhd(X_{2}Z_{2}-Z_{1}X_{2})X_{2}=0.$
-This combines (6.8) implies $\operatorname{ad}(W_{1})^{2}(W_{3})=0$. As in the first case, we can derive the following identities:
+This argument, combined with (6.8), implies $\operatorname{ad}(W_{1})^{2}(W_{3})=0$. As in the first case, we can derive the following identities:
 $\operatorname{ad}(W_{3})^{2}(W_{1})=$ $\operatorname{ad}(W_{1})^{2}(W_{5})=\operatorname{ad}(W_{3})^{2}(W_{5})=\operatorname{ad}(W_{5})^{2}(W_{1})=\operatorname{ad}(W_{5})^{2}(W_{3})=0.$
@@
 $\operatorname{ad}(W_{3})(W_{5})\cong\operatorname{ad}(W_{5})(W_{3})\cong W_{1}.$
+Step3
 Now we turn to proving $\operatorname{ad}(W_{2})^{2}(W_{3})=0$:
@@
 $\operatorname{ad}(W_{3})(W_{6})\cong\operatorname{ad}(W_{6})(W_{3})\cong W_{2}.$
+Step4
 The final case to confirm is $\operatorname{ad}(W_{4})^{2}(W_{5})=0$:
@@
 $R_{i_{l-1}}(...(R_{i_{1}}(W))...)=(k^{\prime\prime},k^{\prime}k^{\prime\prime},k^{\prime\prime}).$
-All cases are leading to contradictions. This completes the induction. ∎
+All of these cases lead to contradictions. This completes the induction. ∎
 Proof of Theorem 6.11.
@@
 $F(g\rhd v_{i})$ $=\pi(g)\rhd F(v_{i}).$
-By [QQG, Lemma 4.4], we deduce that
+By [Huang et al., $20$, Lemma 4.4], we deduce that
 $\mathcal{B}(V)\cong\mathcal{B}(W)$
-in the sense of [QQG, Definition 4.3]. It follows that $\mathcal{B}(V)$ is infinite-dimensional. ∎
-Remark 6.15.
-For the sake of simplicity, this article studies only the cosemisimple pointed coquasi-Hopf algebras under minor conditions.
-We hope to extend this theory to all coquasi-Hopf algebras with bijective antipodes.
+in the sense of [Huang et al., $20$, Definition 4.3]. It follows that $\mathcal{B}(V)$ is infinite-dimensional. ∎
 Acknowledgment This work is supported by the National Key R&D Program of China